A Complete Bibliography of Publications in Technometrics for the decade 2010–2019

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

13 December 2019 Version 1.20

Title word cross-reference

#9 [Jen11].

$100.00 [Mar19a]. 2^{k−p} [LL16]. 3 [CG16, MQ11, VGS19, ZQ18, ZZD^{+}14]. $46.36 [Mar19a]. C[0, 1] [BLK19]. C_{p} [CHS11]. D [ZDH14, LMW19]. G [HN18]. I_{λ} [HN18]. k [Wył10]. N [RB10]. P [AGV12].

-D [MQ11]. -Exploring [CCCC13]. -Optimal [HN18, LMW19, ZDH14]. -Optimality [HN18]. -QPSO [LMW19]. -sample [Wył10]. -Splines [AGV12]. -Valued [BLK19].

1 [CZ15a]. 10.1007 [Ahm19b]. 10.1007/978 [Ahm19b]. 10.1007/978-3-319-91385-8 [Ahm19b]. 11th [Ano11u]. 1st [Sau10, Sau11, Sau12].

2009 [Cob10]. 2014 [Ano16a]. 2015 [Ano16c]. 2016 [Ano16k, Ano17a, Ano17b]. 2017 [Ano17i]. 2018 [Ano18g]. 2nd [Ahm19b, Ano10s, Ano10g, Ano10n, Ano10o, Ano10i, Ano10l, Ano11d, Ano11p, Ano11h, Ano11j, Ano11i, Ano11c, Ano11k, Ano12n, Ano12i, Ano12c, Ano12d, Ano12p, Ano12o, Guo19, Jar12b, Oli11a, Pfa12, Qia10b, Qu12, Rut12, San11].

33rd [Ano11u]. 39 [Hor11b]. 3rd [Ano12e, Ano12q, Hly11, Mar19b, Ors19].

4th [Ano10r, Ano11g].

5th [Ano12h].

6 [Mar19a].

8 [Ahm19b].

978 [Mar19a]. 978-1-4987-7606-6 [Mar19a]. 978-3-319-91384-1 [Ahm19b]. 978-3-319-91385-8 [Ahm19b, Ahm19b].

Abdelmomem [Ano12h]. Aberration [RB10]. Accelerated [LHTD18, Liu12, TL16, XKHY18]. Accelerating [NHQ17]. Acceptance [LRL^{+}12b, LRL^{+}12a]. Accessibility [CS12, DF12, JSQ12, JS12a, JS12b, LW12, WSZ12]. Accomplish [Nat12]. Account [DLY^{+}17]. Accumulated [YZW19]. Accuracy [XQW13]. Accurate [PT14]. across [CCCC13]. Activity [ZHOY13]. Actuarial [Mye11b]. Adaptive [Che11a, KKS^{+}14, LMS15, RMZ15, SWF16, TH18, XWL18, Gol12a]. Addendum [Got10b, Got10a]. Additive [AGV12, DLLR17, PH16, SHD18]. Additivity [Fra19]. Adjustment [JY15]. Adjustments [CJ14]. Admissible [WMT12]. ADMM [GFK^{+}18]. Adrian [Fot10b, Hor11a]. Advanced [Hor10b, Hor10a]. Advances [Ano12j, Gra13, Ano10d, Ano10e, Lip19a]. Aeilko [Ano10r]. Afifi [Ano12h]. AG [Ahm19b]. Aggregate [CY17, SWW^{+}18]. Aggregated [DGS13]. Aggregation [SWW^{+}18]. Agreement [SAC17]. Aitken [Bur11]. Akinori [Ano10j]. al [Han10]. al. [CL16]. Alain [Bai11]. Alan [Bar11, Bur12]. Albrecht [AA12]. Alex [Ano12r]. Alexander [Mye11c]. Alexey [Mye11c]. Algebraic [Hor11b, Ano10f]. Algorithm [CG16, cCOP^{+}13, DASS17, LKS16, NHQ17, XDHQ16, XC10]. Algorithmic [Qiu12]. Algorithms [Par11, SdCR12]. Aliasing [JN11]. Aliev [Sen11]. Alignment [SRZ^{+}19]. All-in-One [HRV19]. Allan [San11]. Allison [Bar11]. Allocation [LCAc13, SNM10, TL16]. Allowing [HRV19]. Almost [GACH10]. Almost-Parallel [GACH10]. Almost-Series [GACH10]. Along [LM15]. Alpaydin [Ano10l]. Alternative [PJW13]. Alternatives [KSB16]. Ambiguity [WC12]. Ameliorated [ZX17]. Analyses [LPVW13, Pfa12].

Analysing [Hor11a]. Analysis [Ahm19a, ANAC13, AGMRS11, ACH14, Ano10o, Ano10d, Ano10e, Ano11h, Ano11f, Ano11e, Ano11u, Ano12n, Ano12j, Ano12g, Ano12o, BFK^{+}12, Bur10, CMR10, CM17, CJL14, CCCC13, CHWE11, CFF13, CRCH18, DS17, Dav13, EDR16, FK14, Fre12, Fre13, GLLS14, HS14, Hec11, Hly10a, HP12, HJM15, JAP15, Ke10, Ke11a, LKS16, LD11, LSP13, LS13, Lip10b, Lip12a, LWM13, Mag11, MZ13, Mai11, MW19, Mar19a, Nku10, Oli10c, Oli12a, PZQ16, Pie13, PHBF17, Qia10b, Qiu12, Ruk15, SL18, SDR15, WYHT18, WMP18, Wie11, XKHY18, XQW13, XC10, YPP19, YZW19, Yu12, ZHOY13, ZM19, ZS15, ZCM^{+}18, dCC11, Ano10s, Ano10j, Ano11c, Ano12h, Ano12q, Ano10g, Ano12d, Gho19, Hor10a, Oli10a]. Analysis-of-Marginal-Tail-Means [MW19]. Analytic [Ano11i, Ban10a]. Analytical [Tan18]. Analytics [McC12a, Mye12b, Nku10]. Analyzing [GG12, Gho10a, Sar11]. Andersen [Lip11c]. Andreas [Hor11f]. Andrew [Ban10a, Che12a, Li10, Mye12b, Ng11]. Andrey [Mye11c]. Angel [Ano12g, Mar19a]. Angle [CM11, LM13]. Anirban [Ano11m, McC12b]. Annealing [WB11]. Annette [Hor11a]. Announcement [Ano10c, Ano11b, Ano13b, Ano14a, Ano16a, Ano16c, Ano17b, Ano19d, Chi12a, Cho11, Cho12, Edi15, Mit10]. Annual [Ano11u]. Anomalous [NHB^{+}13]. Anomaly [Ste10b, YPS17]. ANOVA [Dri10]. Antedependence [Gho12a]. Antón [Gho12a]. Application [AGMRS11, AGRI19, CGR^{+}15, CM14, CGB^{+}14, CS12, DF12, FK14, GGP10, HDM^{+}15, JSQ12, JS12a, JS12b, KKS^{+}14, LW12, LZHZ18, MGSC18, RKE10, SWW^{+}18, SRZ^{+}19, WSZ12, ZCW19, ZS15]. Applications [Ano10d, Ano10e, Ano12i, DNV16, GJL15, HNKF18, Hec10, Hly10b, HW10, ILS^{+}17, Jar12a, JK11, Mal19, Mye11c, Qia10b, SHL15, TW13, You10, YZW12, Ano11o, Hly10a, Hor10b, Hor10c, Laz10, Lip12b, Mye11b, Mye12d, Nag11, Nku10, Par11, Lip19a]. Applied [Ano10r, CJL14, cCOP^{+}13, Fot10b, Yu12, Ano11d, Ano12e, Lip19b, Ors19]. Applying [JFEH14]. Approach [And11, Apl12, BA18, Car14, DS17, HLP13, JAP15, Lip12a, Oli10a, PZQ16, PGK16, QLS^{+}19, QX14, RHK11, SAH^{+}17, TS12, TW13, WLWZ16, ZW17, ZQZ11, ZSGM14, ZZZ^{+}15, Atk12, Bur11, Neu12b, Qiu12]. Approaches [Gru11, WMI11]. Appropriate [SWWR14]. Approximate [EW10, OW17, SD15]. Approximation [Bor12, Jos13, She17]. Approximations [Gui18, Tan18]. Archetypal [DS17]. Aris [Ano10k]. Array [HQH16, ZS15]. Array-Based [HQH16]. Arrays [VGS19]. Assay [LSP13]. Assessing [DSM13, HLBS19, LD11]. Assessment [Hor12b, JLLM^{+}16]. Assisted [LMT14]. Asymmetric [QAHS14]. Asymptotics [Gho12b]. ATM [MW19]. Atmospheric [BFK^{+}12]. Augmented [CL16, GGL^{+}16a, PGK16]. Augmenting [YT16]. Autobiography [Ano10t]. Autocorrelated [TH10]. Autocovariance [NAEK15]. Automated [Ke10]. Automatic [RMZ15, TMMH17]. Autoregression [NS15]. Autoregressive [BLK19, NSS11]. Auxiliary [HM10]. Availability [CMR10]. Available [DSM13]. Aven [Hor12b].

Averaging [Fra19, RKE10]. Award [Ano13b, Ano14a, Cho11, Cho12, Mit10, Ano16k, Ano16a, Ano16c, Ano17i, Ano17b, Ano18g, Bre13, Bre14b, Hla15].

B [Ano12i, Fot10b, Fot10a, Hor11e, Li10, Yu12, YCZ17]. B-Splines [YCZ17]. Babatunde [Cin11]. Background [YPS17]. Baddeley [Fot10b]. Bai [Ano12n]. Bailer [Ng12]. Bailey [Neu10a]. Balakrishnan [Ano10e]. Balanced [Hof10]. Band [TMAZ18]. Bank [LVB13]. Banking [Han10]. Baragona [Cha12]. Barnett [Hor11a, Mye12a]. Basawa [Ano11l]. Based [cCOP^{+}13, CY17, CWW17, CGB^{+}14, DASS17, FPG19, LC15, GLLS14, HLP13, HM10, HM13, HQH16, JN17, JK11, Jos12a, LZ15, Liu12, LACR11, MDS12, PGRC13a, SAH^{+}17, SG18, SCH17, XDHQ16, YXTS12, ZZD^{+}14, ZJT11, KMM11]. Basic [Jen10]. Basis [CBD^{+}17, SWMW13]. Basu [Ano11m]. Batch [DASS17]. Battaglia [Cha12]. Bayes [LPVW13]. Bayesian [Ano11f, Ban10b, Bur11, Gru11, Hey12, Lip19f, Apl12, AHS16, BWMG14, CMR10, Car14, CGR^{+}15, CWW17, DNV16, EDR16, FK14, Fra19, GB18, GW13, GWN13, HN12, HS14, HLSR15, HP12, Jos12a, KRSA15, KKS^{+}14, LHTD18, LG13, LHDP14, LKS16, MMEDG14, MBR14, OW17, PD11, PHK^{+}18, PH16, PR17, SHD18, TS12, TW13, TH10, WC12, WLWZ16, YZW19, ZCM^{+}18, Ban11, Bar10, Boo10a, Che11a, Guo19, Li10, Lip12a, Mag11]. Bayesian/Likelihood [Bur11]. Becker [Han10]. Beginner [Ano12c, Bai11]. Behaved [LMW19]. Behavior [QX14, TMAZ18]. Behavioural [Ahm19c]. Belief [Lip19f]. Benefits [EJLN17]. Berger [Ano11f]. Bernd [Hor11b]. Bernhard [Ano10g]. Bernstein [Tan16]. Berry [Che11a]. Best [LCAc13]. Bettencourt [WMI11]. Between [FQ18, QWKR14]. Bhat [Hly10a]. Bhatia [Lu11]. Bias [Gol11a]. Bidirectional [QL19]. Biennial [Ano11u]. Bifocal [WC12]. Big [ILS^{+}17, LKS16, Qiu16a, SWW^{+}16, XWL18, ZY17]. Biggeri [Ano10m]. Binary [DSM13, LMW19]. Bini [Ano10p]. Binomial [CMAC12, Qu12]. Bioconductor [Ano12o]. Bioinformatics [Ano10d]. Biological [FD15, ZZZ^{+}15, Ano10h]. Biology [Ano12p, Hor11a]. Biomedical [Che12b, SHL15]. Biostatistics [Boo10a]. Biosurveillance [SB10]. Birnbaum [ZLD16]. Bissett [Ke10]. Bit [ZHOY13]. Biweight [CRW18]. Black [MW19]. Black-Box [MW19]. Blackbox [CW16, CL16, GGL^{+}16a]. Blackjack [Ahm19b]. Blending [BDB15]. Blind [KMQ18]. Block [Bag10, Mic10]. Blocked [MGJ14, Wv11]. Blocking [JN16, SSG15]. Blurred [KQ14]. Board [Ano12b, Ano13j, Ano14h, Ano15e, Ano16i, Ano17h]. Boca [Mar19a]. Bocard [Mye12c]. Bock [Ano10j]. Boelen [Bar10]. Book [AA12, Ahm19a, Ahm10, Ahm19b, And11, Ano10s, Ano10g, Ano10j, Ano10r, Ano10h, Ano10n, Ano10q, Ano10p, Ano10t, Ano10k, Ano10o, Ano10d, Ano10e, Ano10f, Ano10i, Ano10m, Ano10l, Ano10u, Ano10v, Ano10w, Ano10x, Ano11d, Ano11o, Ano11p, Ano11h, Ano11n, Ano11f, Ano11g, Ano11l, Ano11e, Ano11j, Ano11m, Ano11i, Ano11c, Ano11k, Ano11q, Ano11r, Ano11s, Ano11t, Ano12h, Ano12n, Ano12f, Ano12j, Ano12e, Ano12i, Ano12c, Ano12g, Ano12k, Ano12d, Ano12q, Ano12p, Ano12o, Ano12l, Ano12r, Ano12m, Ano12s, Ano12t, Ano12u, Ano12v, Ano13c, Ano13d, Ano13e, Ano13f, Ano14b, Ano14c, Ano14d, Ano14e, Ano15d, Ano15a, Ano15b, Ano15c, Ano16d, Ano16e, Ano16f, Ano16g, Ano17f, Ano17c, Ano17d, Ano17e, Ano18a, Ano18b]. Book [Ano18c, Atk12, Bai11, Bai12, Ban10b, Ban10a, Ban11, Bar10, Bar11, Bha12, Boo10b, Boo10a, Bur10, BL11, Bur11, Bur12, Bzi11, Cha11, Cha12, Che11c, Che11a, Che11b, Che12a, Che12b, Cin11, Cob10, Esp11, Fot10b, Fot10a, Fre12, Gho10b, Gho10a, Gho11a, Gho11b, Gho12a, Gho12b, Gho19, Gol11b, Gol11a, Gol12c, Gol12b, Gol12a, Gru11, Guo19, Hec10, Hec11, Hey12, Hin11, Hly10a, Hly10b, Hly11, Hly12, Hor10b, Hor10d, Hor10c, Hor10a, Hor11f, Hor11d, Hor11b, Hor11c, Hor11a, Hor11e, Hor12c, Hor12b, Hor12d, Hor12a, Jar12b, Jar12a, Jen10, Jen11, Kat11, Ke10, Ke11a, Kuh10, Laz10, Laz11, Li10, Lip10a, Lip10b, Lip11a, Lip11c, Lip11b, Lip12b, Lip12d, Lip12c, Lip12a, Lip19a, Lip19b, Lip19c, Lip19d, Lip19f]. Book [Lip19e, Lu11, Mag11, Mai11, Mar19a, Mar19b, McC10, McC12b, McC12c, McC12a, Mic10, Mye11a, Mye11b, Mye11c, Mye12a, Mye12c, Mye12b, Mye12d, Nat12, Neu10b, Neu10a, Neu11, Neu12c, Neu12b, Neu12a, Ng10, Ng11, Ng12, Nku10, Oli10a, Oli10c, Oli10b, Oli11b, Oli11a, Oli12c, Oli12a, Oli12b, Ors19, Par11, Pen10, Pfa12, Qia10b, Qia10a, Qiu12, Qu11, Qu12, Ros11, Rut12, San11, Sar11, Sau10, Sau11, Sau12, Sea11a, Sea11b, Sea12, Sen11, SB19, Sym11, Van10, Wie11, WMI11, Wlu12, Ye10, You10, Yu12]. Bootstrap [AGMRS11, LKS16]. Both [DLY^{+}17, HLLY16]. Boundary [Tan18]. Bounded [DSD12]. Bounds [Lu11]. Box [MW19, Hun13, ZY17]. Bradley [Atk12, Che11a]. Brajendra [Gol12c]. Brenda [Sar11]. Bretz [Che11b]. Brian [Ke10, Pfa12]. Bridge [JSMS15]. Brockhoff [Hor11e]. Broemeling [Boo10a]. Brooks [Che12a]. Brownian [PA17, TQW13]. Bruce [Laz11]. Bruggemann [Lip12b]. Bug [SDR15]. Bug-Database [SDR15]. Building [PT14, SHD18]. Burgess [Hor12c]. Burkholder [Hor12c]. Burn [YXTS12]. Burn-In [YXTS12]. Business [Mye12a, Neu10b, Wlu12, McC12a]. Byleen [Mye12a].

C [Ano10s, Ano10t, Ano11l, Ano12e, Ban10a, Gol12c, Wie11]. Calibration [CBBJ18, CJ14, EPD18, GBH^{+}13, HGL^{+}13, JY15, OY17, PSB^{+}13, PH16]. CALL [Ano13h]. Campaign [ZCW19]. Capture [KKS^{+}14]. Capturing [SHD18]. Carbon [KKS^{+}14, SRZ^{+}19]. Care [Ban10a]. Carlin [Che11a]. Carlo [Ano10n, Ano11e, Mar19b, Neu11, Ng10, SH14, Che12a]. Carlos [WMI11]. Carol [Ano10e]. Case [Atk12, Cob10, LD11, LCAc13, PFF14, RMZ15]. Case-Deletion [PFF14]. Casella [Neu11, Ye10]. Casino [Ahm19b]. Castillo [WMI11]. Castillo-Chavez [WMI11]. Castle [Hor10d]. Catalyst [Fre13]. Categorical [Gho19, PGS11]. Catherine [Ano10e]. Causality [Ano10i]. Causation [Gol11a]. Cell [Zha11]. Cells [RV18]. Cellwise [HRV19]. Censored [CBBJ18, RZ16, ZYH14, WYJ10]. Censoring [EW10, GLLS14, KMM11]. Central [MM19, SGH13]. Centric [Lip19b]. certain [WYJ10]. Chain [Che12a]. Chains [AHS16, CMR10, Hly10b]. Challenges [Qiu16a, SB10]. Cham [Ahm19b]. Chambers [Oli10c]. Chan [Qia10b]. Chang [Hor10c]. Change [AG12, HN12, Lip19e, PZQ16, WB18]. Change-Point [HN12, PZQ16]. Changepoint [BFE^{+}19]. Changepoints [AHS16, NAEK15]. Changes [NAEK15, RTA11]. Chaoqun [Par11]. Characteristic [Gho10a]. Characteristics [XTP17]. Characterizing [SFH^{+}13]. Charles [Nat12]. Chart [BJFB14, CM11, CZZ16, JFEH14, SD12, ZT11]. Charts [Apl12, QWKR14]. Chavez [WMI11]. Chemometrics [Lip10b]. Chen [Ano10h, Ano11f]. Chin [Gol12a]. Choice [WND^{+}18]. Chow [Gol12b]. Chowell [WMI11]. Christian [Neu11]. Christiane [Ng10]. Christman [Hor11f]. Christopher [Gho12b]. Christos [Ano10d]. Chung [Gol12b]. Civil [HLSR15]. Cizek [Jar12b]. Claims [LCL12]. Clarification [Ano18e]. Clarifications [HS17]. Clark [Cob10, Ano12h]. Class [DNV16, TL16]. Classes [DNV16]. Classification [Ano10j, Ano11u, GJL15, KRSA15, WZ19, Ano11e]. Classifications [dMvW10]. Classifiers [QL19]. Claus [Ano11u, Neu12c]. Clear [Grö14, WMT12]. Cleophas [Ano10r, Ano10r]. Clinical [Ano10r, Ano11g, Che11a, Gol12b, Gol12a, And11]. Cloud [YPP19]. Cluster [WSZ14]. Clustering [cCOP^{+}13, CS12, DF12, Far14, HCT17, JSQ12, JS12a, JS12b, KMQ18, KRSA15, LW12, Par11, TMAZ18, WSZ12, ZZD^{+}14, Hor11d]. Coastal [AGRI19]. Codes [LC15, RBM08, XQW13, RBM11]. Coefficient [GAZ15, KL17]. Cointegrated [Ano10g]. Cokriging [LC15]. Cokriging-Based [LC15]. Cold [RKE10]. Collaborators [Ano10y, Ano11v, Ano12w, Ano13g, Ano14g, Ano16h, Ano17g, Ano18d, Ano19b]. Collection [ZTEM17]. Colorimetric [ZS15]. Column [LL16, Qu10]. Combination [HS14]. Comment [ACH14, Apl10, Bor12, Bre14a, CW16, CL16, CMS10, CS12, DM12, Dav13, DF12, For13, Fre13, GJ10, Gra13, HWL16, JSQ12, Jon13, Jos12b, Kle13, Lee12, LW12, LZ14, LS13, O’C10, OW12, PGK16, Pie13, PJW13, Ran13, SWWR14, SF10, Sin10, SJ12, Tsu10, TQW13, WSZ12, WAC10, WBD10]. Common [ZY18]. Communication [PGS11, Cha11]. Community [GWN13]. Companion [Sea11a]. Comparative [Ano11j, Lip12a, Neu10a]. Comparing [HG11, KSB16, SAC17, SFH^{+}13, Gho11a]. Comparison [Bor12, CMAC12, Gru11, LM13]. Comparisons [Che11b, Gol11a]. Competing [SS15, SNM10, YXTS12]. Complete [Lip19d]. Complex [HNKF18, JDTW15, Kat11, LAMO^{+}16, Lip11a, Lip19f, PH16, RBM08, RBM11, Ano12f]. Complex-Valued [HNKF18]. Compliance [LRL^{+}12b, LRL^{+}12a]. Component [CFF13, MGJ14, WMP18, ZTEM17, ZW17]. Composite [GLLS14, MM19, Nat12]. Comprehensive [Van10]. Compressing [CG16]. Compromise [Jon13]. Computation [Got10b, Got10a, HN18, Jos12a, Lip19b, NHQ17, Sea11b, You10]. Computational [Che11c, Neu12b, Oli10a, Boo10b, Hin11, Lip10a]. Computationally [Ano12f, CMAC12, FNL11, RHK11]. Computer [BGMA14, DLLR17, DP10, Dri10, EPD18, FK14, GBH^{+}13, GL12, LC15, HS14, HTW17, HGL^{+}13, HLLY16, HJM15, JLLM^{+}16, JDLH16, JK11, KAT18, LZ16, MDS12, Mor12, Mor15, OY17, PGRC13a, PSB^{+}13, PH16, RBM08, RHK11, Ran13, She17, Tan15a, Tan15b, Tan17, TWY14, VSM13, XQW13, ZQZ11, RBM11]. Computing [HS16, HS17]. Concatenation [VGS19]. Concepts [Jen10, Lip19b]. Conceptual [Oli10a]. Concerns [LS13]. Concise [Neu12a]. Condition [KZS^{+}18]. Conditional [CM14]. Conditions [Tan18]. Conference [Ano11u, CCCC13]. Confidence [Ruk15, WLHB15]. Conjugate [SL18]. Connected [Lip11b]. Connection [MZ13]. Connectivity [ZCG^{+}16]. Consistent [FP17]. Constable [Rut12]. Constrained [CW16, CL16, Far14, GGL^{+}16a]. Constraints [LZ14]. Constructing [MKBN19, SGS15, VGS19]. Construction [EJLN17, VX19]. Consumer [Hor11e]. Contamination [PD11]. Continuous [Ahm19c, CMR10, JFJ^{+}19, LMW19, Qia10a, Sar11, SH14]. Contour [RBM08, RBM11]. Control [Ahm19d, BJFB14, CM11, CZZ16, Hor10c, Jen11, JFEH14, LQ14, QL11, QWKR14, SD12, TS12, XTP17, ZJT11, ZT11]. Controller [TML16, TC17]. Controlling [WH19]. Controversial [Gol12b]. Convex [HS14, QL19]. Convolution [KZS^{+}18]. Cook [GAZ15]. Cooper [Ano10s]. Cooperation [Ano10j]. Coordinate [OW17, SAH^{+}17, SdCR12]. Coordinate-Descent-Based [SAH^{+}17]. Copula [XHX17, ZW17]. Core [Qu11]. Correction [Ano19c]. Correlated [HWL16, JP13, NS15, Wv11, Nku10]. Correlation [Pie13]. Corrigendum [Ano13i, Ano18f]. Count [STZW16]. Counterfactual [Lip19e]. Counting [XMM16]. Course [Ano11c, Hey12, San11, McC12b, Qu11]. Covariance [FOU13, HS16, HS17, LZ16, Zha11]. Covariate [HM13]. Covariates [HDM^{+}15, HW10]. Cox [ZY17]. CRC [Mar19a]. Crimes [SNY^{+}10]. Criminology [Ano12r]. Criteria [LRL^{+}12b, LRL^{+}12a, LACR11, Ran13]. Criterion [CRW18, CHS11, HN12, HTW17, SdCR12, TG10]. Cross [LC15, ZZD^{+}14]. Cross-Sectional [ZZD^{+}14]. Cross-Validation [LC15]. Crowds [Lip11b]. Cryer [Qia10b]. Curt [Ano11g]. Curve [QZZ18, ZDH14]. Curves [CS12, DF12, Gho10a, HSL^{+}13, JSQ12, JS12a, JS12b, LW12, Qia10a, SFH^{+}13, WSZ12]. CUSUM [CZ15a]. Cusums [LCL12]. Cyber [XHX17]. Cylinder [GGP10].

D [Ano11g, Ano12c, Bai12, Boo10a, Che12b, Fre12, Hey12, Mye12b, Nat12, Qia10b, CG16, MQ11, ZQ18, ZZD^{+}14]. Dale [Ban11, Gho12a]. Damage [YZW19]. Daniel [Ano11i, Mye12c]. Daphne [Bha12]. Darren [Ano12p]. DasGupta [Ano11m, McC12b]. Data [Ano10j, Ano10h, Ano10d, Ano11o, Ano11u, Ano12f, Ano12j, Ano12g, Ano12o, Bai12, BFE^{+}19, BJFB14, BDB15, Bur10, CBBJ18, CM11, CM17, CCCC13, CMAC12, cCOP^{+}13, CY17, CRBM19, DGS13, Fre13, GJL15, Ger15, Gho19, GG12, Guo19, Hec11, HM13, HDM^{+}15, Hor10b, Hor11a, HRSV16, HW10, ILS^{+}17, JAP15, JDLH16, JFEH14, JP13, Ke10, Ke11a, Kuh10, LG13, LKS16, LSP13, Lip12a, Lip19b, LCAc13, Mar11, Mar19a, MGSC18, MLBM19, MBR14, NKCB14, Nku10, Oli10c, Oli12a, Par14, PHK^{+}18, Pen15, Qiu16a, RHK11, RMVH11, RTA11, RV18, SHD18, Sar11, SWW^{+}16, STZW16, SAR16, SZ19, Sym11, WX10, WYHT18, Wył10, XWL18, XKHY18, XHM^{+}17, YPS18, YPP19, YT16, ZQ18, ZLD16, ZY17, ZS15, ZYH14, ZCM^{+}18, ZWZJ15, Ano10o, Che12b, Fre12, Gho11b, Gho12a, Gol12c, Kat11, Mai11, McC10, Qia10a, Ano11e]. Data [Par11, Sau10, Sau12]. Data-Centric [Lip19b]. Data-driven [Wył10]. Data-Focused [CCCC13]. Database [SDR15]. Dataset [BFK^{+}12]. Datasets [GB18, KH19, NKCB14, ZHOY13]. David [Ano10t, Ano11g, Ano12e, Ano12q, Ano12r, Bai12, Hor10d, Lip11b, Pen10, Qia10a]. Davis [Hor12c, Hor12d]. Debabrata [Ano11m]. Deblurring [KMQ18]. Deborah [Ano10k]. Decision [Ano11f, Lip19f]. Decision-Making [Lip19f]. Decisions [Guo19]. Decomposition [HS19, LAMO^{+}16, YPS17, YPS18, Oli12c]. Deepayan [Kuh10]. Defect [Sen11, SDR15]. Defense [XHX17]. Definitive [JN16, JN17]. Degenerate [ZZZ^{+}15]. Degradation [Ano10e, HDM^{+}15, Pen15, SYW19, TL16, WX10, WMEW13, VWGR11, XKHY18, YXTS12, YC14, ZY18, ZSGM14]. Degradation-Based [YXTS12]. Degrees [CHKC10]. Deletion [PFF14, RMZ15]. DeMets [Ano11g]. Dempster [Mye11a]. Denoising [MQ11]. Densities [TWY14]. Density [CGR^{+}15, MGSC18]. Department [CCCC13]. Dependence [FP17, ZW17]. Dependent [GGP10, HRC10, LD11, Liu12]. Depth [HS19, TMAZ18]. Derivative [CHS11, HSL^{+}13]. Derived [TW12]. Descent [SAH^{+}17]. Described [CMR10]. Design [AA12, ANAC13, Atk12, BN12, BHGL19, DMW10, DP10, DASS17, EPD18, GJ19, LC15, JLLM^{+}16, JN17, Jos12a, KJB11, KAT18, KKS^{+}14, LHTD18, LZ15, MSE17, Mor15, MGJ14, OW17, PGS11, Pie13, PJW13, PHBF17, RBM08, SD15, Tad13, TW12, TW13, Tan13, Tan15a, XQW13, XDHQ16, ZDH14, RBM11, Ye10, Gol11b, Neu10a, Wie11]. Design-Based [JN17, XDHQ16]. Designed [LACR11]. Designs [ASDMLF16, ACH14, AG12, BMB15, CMAC12, CWW17, DSD12, DASS17, Edw11, ES17, EJLN17, GG12, Got10b, Got10a, Grö14, He19, HN18, HQH16, JN11, Jon13, JSMS15, JN16, JN17, JDTW15, JWG^{+}19, JGBM19, LL16, LMZ19, LMW19, MM19, ME12, Oli10a, PCWW16, Qu10, RB10, SSG15, SGS15, SWWR14, She17, SdCR12, Tan13, TG15, TG17, TG10, VX19, VGS19, Wv11, WMT12, Mic10]. Destination [Car14, Haz10]. Destructive [XKHY18]. Detecting [RMZ15, RV18, ZZD^{+}14]. Detection [AHS16, BFE^{+}19, BVW10, BH13, FQ18, Fra19, GWN13, Han10, JFJ^{+}19, KQ14, KRSA15, LMT14, NHB^{+}13, SB10, Ste10b, TMMH17, WB18, YPS17, YZW12]. Deteriorating [YXTC14]. Deterministic [HS14, JWG^{+}19, RHK11, Tan17]. Development [And11, Mor15]. Deviating [RV18]. Deviation [SHD18, SD12]. Dey [Ano11f]. Diagnosis [ILS^{+}17, PZQ16]. Diagnostic [Boo10a, BA18, ZJT11]. Diagnostics [PFF14]. Dieter [AA12]. Difference [FQ18]. Differences [HCT17, HP12]. Different [MM19, TWY14]. Differential [LZW15, SCH17]. Diggle [Bur12]. Digital [Lip19c, Sen11]. Dimension [CSY18, MZ15, WX14, ZM19]. Dimensional [ANAC13, Ano12n, CGR^{+}15, Dav13, DNV16, Fre13, GFK^{+}18, HRSV16, LS13, LMS15, Mar11, PTZ19, Pie13, PH16, SL18, WSZ14, YPS18, ZHOY13, ZLD16, ZWZJ15]. Dimensions [QLS^{+}19]. Dimitris [Hor12d]. Dipak [Ano11f]. Directional [SWW^{+}18]. Dirichlet [HRC10, WB18]. Dirk [Ano10n]. Discovering [SAR16]. Discrete [Ano12c, GJ10, HEM10a, HEM10b, Kle13, LMW19, MW19, O’C10, SF10, Sin10, WAC10]. Discrete-Event [Kle13]. Discrete-Use [GJ10, HEM10a, HEM10b, O’C10, SF10, Sin10, WAC10]. Discriminant [CHWE11, MZ13, ZM19, ZS15]. Discrimination [WZ19]. Discussion [Han10]. Disease [Hor11d, Li10]. Disparate [SHD18]. Dispersion [HL17, TL16]. Distance [CYZ17, GJL15, GAZ15, HS16, HS17, JK11, WZ19]. Distributed [AS16]. Distribution [AMY14, Apl12, BJFB14, CM17, CZZ16, CYZ17, HRC10, Oli12b, WW18, Zha10, ZTEM17]. Distribution-Free [BJFB14, CM17, CZZ16]. Distributions [BJFB14, CM14, DNV16, FNL11, KRSW19, WYJ10, Gho11a, Pen10]. Dmytro [Mye11c]. Do [DM12, Jos12b, Lee12, OW12, SJ12]. Dobson [Hor11a]. Doganaksoy [Neu10b]. DOI [Ahm19b]. DoIt [Jos12b, Lee12, OW12, Jos13, SJ12]. Dolt [DM12]. Domenico [Ano10p]. Dominating [Grö14]. Donald [Hor12c]. Dongchu [Ano11f]. Down [XTP17]. Downscaling [MKBN19]. Dräghici [Ano12o]. Draief [Lip11a]. Driven [JY15, Wył10]. Drton [Hor11b]. Dudley [Ano11n]. Dudoit [Ano12l]. Durrett [Laz10]. Dynamic [Ahm19d, AGMRS11, CJL14, Ger15, HM13, HDM^{+}15, LQ14, PTZ19, QX14, QZZ18, RKE10, TML16, TC17, YR11, ZY18, Gol12c]. Dziga [Lip19c].

e-book [Ahm19b, Mar19a]. Early [SB10, XHX17]. Early-Warning [XHX17]. Easley [Lip11b]. Ecology [Hor12a, Mye12c, Oli10b]. Econometric [Ban11]. Econometrics [Hor10d]. Economic [Lip19f, McC10]. Economics [Mye12a]. Economy [Nat12]. Ecosystems [XMM16]. ed [Ahm19b, Ano10s, Ano10g, Ano10r, Ano10n, Ano10o, Ano10i, Ano10l, Ano11d, Ano11p, Ano11h, Ano11g, Ano11j, Ano11i, Ano11c, Ano11k, Ano12h, Ano12n, Ano12e, Ano12d, Ano12q, Guo19, Hly11, Jar12b, Mar19b, Oli11a, Ors19, Pfa12, Qia10b, Qu12, Rut12, San11, Sau10, Sau11, Sau12, Ano13j, Ano14h]. Edition [Ano12g, Ano12i, Ano12c, Ano12p, Ano12o]. Editor [Qiu15, ANAC14, Chi11, Chi12b, Chi13, Con10, Apl17, E19, Qiu14, Qiu16b, Ste10a]. Editorial [Ano10y, Ano11v, Ano12b, Ano12w, Ano13g, Ano14g, Ano15e, Ano16i, Ano16h, Ano17h, Ano17g, Ano18d, Ano19d, Ano19b, CJDN15, Chi12a, Edi15]. Educational [Ano10p]. Edward [Mye11b]. Effect [Bag10]. Effective [JN17]. Effectiveness [XHX17]. Effects [Apl10, CMS10, DLY^{+}17, ES17, HL17, Hof10, Kat11, LRL^{+}12b, LRL^{+}12a, MGJ14, Oli10b, Pen15, QZW10a, QZW10b, SGC14, Tsu10, WBD10]. Efficiency [Ban10a]. Efficient [BGMA14, CMAC12, CW16, CSY18, EJLN17, FNL11, JN11, KMQ18, SH14, ZM19, ZWZJ15]. Eigenvalues [Lu11]. Ekstrøm [Neu12c]. Elastic [SAH^{+}17, WSZ14, ZLD16]. Elastic-Net [ZLD16]. Electricity [AGMRS11, CJL14]. Electromagnetic [JLLM^{+}16]. Electron [CGR^{+}15]. Elementary [Laz10]. Elements [Oli12b]. Elena [Bai11]. Elliptical [BJFB14]. Emissivity [BWMG14]. Emmanuel [Ano12d]. Empirical [PTZ19]. Emulation [CBD^{+}17, OY17, WHQ18, YR11, ZQZ11]. Emulations [KAT18]. Emulators [FOU13, GL12, PT14]. Enabled [CM14]. Encyclopedia [Ahm10]. Energy [CCCC13, JDTW15, JWG^{+}19]. Engine [GGP10]. Engineering [ANAC13, Dav13, JY15, Pie13, Gho10b]. Engineering-Driven [JY15]. Engineers [Bzi11, Cin11]. Enhance [JFEH14]. Enhancements [PGK16]. Enhancing [Jon13]. Enough [Bzi11]. Ensemble [CG16, HGL^{+}13]. Enterprise [Rut12]. Entry [Far14]. Entry-Wise [Far14]. Envelope [ZL17]. Envelopes [CZ15b, CS16]. Environmental [Nat12]. Environments [ZY18]. EOV [Ano13j, Ano14h, Ano15e, Ano16i, Ano17h]. EPIC [LD11]. Epidemics [Lip11a]. Epidemiology [Li10, WMI11]. EQIE [HTW17]. Equations [LZW15, SCH17]. Equivalence [Ano11p, DNV16, MZ13]. Era [ILS^{+}17]. Eric [Fre12, Laz11]. Erik [Bai11]. Errata [RBM11]. Erratum [Ano12x, Ano16j, LRL^{+}12a]. Error [LG13, Ano10k]. Errors [ZCW19]. Errors-in-Variables [ZCW19]. Estimate [CSY18, ES17]. Estimating [GW13, LM15, Par14, SNM10, TBMM15, ZTEM17]. Estimation [Ahm19a, CGR^{+}15, CHS11, CY17, CYZ17, JWFW19, LVB13, LZW15, LZHZ18, LCAc13, MGJ14, PBH13, RBM08, RZ16, SCH17, SGH13, SWF16, WC12, WMI11, YXTC14, Zha10, ZDH14, ZYH14, RBM11, Gru11, Ahm19a]. Estimators [AMY14, Ano11j, Wad16]. Ethem [Ano10l]. Eugene [Ano10r, Ano11l]. Eva [Ano10f, Fot10b]. Evaluating [dMvW10]. Evaluation [Ano10p, Bar10, CBC15, Guo19, LMT14, TMMH17]. Evarist [Ano11n]. Event [GLLS14, HW10, Kle13, LÇC12, QWKR14]. Events [RMVH11, WYHT18]. Everitt [Pfa12, Yu12]. Evolution [Ano12d]. Evolutionary [Cha12]. EWMA [TML16, ZT11]. Ewout [And11]. Exact [HN18, PSR17, WLHB15]. Example [Hly10b, Ke11b]. Excel [Ke10]. Exchange [OW17, SdCR12]. Exchanges [Ano10k]. Excursion [AGRI19, CGB^{+}14]. Expanded [XC10]. Expansions [Gho12b]. Expected [CW16, For13, Gra13, WC12]. Expectile [YZZ18]. Expensive [FNL11, JWG^{+}19]. Experiment [GLLS14, Mic10, RBM08, SD15, RBM11]. Experimental [AA12, ANAC13, ASDMLF16, Ano10k, CMAC12, GG12, GJ19, MM19, Mor15, Pie13, PJW13]. Experimentation [Fre13, Van10]. Experiments [AG12, Atk12, BN12, BHGL19, BGMA14, BDB15, DP10, EDR16, GL12, HS14, HTW17, HLLY16, HJM15, JLLM^{+}16, JK11, Jos12a, JGBM19, KJB11, KAT18, LWM13, LACR11, MKBN19, MDS12, MGJ14, OW17, PGS11, PGRC13a, PHBF17, Qu10, She17, TW13, Tan15a, Tan15b, Tan17, TWY14, WB11, Wv11, dCC11, Neu10a]. Experiments-Based [Jos12a]. Explanatory [Pen15]. Exploiting [Wad16]. Exploration [BHGL19, cCOP^{+}13, JDTW15]. Exploratory [Ano12g, Mar19a]. Exploring [Apl12, CCCC13]. Exponential [CJL14, TL16, ZDH14].

Exponential-Dispersion [TL16]. Exponential-Linear [ZDH14]. Exposure [JLLM^{+}16]. Extensions [CZ15a, Oli10b]. Extrema [AGRI19]. Extreme [Wad16, Jar12a].

F [Ano10r, Bai11, Hly12, Hor10d, Oli10b]. Fabrizio [Ahm10]. Facing [SB10]. Factor [AGMRS11, AG12, Bag10, CJL14, ES17, Grö14, LMZ19, Tad13, WMT12, ZX17]. Factorial [DASS17, EDR16, Fra19, Grö14, LMZ19, RB10, SGS15, TG10, Van10, WMT12]. Factorial-Based [DASS17]. Factorials [HL17]. Factorization [DS17]. Factors [DLLR17, GJ19, LMW19, ZQZ11]. Fahrmeir [Ano10q]. Failed [WYHT18]. Failed-But-Not-Reported [WYHT18]. Failure [HM10, HM13, Liu12, SYW19, SNM10]. Failure-Time [HM13]. Failures [YT16]. Faltin [Ahm10]. Familial [Gol12c]. Families [HLP13]. Farm [PHD16]. Fast [CGB^{+}14, EJLN17, Got10b, Got10a, LC15, HN18, HS16, HS17, PSB^{+}13]. Fault [ILS^{+}17]. Feature [Lip10a, PJR15]. Federer [Mic10]. Fernandez [Ano11o]. Ferrari [Ano10m]. Ferraty [Ano12j]. Ferreira [Ban10b]. Festschrift [Ano10q, Hor10d]. Fetus [JLLM^{+}16]. Fibrous [GLLS14]. Fidelity [LC15, HTW17]. Fieguth [Lip12c]. Field [CY17, HM10, HM13, MLBM19, WYHT18, YT16]. Field-Failure [HM10, HM13]. Fields [JLLM^{+}16]. Fienberg [Ano10t]. Fighting [SNY^{+}10]. Filling [DSD12, JGBM19]. Film [Lip19c]. Filter [HGL^{+}13]. Filtering [Ahm19a, SPS18]. Filzmoser [Lip10b]. Finance [Ano10e, Hor10a, Jar12b, Jar12a, Ors19]. Financial [Gho10b, Mye11b, Mye11c, SNY^{+}10]. Finding [LMW19]. Fingerprint [LD11]. Finite [AHS16, HCT17, Hec10, LX18, Tan13, Mye12a]. Finite-State [AHS16]. Finkelstein [Jen10]. First [Ano11c, Hey12, McC12b]. Fit [PSR17, ZX17]. Fixed [MGJ14, TH18]. FL [Mar19a]. Flaw [TMMH17]. Flaws [LMT14]. Flexible [Gol12a, YZZ18]. Flooding [AGRI19]. Flow [ASDMLF16]. Flowgraph [HW10]. Focused [CCCC13]. Foldover [EJLN17, LL16]. Follow [KAT18]. Follow-Up [KAT18]. Forecast [PHD16]. Forecasting [AGMRS11, HLSR15]. Forensic [LM13]. Forest [Mye12b]. Forgings [LMT14]. Form [dCCT15]. Foundations [SB19]. Fractional [Grö14, LMZ19, RB10, SGS15, WMT12]. Fractionated [TW13]. Frailty [Bai12, Liu12]. Framework [DASS17, TMMH17, XMM16, ZJT11]. Franceschetti [Cha11]. Francesco [Ano11e, Cha12]. Francisco [Gru11]. Francois [Mye12c]. Frank [Che11b]. Fraser [Qiu12]. Fraud [BVW10, Han10]. Frederic [Ano12j]. Frederick [Ahm10, Ano10t]. Free [BJFB14, CM17, CZZ16, ML16, dCCT15]. Free-Form [dCCT15]. Freedom [CHKC10, Mic10]. Frees [Mye11b]. Frequentist [Gru11]. Friedman [Ano11g, Bha12]. Frontier [LACR11]. Frontiers [Ano11f, Jen11]. Fronts [LCAc13]. Fuentes [Bur12]. Full [DASS17]. Function [HLP13, LZ16, SWF16, TBMM15, WLHB15]. Function-Based [HLP13]. Functional [Ano12j, CM14, DGS13, Dri10, EPD18, GJL15, GAZ15, Ger15, HS19, HJM15, Mai11, MGSC18, SZ19, SGC14, Tan18, YPS18, YZW12, ZCG^{+}16, ZSGM14, ZCM^{+}18, Bur10]. Functions [CBD^{+}17, JWFW19, MGSC18, WHQ18]. Fund [Mye11a]. Fundamentals [Cin11, Ano11g, McC12b, Oli11a]. Furberg [Ano11g]. Fused [SWF16]. Fusion [HLSR15, NKCB14]. Future [ML16].

G [Ano10k, Ano12q, Ano16k, Ano17i, Ano18g, Bai11, Bre13, Bre14b, Bur10, BL11, Che12b, Gho12b, Hla15, Hor11a, Oli10a, Oli10b, Sar11]. Galin [Che12a]. Gamma [AMY14, Liu12, WW18, YXTC14]. Gamst [Oli10a]. Gan [Par11]. Ganapati [Lip12b, Nat12]. Garrote [PJR15, Xio10]. Gary [Ban11]. Gatignon [Ano10o]. Gauge [BMS10]. Gaul [Ano10j]. Gaussian [Fot10a, DLLR17, FOU13, GL12, GH16, Gui18, HS14, HCT17, HSL^{+}13, KKS^{+}14, KZS^{+}18, LX18, LZHZ18, Mor12, Pen15, PSR17, RHK11, SCH17, TW12, Tan18, TL19, WX10, WB18, YC14, dCCT15]. Gaussian-Process-Based [SCH17]. Gautam [Mye11a]. Gebhardt [AA12]. Geer [Ano12m]. Geert [Gho11b]. Gelfand [Bur12]. Gelman [Che12a]. General [BDB15, GG12, SGS15, TG10]. Generalized [AMY14, CHS11, CYZ17, DP10, KRSW19, Oli12c, TC17, Zha10]. Gentle [Hin11]. Geodesic [dCCT15]. Geometric [Ano10f, DS17]. Geometry [Hor10a]. Georg [Mye11a]. George [Ano11o, Neu11, Ye10, Hun13]. Geostatistics [Cob10, LZ15]. Gerald [Neu10b]. Gerardo [WMI11]. Gerhard [Ano10q]. Gesellschaft [Ano11u]. Gibbs [Laz11]. Gibilisco [Ano10f]. Gillet [Mye12c]. Gine [Ano11n]. Given [WLHB15]. Glaz [Nag11]. Glenn [Hor10b]. Global [CG16, LWM13, MKBN19]. GOEM [NHQ17]. Gold [DSM13]. Gonena [Gho10a]. Good [Oli12a]. Goodness [ZX17]. Goos [Atk12]. Gore [Nat12]. GPS [CM14]. GPS-Enabled [CM14]. Grading [WLWZ16]. Graham [Sau12]. Gramacy [CL16]. Graphical [Bha12, LX18]. Graphically [Apl12]. Graphics [Hec11]. Graves [Bur10]. Greedy [QLS^{+}19]. Greenacre [Ano11e]. Gregory [Lip12a]. Grids [LVB13]. Group [MDS12, QLS^{+}19, SAH^{+}17]. Grouping [Gui18, WSZ14]. Growth [GJ10, HEM10a, HEM10b, O’C10, Par14, SF10, Sin10, WAC10, WHQ18, ZDH14]. Gruber [Ano11j]. Guarino [Oli10a]. Guide [Ano12c, Bai11, Bzi11, Hor10b, Neu12a, Oli12a, Van10, Sar11]. Guidelines [ACH14, SWWR14]. Guide(R) [Rut12]. Guido [Ano10m]. Gunther [Boo10b]. Guojun [Par11]. Gusak [Mye11c]. Gut [San11]. Guttorp [Bur12].

H [Ano10r, Ano10d, Ano11j, Bai11, Ban10b, Hor12a, Jen11, Mye11a, Oli12c, Syl12]. Hahn [Neu10b]. Half [EG12]. Half-Half [EG12]. Hall [Ano11l]. Halloran [Wie11]. Hamann [Mye12b]. Hanagal [Bai12]. Hand [Qia10a]. Handbook [Ano10s, Ano12r, Pfa12, Bur12, Che12a]. Hans [Ano10j, Neu12a]. Hans-Hermann [Ano10j]. Hans-Michael [Neu12a]. Hard [AG12]. Hard-to-Change [AG12]. Hardback [Mar19a]. Hardle [Jar12b]. Hardware [CMR10]. Harper [Cob10]. Harris [Ano10s]. Harry [Chi12c]. Hastings [LKS16]. Health [Ban10a, Hor11a]. Hedges [Ano10s].

Hendry [Hor10d]. Henry [Ano10f, Hor12a, Hor10a]. Henry-Labordère [Hor10a]. Herbert [Ban10b, Bar10, Gol11a]. Hereditary [Hor10c]. Heredity [LZ14]. Hermann [Ano10j, Ano11u]. Heterogeneous [GW13, MLBM19]. Hettmansperger [Ano11k]. Heyde [Ano11l]. Hidden [Fra19]. Hierarchical [CHKC10, DGS13, HCT17, KL17, Li10, MLBM19, MMEDG14, PJR15]. High [AS16, CSY18, GFK^{+}18, HRSV16, LSP13, LMS15, MKBN19, Mar11, PTZ19, PH16, QLS^{+}19, Qu10, VWGR11, WSZ14, XQW13, YPS18, ZHOY13, ZLD16, ZWZJ15]. High-Accuracy [XQW13]. High-Dimensional [GFK^{+}18, HRSV16, LMS15, Mar11, PTZ19, PH16, WSZ14, YPS18, ZHOY13, ZLD16, ZWZJ15]. High-Performance [AS16]. High-Reliability [VWGR11]. High-Resolution [MKBN19]. High-Throughput [Qu10]. Highly [Lip11b]. Hilbe [Gol11c, Qu12, Ros11]. Hilbert [BLK19, YZZ18]. Hiroshi [Lip10a]. Hirukawa [Gho10b]. History [BVW10, WND^{+}18]. Hoaglin [Ano10t]. Hoff [Hey12]. Hoijtink [Bar10]. Honor [Ano11f]. Honour [Ano10q, Hor10d]. Hooker [Bur10]. Horton [Hec11]. Hothorn [Che11b, Pfa12, Yu12]. Hourly [CJL14]. Howard [Ke11b]. Hsiung [Hor10c]. HTS [LSP13]. Huan [Lip10a]. Huber [Ano10e, JWFW19]. Huber-Carol [Ano10e]. Hubert [Ano10o]. Hui [Ano11f]. Humanities [Lip19c]. Hunter [Ano16k, Ano17i, Ano18g, Bre13, Bre14b, Hla15, Cho11, Cho12, Mit10]. Hygiene [MBR14]. Hyman [WMI11]. Hypercube [BMB15, DP10, DASS17, HQH16]. Hyperspectral [BWMG14, SHL15]. Hypotheses [Ano11p, Bar10]. Hypothesis [HLP13, SZ19].

Identification [CGB^{+}14, HL17]. Identifying [QX14]. Ieno [Bai11, Oli10b]. IFCS [Ano11u]. II [Hor10b, WYJ10]. Image [FPG19, FQ18, KMQ18, Lip12c, MGSC18, MQ11, Par14, Qiu12, QX13]. Image-Based [FPG19]. Images [BWMG14, CGR^{+}15, FQ18, HP12, SHL15, SYW19, TEN14, TMMH17, YPS17, ZZD^{+}14]. Imaizumi [Ano10j]. IML [Sea11a, Sea12]. Imperfect [SB19, ZX17]. Implicit [AS16]. Importance [CBC15, LVB13, ZW17]. Improved [HN12, TMMH17, TG15]. Improvement [CW16, For13, Gra13, ZX17]. Improving [LCAc13, Zha10]. In-Plane [SHD18]. Inaccuracy [WH19]. Incorporating [HRC10, HW10]. Independence [SG18]. Index [Ano10z, Ano11w, GL12, WX14]. Indexes [Ano10m]. indicator [Lip12b]. Indirect [Esp11]. Individual [LMZ19]. Individuality [LD11]. Individuals [QX14]. Industrial [MBR14]. Industry [Ano10e, Neu10b]. Inference [AGMRS11, Ano10k, Ano10d, Bur11, Car14, DNV16, FK14, Fra19, Gho10b, GB18, Guo19, Haz10, KMM11, LAMO^{+}16, SWW^{+}16, SB19, SS15, SHG12, TG17, WYJ10, WW18, ZCW19, Ano10i, Sea11b]. Inference-Data [Guo19]. Influence [RMZ15, XMM16]. Info [SB19]. Info-Metrics [SB19]. Informatics [Ano13h]. Information [GW13, HN12, HM10, HM13, HDM^{+}15, LMZ19, SB19, Tan18, YT16]. Informative [Bar10, PJR15]. Infrared [BFK^{+}12]. Ingo [Hor11f]. Initial [Tan18]. Inputs [Mor12]. Inquiry [Lip19e]. Inspection [QWKR14]. Insurance [Jar12b]. Integer [SSG15]. Integrated [Ano10g, Bur11]. Integration [CRBM19, ZM19]. Intelligence [PCWW16]. Intensive [Ano12f]. Interactions [DWD^{+}14, ES17, Grö14, WMT12]. Interactive [QLS^{+}19]. Interdependence [CS12, DF12, JSQ12, JS12a, JS12b, LW12, WSZ12]. Intermediate [Ke11a, San11]. Internal [SRZ^{+}19]. International [Ahm19b]. Interpolation [Jos12a, RHK11, Tan15b]. Interpreting [Sar11, TS12, McC10]. Intertwining [Lip12d]. Interval [JAP15, WLHB15]. Interval-Valued [JAP15]. Intervals [Ruk15]. Introducing [Neu11]. Introduction [Boo10b, Hly10a, Lip12b, Yu12, Ano10l, Laz11, Lip10b, Ng11, Sym11]. Inverse [JK11, Oli12c, Pen15, Tad13, TBMM15, THS15, WX10, YC14]. Inversion [MMEDG14]. Inversions [CGR^{+}15]. Irene [Bar10, Cha12]. Irregular [QX14]. ISBN [Ahm19b, Mar19a]. ISBN-13 [Mar19a]. Ishwar [Ano11l]. Isobel [Cob10]. Isotonic [THT11]. Issue [Ano13h, CCCC13, Ste10b]. Issues [Gol12b]. Iterated [Bor12].

J [Ano10r, Ano11e, Ano11j, Ano12p, Ban11, Bur10, Che11a, Gru11, Hec11, Hly11, Hor10b, Hor11a, HS17, Jen11, Neu10b, Oli10a, Oli10c, Oli10b, Qia10a, Sea11a, Wie11]. Jack [Ano12n, Che11a]. Jackman [Mag11]. Jacobs [Ban10a]. Jake [Ano10h]. James [Ano11f, Che12b, Hin11, WMI11]. Jamis [Sea11a]. Jane [Wlu12]. Jank [McC12a]. Javier [Ano12k]. Jay [Fot10a]. Jeff [Mye12b]. Jeffery [Mar19a]. Jeffrey [Ano10s, Ano12g]. Jennifer [Hor10d]. Jensen [Fot10b]. Jiang [McC12c]. Jianhong [Par11]. Jiming [McC12c]. JMP [Sar11, Wlu12]. John [Ng12]. Johnson [Hor10b]. Joint [DJ15, HL17, LRL^{+}12b, LRL^{+}12a]. Jon [Lip11b]. Jonathan [Qia10b]. Jones [Che12a, Ng11, Atk12]. Jose [Sar11]. Josef [Wlu12]. Joseph [Ano11k, Gol11c, Nag11, Qu12, Ros11]. Jr. [Chi12c]. Judea [Ano10i]. Judgment [Gol11a]. Judith [Ano10t]. Jump [KQ14]. Jumps [TH10]. Junge [Ano11u]. Junichi [Gho10b]. Jurgen [AA12]. Just [Bzi11]. Justin [Ban11].

Kahle [Ano10e]. Kalman [HGL^{+}13]. Kaltenbach [Neu12a]. Karen [Che12b]. Karl [Mye12a, Qiu12]. Keener [Qu11]. Keh [Hor12d]. Keh-Shin [Hor12d]. Kelbert [Hly10b]. Kelly [Hly12]. Ken [Hec11]. Kenett [Ahm10]. Kenichiro [Gho10b]. Kennedy [Ke11a]. Kenneth [Ano11d, Ano11h]. Kernel [AS16, BLK19, JAP15, WZ19, YZZ18]. Keying [Ano11f]. Khuri [Oli11b]. King [Mic10]. Klein [Ano12q]. Kleinbaum [Ano12q]. Kleinberg [Lip11b]. Kleinman [Hec11]. Klugkist [Bar10]. Kneib [Ano10q]. Knowledge [Ano11u]. Koehn [Hor11c]. Kolaczyk [Fre12]. Koller [Bha12]. Koltchinskii [Ano11n]. Koop [Ban11]. Kragh [Lip11c]. Kriging [CGB^{+}14, GB18, PT14, PA17, PR17, TH18, WH19]. Kriging-Based [CGB^{+}14]. Krishnamoorthy [You10]. Kroese [Ano10n]. Krzanowski [Qia10a]. Kukush [Mye11c]. Kulik [Mye11c]. Kung [Qia10b]. Kung-Sik [Qia10b]. Kurt [Lip10b].

L [AA12, Ano11g, Ano12g, Ano12k, Ban11, Gho12a, Hor10d, Hor12c, Mar19a, Oli10a]. Labordère [Hor10a]. Lagged [TS16]. Lagrangian [CL16, GGL^{+}16a, PGK16]. Lang [Kat11]. Lange [Ano11d, Ano11h]. Laplace [Bor12]. Large [Ano12n, cCOP^{+}13, KH19, MLF^{+}16, NS15, NKCB14, NHQ17, PSB^{+}13, QL19, RB10, VX19, YR11, McC12c]. Large-Scale [NHQ17]. Large-Vector [NS15]. Larry [Ano10s]. Lasso [CRW18, SWF16, TS16, SHL15, ZJT11]. LASSO-Based [ZJT11]. Latent [Lip19a, XTP17]. Latin [DASS17, BMB15, DP10, HQH16]. Lattice [QAHS14, Kuh10]. Laurent [Lip11a]. Lauro [Ano11e]. Lavallée [Esp11]. Law [SS15, Jen10]. Lawrence [Ano11g]. Lawson [Li10]. Lawton [Syl12]. Lazar [Mai11]. Learned [BVW10]. Learning [Che12b, Fre13, Lip19d, The13, ZCG^{+}16, ZZZ^{+}15, Ano10l]. Least [Ahm19e, CM11, CS16, SPS18, XDHQ16, ZL17]. Least-Squares [CS16, ZL17]. Lectures [Gho19, Hor11b]. Lee [Che11a, Ban10b]. left [KMM11]. Legendre [Mye12c]. Lehmann [Ano12k]. Lemieux [Ng10]. Lene [Lip11c]. Lenz [Jen11]. Lessons [BVW10]. Letter [ANAC14, Con10]. Level [AG12, Edw11, ES17, EJLN17, Grö14, HL17, MSE17, PCWW16, SdCR12, Van10, VX19, VGS19, XHM^{+}17]. Leveraged [BMS10]. Li [Che12a]. Libraries [SWMW13]. Life [LHTD18, Liu12, Mye12a]. Lifetime [Ano10d, CY17, ZTEM17, WYJ10]. Lifted [PA17]. Lifting [HNKF18]. Lii [Hor12d]. Likelihood [Bur11, OY17, SCH17, SGC14, Wad16, XC10]. Limits [EW10, Hof10, ZPM14]. Limnios [Ano10e]. Lin [Che11c]. Linear [CZ15b, Lip11c, MZ13, PFF14, PSR17, SSG15, Sea11a, SZ19, SGC14, TML16, YR11, ZLD16, ZDH14, Gho11b, Oli11b]. Liners [GGP10]. Liu [Lip10a, Qiu12]. Local [Fot10a, GH16, KMQ18, MQ11, PD11, PR17]. Locally [NHB^{+}13]. Locarek [Ano11u]. Locarek-Junge [Ano11u]. Location [BJFB14, DLY^{+}17, EW10, HLP13, HL17, RTA11]. Location-Scale [HLP13]. Log [AMY14, EW10, ZLD16]. Log-Gamma [AMY14]. Log-Linear [ZLD16]. Log-Location-Scale [EW10]. Logical [Lip12a]. Logistic [ZPM14, Ros11]. lognormal [KMM11]. Lonardi [Ano10h]. Long [BWMG14]. Long-Wave [BWMG14]. Longini [Wie11]. Longitudinal [Gho11b, Gho12a, Gol12c, QX14]. Lorenzo [Lip12d]. Loss [TW12]. Low [JSMS15, XQW13]. Low-Accuracy [XQW13]. Ludwig [Ano10q]. Luigi [Ano10p, Ano10m]. Luis [WMI11]. Lumber [WLWZ16, YZW19]. Lyle [Boo10a].

M [Ano10t, Ano10e, Ano11n, Ano11g, CZ15a, Che11a, Gol11c, Hor12d, Mye11a, Oli10c, Oli10b, Qu12, Ros11, Wie11, WMI11, CZ15a]. M/M/1 [CZ15a]. Ma [Par11]. Machine [Ano10l, Hor11c, The13, WB18, Lip19d]. Machines [AS16, Hor11f]. MacroPCA [HRV19]. Magnani [Lip12d]. Mahalanobis [GJL15]. Maillardet [Ng11]. Main [Bag10, ES17].

Main-Effect [Bag10]. Maintainability [CMR10]. Maintenance [ZX17]. Making [Ano11f, Lip19f, Hor10b]. Maller [Ano11l]. Malley [Che12b]. Malware [KRSA15]. Management [Ano10o, GJ10, HEM10a, HEM10b, Hec11, O’C10, SF10, Sin10, WAC10, Mye11a, Wlu12]. Managers [McC12a]. Manifolds [CRBM19]. Manipulation [Sau10]. MANOVA [Zha11]. Mantovan [Ano12f]. Manufacturing [DJ15, SHD18, dCC11]. Mapping [Li10]. Marasinghe [Ke11a]. Marco [Ban10b]. Marcus [Fot10a]. Margin [QL19]. Marginal [MW19]. Maria [Ano10f]. Marine [GGP10]. Mark [Hly10b]. Marked [LD11]. Market [AGMRS11]. Markets [Lip11b]. Markov [Che12a, Hly10b, AHS16, CMR10, Fot10a]. Marks [LM13]. Markus [Neu12b]. Marten [Ano12m]. Martin [Hor12a]. Martinez [Ano12g, Mar19a]. Marvin [Ano11j]. Masanobu [Gho10b]. Massimo [Cha11]. Massive [BFK^{+}12, GB18]. Massoulie [Lip11a]. Matched [The13]. Matched-Pair [The13]. Matching [PSR17, WND^{+}18]. Material [SYW19]. Materials [DNV16]. Mathematical [WMI11]. Mathematica^{T M} [Lip12a]. Mathematics [Ano12c, Mye11c, Mye12a, Bar11]. Mathew [You10]. Mathias [Hor11b]. Matilde [Ano10p]. MATLAB [Ano12g, Bur10, Mar19a]. Matrices [Car14, Haz10, Oli12c, Zha11, Ano12n]. Matrix [DS17, Lu11, ZS15]. Maximin [SdCR12]. Maximum [SCH17, Wad16, WLHB15]. May [Ano12h]. Mayo [Ano10k]. MCEM [XC10]. McKean [Ano11k]. MCMC [FNL11]. Mean [Apl12, TS12, TH10, KMM11]. Means [MW19, Wlu12]. Measure [WC12]. Measurement [DSM13, DMW10, LG13]. Measurements [CRCH18, PGS11, PHD16]. Measures [GAZ15, WMEW13]. Measuring [Tad13, XMM16, Ban10a]. Medhi [Hly11]. Medical [Hec10]. Medicine [Lip19d, Boo10a]. Mee [Van10]. Meester [Cha11]. Meesters [Bai11]. MEG [THS15]. Memoriam [Hun13, Syl12, Chi12c]. Meng [Che12a]. Mervyn [Ke11a]. Mesh [TWY14]. Meta [GB18, LZ16, Ruk15, Ano10s]. Meta-Analysis [Ruk15, Ano10s]. Meta-Kriging [GB18]. Meta-Modeling [LZ16]. Metamodels [Tan17]. Method [BGMA14, HRV19, MW19, Nat12, THS15, WZ19, ZWZJ15, Ano10n, Mar19b]. Methodology [Hor10d, Oli11b]. Methods [Ano10p, Ano10f, Ano10m, BL11, Che11a, GJ10, Gui18, GW13, HEM10a, HEM10b, HDM^{+}15, Hor10b, Hor10a, Hor11d, Jar12a, JFEH14, Lip10a, Lip19a, LM15, MZ13, Mye12d, Nag11, Neu11, O’C10, SF10, Sin10, SFH^{+}13, SNY^{+}10, WMEW13, WAC10, Ano11k, Ano12f, Ban11, Hey12, Ke11a, Oli11a]. Metrics [GJ10, HEM10a, HEM10b, O’C10, SF10, Sin10, SB19, WAC10]. Metropolis [LKS16]. Meyers [Oli10a]. Michael [Ano11e, Fot10a, Jen10, Mye12a, Neu12a]. Microarray [Qiu12]. Microarrays [Ano12o]. Microscopy [CGR^{+}15]. Microstructure [SYW19]. Mike [Sea11b]. Mill [RKE10]. Min [Che11c]. Ming [Ano11f]. Ming-Hui [Ano11f]. Mini [Ste10b]. Mini-Issue [Ste10b]. Minimal [CJ14, JN11, LHDP14]. Minimax [Tan13]. Minimum [CYZ17, JDTW15, JWG^{+}19, LM15, RB10]. Mining [Ano10d, Ano11o, Hor10b, Oli12a, Sau12, Ano10h]. Misaligned [LG13, MBR14]. Misclassification [DSM13]. Mishura [Mye11c]. Missing [HRV19]. Mitchel [Ano12q]. Mithat [Gho10a]. Mitra [Mye11a]. Mixed [Apl10, CMS10, Gho11b, Gol12c, PFF14, QZW10a, QZW10b, SSG15, Tsu10, WBD10, ZCM^{+}18, Kat11, Oli10b]. Mixing [DLY^{+}17]. Mixture [BDB15, Hec10, KJB11, KSB16, KH19, LX18, LWM13]. Mixture-of-Mixtures [KJB11]. Mixtures [KJB11]. MMSE [TC17]. Mobile [CM14]. Model [BWMG14, CBBJ18, CJL14, CJ14, cCOP^{+}13, DGS13, Fra19, GGP10, GBH^{+}13, HGL^{+}13, HLSR15, JDLH16, JN17, KSB16, Kat11, KL17, LVB13, LG13, LZ14, LMT14, LZ15, Liu12, MLBM19, MMEDG14, Mor15, Oli11b, PSR17, PSB^{+}13, RKE10, Ruk15, SHD18, SL18, SdCR12, SD15, SDR15, SGC14, TW13, TG10, TQW13, WX10, WX14, WB18, VWGR11, WB11, XHX17, YC14, ZLD16, ZX17]. Model-Assisted [LMT14]. Model-Based [cCOP^{+}13, LZ15]. Model-Robust [SdCR12, SD15]. Modeling [Ahm19a, Ahm19c, Ahm19d, Ahm19e, Ano10e, Ano12f, Apl10, Ban10b, CM14, CWW17, CMS10, DJ15, GGL^{+}16a, GB18, Hly10a, HSL^{+}13, Hor10a, JSMS15, KJB11, KRSW19, KZS^{+}18, Li10, LZ16, MBR14, Mye11b, Nku10, PHK^{+}18, Pen10, QZW10a, QZW10b, RMVH11, Sea11b, SYW19, SB19, Tan18, TL19, Tsu10, TWY14, VSM13, WBD10, XMM16, XHM^{+}17, YPP19, YCZ17, ZSGM14, ZZZ^{+}15, dMvW10, CL16, Lip12c, Oli12a, Bai12]. Modelling [Ano10q, Ano12p, Lip19e]. Models [And11, Ano10i, AGV12, Bha12, BDB15, CG16, CBD^{+}17, CBC15, CHKC10, DJ15, DLLR17, Dri10, EW10, EPD18, FK14, GAZ15, Gho11b, Gho12a, Gol11a, Gol12c, GL12, GACH10, GWN13, HN12, HRC10, HCT17, Hof10, HW10, ILS^{+}17, JP13, KH19, LX18, LZ16, LZHZ18, LRL^{+}12b, LRL^{+}12a, Lip19a, MM19, ME12, Mor12, Oli10b, PD11, PFF14, PA17, PH16, TW12, TL16, TC17, Wlu12, XKHY18, XC10, YZW19, YR11, ZCW19, ZQZ11, ZCM^{+}18, Bai12, Hec10, Ros11, Sea11a]. Modern [Oli11a]. Modes [Liu12, SNM10]. Modified [JWFW19]. Moez [Lip11a]. Molenberghs [Gho11b]. Monari [Ano10p]. Monitoring [Apl10, BA18, BH13, CM14, CZ15a, CMS10, FQ18, KZS^{+}18, LCL12, LZHZ18, LMS15, MGSC18, PZQ16, PBH13, QZW10a, QZW10b, QZZ18, RTA11, Sen11, STZW16, SH14, TH10, Tsu10, VSM13, WMP18, WBD10, XWL18, YPS18, YZW12, ZQ18, ZSGM14, ZC18, ZWZJ15]. Monograph [Fot10b]. Monotone [LÇC12, TBMM15, YCZ17]. Monotonic [Tan16, Tan17]. Monte [Ano10n, Che12a, Mar19b, Neu11, Ng10, SH14, Ng10]. Montserrat [Bur12]. Morality [Lip12d]. Most [BFE^{+}19]. Mosteller [Ano10t]. Motion [TQW13]. Motoda [Lip10a]. Mou [Hor10c]. Mou-Hsiung [Hor10c]. Moving [BH13]. MR2655651 [RBM11]. MR2664866 [Got10b]. MR2752105 [Han10]. MR2752106 [Han10]. MR2967968 [Bor12, DM12, Lee12, OW12, SJ12]. MR2967975 [LRL^{+}12a]. MRI [Ahm19a, Mai11]. Muenchen [Gol11c, Sau11]. Muller [Ano11f, Che11a]. Multi [LC15, HTW17, Lip12b, TG17, XHM^{+}17, ZC18]. Multi-Fidelity [LC15, HTW17]. Multi-indicator [Lip12b]. Multi-Level [XHM^{+}17]. Multi-Parameter [ZC18]. Multi-Stratum [TG17]. Multicategory [WZ19]. Multichannel [WMP18]. Multiclass [WZ19]. Multidimensional [CM11, Lip12c]. Multifidelity [GBH^{+}13]. Multilayer [NS15]. Multilevel [GW13]. Multinomial [Tad13]. Multiple [AHS16, CGR^{+}15, GLLS14, HN12, ILS^{+}17, LACR11, MM19, Par14, PHD16, RMZ15, SWMW13, SDR15, Che11b]. Multiple-Case [RMZ15]. Multipollutant [JP13, PHK^{+}18]. Multiscale [MMEDG14, Ban10b]. Multisite [PHK^{+}18]. Multistratum [GG12, TG15]. Multivariate [Ano12h, BJFB14, CM17, CJL14, CZZ16, CZ15b, FOU13, HRC10, JP13, KRSW19, KKS^{+}14, KZS^{+}18, Kuh10, LZ15, LZ16, LZHZ18, Lip10b, PD11, PHK^{+}18, PZQ16, SG18, TS12, TML16, WMP18, XC10, Yu12, ZJT11, ZT11]. Murray [Bur11]. Murrell [Sym11]. Myatt [Hor10b].

N [Ahm19b, Ano10e, Bai11, Hor12d, Oli10b]. Naes [Hor11e]. Nanocomposites [ZZD^{+}14]. Nanoparticle [DLY^{+}17, SWW^{+}18]. Nanostructures [ZDH14]. Nanotubes [SRZ^{+}19]. Narayan [Hly10a]. NASA [BFK^{+}12]. Nason [BL11]. Natale [Ano11e]. Nature [MKBN19, SAR16]. Nearest [SG18]. Nearly [THT11]. Nearly-Isotonic [THT11]. Necip [Neu10b]. Need [SWWR14]. Negative [Qu12]. Neighborhood [GH16]. Neighbors [SG18]. Neil [Hor10d, Rut12]. Nested [GJ19, QAHS14]. Net [SAH^{+}17, WSZ14, ZLD16]. Network [Fre12, VSM13]. Network-Wide [VSM13]. Networks [Cha11, Lip19f, PGS11, Ano10d, Hly12, Lip11a, Lip11b]. Neuhauser [Neu12b]. Neural [Ano10d]. Neuronal [ZCG^{+}16]. Neuroscience [Ahm19d]. Neutron [BH13]. Niche [Pen10]. Nicholas [Hec11]. Nicole [Mai11]. Nikolaos [Ano10e]. Nikulin [Ano10e]. Nir [Bha12]. Noise [Ahm19a, BN12, JSMS15, Sen11, SRZ^{+}19]. Noisy [PGRC13a, Ran13]. NoMax [HWL16]. Nominations [Ano16k, Ano17i, Ano18g, Bre13, Bre14b, Hla15]. Non [Wv11]. Non-Normal [Wv11]. Nonbipartite [PSR17]. Noncollapsing [DSD12]. Nonconvex [PHBF17]. Nondestructive [LMT14, TMMH17]. Nonhomogeneous [RWP15]. Noninferiority [Ano11p]. Nonlinear [Got10b, Got10a, SAR16, WC12]. Nonlongitudinal [Par14]. Nonnegative [DS17, Jos13, PJR15, Xio10]. Nonorthogonal [MSE17]. Nonparametric [Ano11k, Apl10, BWMG14, CMS10, GWN13, JAP15, KMQ18, KRSA15, KZS^{+}18, LX18, LHDP14, MZ15, MQ11, QZW10a, QZW10b, QL11, QX13, QZZ18, RTA11, SWMW13, SHG12, Tsu10, WBD10, XWL18, Neu12b]. Nonrectangular [DSD12]. Nonregular [VX19]. Nonseparable [FOU13, LZ16]. Nonstationary [HCT17, PSB^{+}13, TMAZ18]. Normal [Wv11]. Norvaisa [Ano11n]. Note [Grö14, Jos13, LL16, MZ13]. Notes [Xio10]. Novak [Jar12a]. Novel [Nat12]. Numerical [CBD^{+}17, Ano11h, Mye12c]. Núñez [Gho12a]. Núñez-Antón [Gho12a].

O [Ano11f, Jen10]. Oberwolfach [Hor11b]. Object [Par14]. Objectivity [Ano10k]. Observational [Gol11b, Nat12]. Observations [LZHZ18, ML16, Pie13, YZW12]. Observing [MKBN19]. Ogunnaike [Cin11]. Okada [Ano10j]. Oliver [Hor11e]. Olivier [Gho11a]. One [AG12, Hof10, HRV19, ZC18]. One-Sided [Hof10, ZC18]. Online [JDLH16, LMS15, NHB^{+}13, RKE10, SWW^{+}16, XWL18, ZWZJ15]. Only [YT16]. Onset [VWGR11]. Operating [Gho10a]. Oppenlander [Wlu12]. Optimal [ASDMLF16, BMB15, Edw11, GJ19, HN18, LMW19, MGJ14, PSR17, Qu10, SRZ^{+}19, Tad13, ZHOY13, ZDH14, AA12, Atk12, Gho10b]. Optimality [HN18]. Optimization [AS16, CW16, CL16, FD15, GGL^{+}16a, HTW17, LACR11, MW19, PGRC13a, SCH17, SRZ^{+}19, TW12, YPP19]. Optimizing [PCWW16]. Optimum [SNM10, TL16]. Option [Hor10a]. Order [Lip12b]. Ordered [TS16]. Ordinal [CRCH18, dMvW10]. Ordinary [LZW15]. Organization [Ano11u]. Orientation [SGH13]. Orientations [DNV16, SWW^{+}18]. Origin [Car14, Haz10, Sen11]. Origin-Destination [Car14, Haz10]. Orthogonal [Bag10, Edw11, HQH16, MSE17, SSG15, SGS15, VGS19]. Orthogonalizing [XDHQ16]. Orthonormal [Tan15a]. Other [CHKC10, KSB16, Pie13]. Othmar [McC10]. Outbreak [SB10]. Outcomes [EDR16]. Outdoor [HDM^{+}15]. Outlier [YZW12]. Outliers [Far14, HS19, HRSV16, HRV19, PD11]. Output [Dri10, JDLH16, LZ16, Tan18]. Outputs [GBH^{+}13, HWL16, Mor12]. Over-Specified [CBD^{+}17]. Overview [Lip19d]. Owen [Ng11].

P [Ano10r, Ano10n, Ano10f, Ano11k, BL11, Che11a, Hly12, Hor10b, Hun13, Jar12b, Jen11, Lip12b, Mye12b, Nat12, Neu11, Oli12a]. Pacelli [Bar11]. Packet [PGS11]. Packets [Mal19]. Packing [He19]. Pair [The13]. Pairwise [FP17, LZ16, LZHZ18]. Pajevic [Che12b]. Palumbo [Ano11e]. Panel [BFE^{+}19]. Paola [Ano10p]. Paolo [Ano10f]. PAPERS [Ano13h]. Paradis [Ano12d]. Parallel [cCOP^{+}13, CGB^{+}14, GACH10]. Parameter [Got10b, Got10a, HLBS19, She17, TW13, Tan15a, WC12, XC10, ZC18]. Parameterized [CHKC10]. Parameters [Ruk15, SNM10, ZDH14]. Parametric [ZTEM17, ZYH14, dCCT15]. Pareto [CYZ17, KRSW19, LACR11, LCAc13, Zha10]. Partial [Ahm19e, CS16, Lip12b, ZL17]. Particle [LPVW13, LMW19, SPS18]. Particles [ASDMLF16, ZZD^{+}14]. Partition [GLLS14]. Partition-Based [GLLS14]. Partitioning [CHKC10, MLF^{+}16]. Path [Ahm19e]. Pathways [Par14]. Patil [Nat12, Lip12b]. Paul [Ano12e, Bar10, Gol11b, Lip12c, Sym11]. PCA [HRSV16, HRV19, MGSC18]. Peaks [KRSW19]. Pearl [Ano10i]. Penalized [FPG19, GFK^{+}18, HLBS19]. Performance [AS16]. Periodic [SHG12]. Permutation [Gui18]. Permutations [LL16, WHQ18]. Perret [Sea11a]. Perspective [EDR16, Kle13, Ban10b]. Perturbation [Lu11]. Peter [Ano11f, Ano11l, Atk12, Ban10a, Bur12, Che11a, Che11b, Hec10, Hey12, Lip10b, Nku10]. Pfaff [Ano10g]. Pflug [Mye11a]. Pharmacogenomics [Che11c]. Phase [TH10, BJFB14, CM17, PZQ16, WMP18, ZQ18]. Phase-I [PZQ16]. Phenomena [Cin11]. Phil [Lip12a, Sau10]. Philipp [Hor11c]. Philosophical [Lip12d]. Phones [CM14]. Phylogenetic [Ano12d]. Physical [LMT14, Lip12a, MBR14, Mor15]. Piccolo [Ano10p]. Piecewise [WX14]. Piera [Ano10f]. Piercesare [Ano12f]. Pierre [Esp11, Hor10a, Mye12c]. Pietro [Ano12f]. Pilipenko [Mye11c]. Pilz [AA12]. Pipeline [LM15]. Piquero [Ano12r]. Pixel [KMQ18]. Plane [SHD18]. Planning [Liu12, WMEW13, YXTS12]. Plans [Bag10, Edw11, TBMM15]. Platform [Hor12b]. Pleasure [Ano10t]. Plot [EG12, GG12, Mic10, MGJ14, SGS15, TW13, TG15, TG17]. Point [HN12, PZQ16, RMVH11, SGH13, WLHB15, YPP19]. Pointwise [ZPM14]. Poirier [Ban11]. Poisson [ME12, RWP15, STZW16]. Poli [Cha12]. Policy [Lip19e, Ban10a]. Political [Tad13]. Politics [Bar11]. Politis [Hor12d]. Polynomial [Tan15b, WLHB15]. Polynomials [Tan15a, Tan16]. Population [STZW16]. Populations [SAC17, SFH^{+}13]. Portal [BH13, PBH13]. Posterior [Apl12, FNL11]. Posteriors [JWG^{+}19]. Potential [EDR16]. Power [Bar11, LVB13, SS15]. Power-Law [SS15]. Pozdnyakov [Nag11]. pp [Ahm19b, Mar19a]. Practical [And11, Bre14a, Hor10b, Ano12h, Cob10]. Practicalities [For13]. Practice [Hor10d, Ano10m]. Practitioner [Oli12a]. Practitioners [ACH14, SWWR14]. Prado [Sea11b]. Precision [LCAc13, PGRC13a, TQW13]. Predicting [XHX17]. Prediction [And11, BLK19, GBH^{+}13, GH16, HM10, ML16, RKE10, SYW19, VSM13, WLWZ16, ZM19]. Predictions [HG11, HM13, Pen10]. Predictive [ZZZ^{+}15]. Predictor [CS16]. Predictors [Lip11c]. Preferential [SWW^{+}18]. Presence [DWD^{+}14, EW10, Far14, GJ19, SRZ^{+}19, ZCW19]. Press [Mar19a]. Price [CJL14, Ano10m]. Pricing [Hor10a]. Primer [Hly10b, Hor12a, Neu12c]. Principal [CFF13, WMP18]. Principle [TL19]. Principles [Bha12]. Printing [ZQ18]. Prior [ML16, WND^{+}18]. Prior-Free [ML16]. Prioritization [Lip12b]. Priors [GLLS14, JFJ^{+}19]. Prize [Ano10c, Ano11b]. Prizes [Ano10a, Ano10b, Ano11a, Ano12a, Ano13a, Ano16b, Ano19a, Ano14f, Ano17a]. Probabilistic [ML16, Bha12]. Probability [Cin11, Hor12c, Jen10, Laz10, Laz11, LMT14, McC12b, MGSC18, SAC17, Ano11d, Fot10b, San11, Ano11i, Ano12i, Hly10b]. Probit [XC10]. Problem [THS15]. Problems [Bre14a, CRBM19, Lip19f]. PROC [Ke11b]. Procedures [Cha12]. Proceedings [Ano11u]. Process [Apl12, DLLR17, FOU13, GL12, GH16, Gui18, HCT17, HSL^{+}13, KL17, KKS^{+}14, LQ14, LZHZ18, LMS15, QL11, RHK11, SCH17, SS15, TW12, Tan18, TL19, TH10, WX10, WB18, XMM16, XTP17, XHM^{+}17, YPP19, YC14, YCZ17, ZJT11]. Processes [BLK19, Fot10a, HS14, Hly10b, KZS^{+}18, LÇC12, LD11, MBR14, Mye11c, NSS11, Pen15, PSB^{+}13, QL11, RWP15, RZ16, TML16, WB18, YXTC14, ZTEM17, dCC11, dCCT15, Hly11]. Processing [Lip12c, SPS18]. Products [YXTC14, Ano10p]. Profile [Apl10, AGRI19, CMS10, LZHZ18, MGSC18, PZQ16, QZW10a, QZW10b, SAR16, Tsu10, WMP18, WBD10, YZW12]. Profiles [HRC10]. Prognosis [KZS^{+}18]. Prognostics [FPG19]. Programming [Ng11, Ng12, Oli10c, Rut12, SSG15, Sea12]. progressively [WYJ10]. Projection [Oli12c]. Proof [Bar11]. Properties [For13]. Protocols [FD15]. Publishing [Ahm19b].

QPSO [LMW19]. QQ [DJ15]. Quadratic [TW12]. Qualitative [DJ15, DLLR17, HLLY16, ZQZ11]. Quality [Ahm10, Ano10p, DJ15, Jen11, JFEH14, TS12, XTP17]. Quantification [CWW17, MMEDG14, Tan15b]. Quantifying [AGRI19, DLY^{+}17, NAEK15, WLWZ16]. Quantile [For13, GFK^{+}18, PBH13, PGRC13a, PT14, SWF16, Tan16, YCZ17]. Quantile-Based [PGRC13a]. Quantiles [PTZ19, KMM11]. Quantitative [Ano12r, DJ15, DLLR17, HLLY16, Ors19, ZQZ11, Hor12b, Mye11a]. Quantum [LMW19]. Quantum-Behaved [LMW19]. Quasi [Ng10, TC17]. Quasi-MMSE [TC17]. Quasi-Monte [Ng10]. Queues [CZ15a]. Queuing [Hly10a]. Quick [Bzi11]. Quick-Start [Bzi11].

R [AA12, Ano11n, Ano12g, Ano12o, Ano12r, Ban10b, Bur10, Gol11b, Hec11, Jar12b, Mar19a, Mye12a, Neu10a, Neu12c, Oli11a, AA12, Ano10g, Ano12d, Bai11, Boo10b, BL11, Che11b, Hor12a, Kuh10, Laz11, Mye12c, Mye12b, Neu11, Ng11, Oli10c, Oli10b, Pfa12, Qia10b, Sau10, Sau12, Yu12, Gol11c, Sau11]. R&R [BMS10]. Radiation [BH13, PBH13, SH14]. Radiative [FK14]. Radiographic [LG13]. Rainer [Lip12b]. Rajendra [Lu11]. Ramirez [Sar11, Sar11]. Ramsay [Bur10]. Random [Ano12n, CS12, DF12, Hly10b, Hof10, JSQ12, JS12a, JS12b, LW12, LRL^{+}12b, LRL^{+}12a, Pen15, SGH13, TH10, WSZ12, VWGR11, Cha11, Cin11]. Randomization [AS16]. Rank [TH18]. Ranking [Lip12b]. Ranks [SG18]. Raquel [Sea11b]. Rasch [AA12]. Rate [ASDMLF16, HM10]. Rates [DSM13]. Ratio [SGC14]. Rationality [Ano10k]. Raton [Mar19a]. Rattle [Sau12]. Raymond [Mye12a]. Real [JDLH16, YPS18]. Real-Time [JDLH16, YPS18]. Reasoning [Ano10k, Ano10i, Lip11b]. Receiver [Gho10a]. Receptor [HRC10, JP13, PHK^{+}18]. Reconstruction [dCCT15]. Recurrence [XHM^{+}17, ZYH14]. Recurrent [HW10, LÇC12]. Reduction [CSY18, CGB^{+}14, MZ15, WX14, ZM19]. Regions [AGRI19, DSD12, PHBF17, Ruk15, Tan13, You10]. Registration [QX13]. Regression [Ano10q, CM11, CRW18, CZ15b, CS16, EW10, FPG19, Ger15, GFK^{+}18, HLBS19, ILS^{+}17, JK11, KQ14, KMQ18, KH19, Mar11, ME12, MQ11, PH16, RMZ15, Ros11, SWMW13, Tad13, Tan16, THT11, TS16, WSZ14, YPP19, YZZ18, YCZ17, ZLD16, ZL17, ZPM14, Qu12, Ano11j, Lip11c, Mye11b]. Regression-Based [JK11]. Regularization [THS15]. Regularized [LZW15, YPP19]. Rejoinder [AANC13, GGL^{+}16b, HEM10b, JS12b, Jos12b, PGRC13b, QZW10b, WLDD14]. Related [Ahm19c, Ano12j, SAC17]. Reliability [Ano10k, Ano10d, Ano10e, CMR10, CMAC12, CY17, CBC15, GJ10, GACH10, GW13, HEM10a, HEM10b, LCAc13, MLBM19, O’C10, SF10, Sin10, SAC17, SDR15, VWGR11, WAC10, WHQ18, YZW19, ZW17, Ahm10, Ano12e]. Religion [Lip12d]. Remote [NKCB14, ZCW19]. Removal [Ano18h]. Renewal [XHM^{+}17, ZTEM17]. Renming [Hor12c]. Repair [LHDP14]. Repeatability [CRCH18, dMvW10]. Repeated [PGS11, WMEW13]. Replicated [LZHZ18]. Replication [BHGL19]. Report [Chi11, Chi12b, Chi13, Apl17, E19, Qiu14, Qiu15, Qiu16b, Ste10a]. Reported [WYHT18]. Representation [CWW17]. Reproducibility [CRCH18, dMvW10]. Reproducing [BLK19, YZZ18]. Resampling [Oli12a]. Research [Ano10s, Ano11u, CCCC13]. Resolution [MKBN19, TH18]. Response [HJM15, LMW19, SRZ^{+}19, Wv11]. Responses [CM14, DJ15, GAZ15, MM19]. Restricted [SGC14]. Retrospective [Ger15]. Return [WYHT18]. Returned [YT16]. Reversibility [Hly12]. Review [AA12, Ahm19a, Ahm10, Ahm19b, And11, Ano10s, Ano10g, Ano10j, Ano10r, Ano10h, Ano10n, Ano10q, Ano10p, Ano10t, Ano10k, Ano10o, Ano10d, Ano10e, Ano10f, Ano10i, Ano10m, Ano10l, Ano11d, Ano11o, Ano11p, Ano11h, Ano11n, Ano11f, Ano11g, Ano11l, Ano11e, Ano11j, Ano11m, Ano11i, Ano11c, Ano11k, Ano12h, Ano12n, Ano12f, Ano12j, Ano12e, Ano12i, Ano12c, Ano12g, Ano12k, Ano12d, Ano12q, Ano12p, Ano12o, Ano12l, Ano12r, Ano12m, Atk12, Bai11, Bai12, Ban10b, Ban10a, Ban11, Bar10, Bar11, Bha12, Boo10b, Boo10a, Bur10, BL11, Bur11, Bur12, Bzi11, Cha11, Cha12, Che11c, Che11a, Che11b, Che12a, Che12b, Cin11, Cob10, Esp11, Fot10b, Fot10a, Fre12, Gho10b, Gho10a, Gho11a, Gho11b, Gho12a, Gho12b, Gho19, Gol11b]. Review [Gol11a, Gol12c, Gol12b, Gol12a, Gru11, Guo19, Hec10, Hec11, Hey12, Hin11, Hly10a, Hly10b, Hly11, Hly12, Hor10b, Hor10d, Hor10c, Hor10a, Hor11f, Hor11d, Hor11b, Hor11c, Hor11a, Hor11e, Hor12c, Hor12b, Hor12d, Hor12a, Jar12b, Jar12a, Jen10, Jen11, Kat11, Ke10, Ke11a, Kuh10, Laz10, Laz11, Li10, Lip10a, Lip10b, Lip11a, Lip11c, Lip11b, Lip12b, Lip12d, Lip12c, Lip12a, Lip19a, Lip19b, Lip19c, Lip19d, Lip19f, Lip19e, Lu11, Mag11, Mai11, Mar19a, Mar19b, McC10, McC12b, McC12c, McC12a, Mic10, Mye11a, Mye11b, Mye11c, Mye12a, Mye12c, Mye12b, Mye12d, Nat12, Neu10b, Neu10a, Neu11, Neu12c, Neu12b, Neu12a, Ng10, Ng11, Ng12, Nku10, Oli10a, Oli10c, Oli10b, Oli11b, Oli11a, Oli12c, Oli12a, Oli12b, Ors19, Par11, Pen10]. Review [Pfa12, Qia10b, Qia10a, Qiu12, Qu11, Qu12, Ros11, Rut12, San11, Sar11, Sau10, Sau11, Sau12, Sea11a, Sea11b, Sea12, Sen11, SB19, Sym11, Van10, Wie11, WMI11, Wlu12, Ye10, You10, Yu12]. Reviews [Ano10u, Ano10v, Ano10w, Ano10x, Ano11q, Ano11r, Ano11s, Ano11t, Ano12s, Ano12t, Ano12u, Ano12v, Ano13c, Ano13d, Ano13e, Ano13f, Ano14b, Ano14c, Ano14d, Ano14e, Ano15d, Ano15a, Ano15b, Ano15c, Ano16d, Ano16e, Ano16f, Ano16g, Ano17f, Ano17c, Ano17d, Ano17e, Ano18a, Ano18b, Ano18c]. Reward [Ahm19b]. Riccomagno [Ano10f]. Richard [Ahm19b, Hor12d].

Richly [CHKC10]. Rick [Laz10, Sea12]. Ridge [Mar11]. right [KMM11, WYJ10]. right-censored [WYJ10]. Rimas [Ano11n]. Rinaldo [Ano12i]. Risk [Hor12b, Mye11c, Ahm19b]. Risks [SS15, YXTS12]. Robert [Bzi11, Gol11c, Neu11, Ng11, Qu11, Sau11, Van10]. Roberto [Cha12]. Robinson [Mye12b, Ng11]. Robust [AMY14, BN12, CRW18, CFF13, DMW10, Far14, FD15, Got10b, Got10a, JWFW19, LSP13, MW19, Mar11, ME12, SD12, She17, SdCR12, SD15, TW12, TW13, Tan15a, THS15, Ano11k]. Robustness [JGBM19]. ROC [Qia10a]. Rogantin [Ano10f]. Rojo [Ano12k]. Role [Neu10b]. Rolling [RKE10]. Ron [Ahm10]. Ronald [Cha11]. Rongling [Che11c]. Rosen [Fot10a, GLLS14]. Rosenbaum [Gol11b]. Rosenblatt [Hor12d]. Ross [Ano11l]. Rotated [He19]. Rotations [SGH13]. Row [Qu10]. Row-Column [Qu10]. Rowena [Ban10a]. Rowwise [HRV19]. Rubinstein [Ano10n]. Rueven [Ano10n]. Ruggeri [Ahm10]. Rule [TL16]. Rumours [Lip11a]. Run [TC17, VX19]. Run-to-Run [TC17]. Runs [MKBN19, MLF^{+}16]. Rutledge [Bzi11].

S [Ahm10, Ano10e, Bur10, Oli10a, Pfa12, Sar11]. Salmaso [Ano10p]. Samaniego [Gru11]. Sample [McC12c, SNM10, Wył10]. Samples [QAHS14, SHD18, KMM11]. Sampling [CBC15, JWG^{+}19, LVB13, Laz11, LQ14, LZ15, LMS15, Nat12, QWKR14, SH14, XWL18, WYJ10, Esp11, Ng10]. Sandrine [Ano12l]. Sara [Ano12m]. Sarkar [Kuh10]. SAS [Ano11o, Sau11, Sea11a, Sea12, Ke11b, Gho10a, Ng12, Oli10a, Hec11, Ke11a]. SAS/IML [Sea11a, Sea12]. SAS(R) [Bzi11, Rut12]. Satellite [ZCW19]. Saunders [ZLD16]. Saveliev [Oli10b]. Sawitzki [Boo10b]. Scalable [GB18]. Scale [EW10, HLP13, HP12, NHQ17, RTA11]. Scaled [CS16]. Scan [NHB^{+}13, Nag11]. Scheme [LQ14, STZW16]. Schemes [FNL11, JN16]. Schinazi [Ano12i]. Schlattmann [Hec10]. Schmee [Wlu12]. Schmid [Jen11]. Schreier [Ke11b]. Science [Ahm19b, DNV16, Ano10k, Hor11e]. Sciences [Ahm19c, Lip12a, Lip19b, Mye12a, Mag11, Mye12a]. Scientific [Hor12b, Ng11]. Scott [Che11a]. Screening [ACH14, BGMA14, DWD^{+}14, JN16, JN17, LSP13, MSE17, MDS12, QX14, Qu10, SWWR14, WB11, ZHOY13]. Search [GH16, Jon13, WB11]. Season [NAEK15]. Seasonal [AGMRS11, Hor11a]. Secch [Ano12f]. Second [Ano12g]. Sectional [ZZD^{+}14]. Selected [Ano11n, Ano11l, Ano11m, Ano12k, Ano12l, Ano12m, Hor12c, Hor12d]. Selecting [LCAc13, MSE17]. Selection [AGV12, BLK19, Fra19, HLBS19, JN17, LZ14, PJR15, SWMW13, TW13, Wad16, WB11, ZLD16, Lip10a]. Self [STZW16]. Self-Starting [STZW16]. Semifoldover [Edw11]. Seminars [Hor11b]. Semiparametric [LZW15, SDR15, XKHY18, YXTC14]. Seneta [Ano11l]. Sense [Hor10b]. Sensing [NKCB14, ZCW19]. Sensitivity [FK14, LWM13, MDS12]. Sensitivity-Based [MDS12]. Sensor [JDLH16, PJR15, ZS15]. Sensory [Hor11e]. Sentiment [Tad13]. Separation [BWMG14]. Sequences [AHS16, SHG12]. Sequencing [WHQ18]. Sequential [CGR^{+}15, CWW17, DASS17, EPD18, LC15, JLLM^{+}16, JDTW15, KAT18, LHTD18, PSB^{+}13, RBM08, RBM11, SCH17, SH14, XQW13, BHGL19]. Serguei [Jar12a]. Series [Ano10g, Cob10, Dri10, FP17, GACH10, Hor11a, NS15, PTZ19, Qia10b, Sea11b, TMAZ18]. Service [CS12, DF12, JSQ12, JS12a, JS12b, LW12, WSZ12]. Services [Ano10p]. Set [CGB^{+}14, WLHB15, WHQ18]. Seth [Hor11b]. Sets [WMT12]. Setting [SWW^{+}16]. Settings [BN12, Got10b, Got10a]. Severini [Oli12b]. Shape [SNM10, dCC11]. Shapes [SWW^{+}18]. Sharad [Nat12]. Sharpening [Gui18]. Shein [Gol12b]. Shein-Chung [Gol12b]. Shephard [Hor10d]. Shifts [TS12]. Shin [Hor12d]. Shock [SFH^{+}13]. Short [PHD16]. Short-Term [PHD16]. Shrinkage [JFJ^{+}19]. Sided [Hof10, ZC18]. Sieving [LPVW13]. Sign [ZT11]. Signal [Ano11c, JFJ^{+}19, SPS18]. Signals [KZS^{+}18, ZSGM14]. Significance [LM13]. Sik [Qia10b]. Silverstein [Ano12n]. Simon [Mag11]. Simple [ZQZ11]. Simplicial [MGSC18]. Simulated [WB11]. Simulating [Lip19e]. Simulation [BHGL19, Bre14a, CBC15, Kle13, Laz11, MKBN19, MLF^{+}16, Ng11, RHK11, Tan16, TQW13, Ano10n, Mar19b]. Simulations [PT14]. Simulators [FK14, GBH^{+}13, OY17, Ran13]. Simultaneous [CZ15b, ZPM14]. Single [GL12, WX14]. Single-Index [GL12, WX14]. Singular [Oli12c]. Sinisa [Che12b]. Situation [SH14]. Size [DLY^{+}17]. Sized [TH10]. Sizes [STZW16, SNM10, VX19, Zha11]. Skiadas [Ano10d]. Skovgaard [Lip11c]. Slack [KSB16]. Slack-Variable [KSB16]. Sliced [BMB15, DASS17, He19, HQH16]. Small [SHD18, Gho12b]. Smith [Ban10a, Chi12c, Oli10b]. Smooth [YPS17, YPS18]. Smooth-Sparse [YPS17]. Smoothing [CJL14, MQ11, Mye12d]. Social [Lip19b, Lip19e, Mag11, McC10, Mye12a]. Societal [Lip19e]. Software [SDR15, Sea12, Oli10c]. Solka [Mar19a, Ano12g]. Solve [Bre14a, Lip19f]. Solving [SAH^{+}17, THS15]. Some [Jos12b, LS13, PGK16, SG18, SJ12, Xio10]. Sonar [ZCM^{+}18]. Sonar-Terrain [ZCM^{+}18]. Song [Hor12c, Nku10]. Sorin [Ano12o]. Sought [Ano16k, Ano17i, Ano18g, Bre13, Bre14b, Hla15]. Sounder [BFK^{+}12]. Source [BH13, HRC10, RMVH11]. Sources [WH19]. Space [Ano10m, DSD12, HP12, JGBM19]. Space-Filling [DSD12, JGBM19]. Spaces [BLK19, SD15, YZZ18]. Spanos [Ano10k]. Sparse [CWW17, CSY18, CHWE11, CFF13, GFK^{+}18, HRSV16, MZ13, SAH^{+}17, TS16, YPS17, YPS18]. Sparsity [QLS^{+}19]. Spatial [Bur12, CS12, DF12, GB18, HCT17, HG11, JSQ12, JFJ^{+}19, JS12a, JS12b, KH19, Li10, LW12, MKBN19, PHK^{+}18, PSB^{+}13, SHL15, SWF16, WSZ12, ZQ18, ZZD^{+}14]. Spatial-Temporal [PSB^{+}13]. Spatially [JP13, KL17, LD11, NS15, NSS11]. Spatio [CG16, cCOP^{+}13, NKCB14, YPS18]. Spatio-Temporal [CG16, cCOP^{+}13, NKCB14, YPS18]. Special [Ano13h, CCCC13]. Specified [CBD^{+}17]. Specimen [SNM10]. Spector [Sau10]. Spectral [Ano12n]. SpectralTemporal [RMVH11]. Speed [PHD16, TMAZ18, Ano12l]. Speeding [GH16]. Sphere [He19]. Splines [AGV12, Mye12d, YCZ17]. Split [GG12, Mic10, MGJ14, SGS15, TW13, TG15, TG17]. Split-Plot [GG12, MGJ14, SGS15, TG15, TG17]. Springer [Ahm19b]. SPSS [Oli10a, Sau11]. Spurious [Pie13]. SQL [Ke11b]. Squares [Ahm19e, CS16, SPS18, XDHQ16, ZL17]. Stable [RHK11]. Stage [MDS12, TBMM15]. Staggered [AG12]. Staggered-Level [AG12]. Stance [Lip12d]. Standard [DSM13, SD12]. Start [Bzi11]. Starting [STZW16]. Startup [TH10]. Stata [Gol11c]. State [AHS16, DLY^{+}17, GGP10, LVB13, WB18]. State-Dependent [GGP10]. Statement [Ano18h]. Static [Car14]. Stationarity [TEN14]. Statistic [Ahm19d]. Statistical [Ano11p, Ano11f, Ano11k, Ano12f, Ano12i, Car14, CG16, Cha12, CZ15a, Dav13, GJ10, Gho10b, Gol12b, HEM10a, HEM10b, Haz10, Hec11, Hey12, HDM^{+}15, Jen11, JY15, Ke11a, LPVW13, LQ14, LS13, Lip10b, LM15, LM13, MKBN19, Mai11, Mal19, Neu12b, NHQ17, O’C10, Oli11a, Pen10, Pfa12, QL11, SWW^{+}16, SF10, SB10, Sin10, SS15, SNY^{+}10, TMMH17, TG17, VSM13, WAC10, WMI11, Wlu12, XTP17, YR11, ZCG^{+}16, ZJT11, dCC11, Ahm19a, Ano10q, Ano10p, Ano10o, Ano11o, Bur11, Che11c, Che12b, Fre12, Hor11d, Hor11c, Jar12b, Lip12c, Ng12, Sea12, Ye10, You10]. Statisticians [Fot10b, Ano11h]. Statistics [Ahm10, Ano10t, Ano10e, Ano11c, Boo10b, BL11, Cin11, Fot10b, Hor11b, Hor12c, Jen10, Nag11, NHB^{+}13, Neu10b, Qu11, SWW^{+}18, Ano10f, Bur12, Gho12b, Hin11, Hly10b, McC12c, Neu12a, Ano10r, Ano12o, Hor11e]. Stefan [Ano11p]. Stefano [Ano10h]. Step [Sar11]. Step-by-Step [Sar11]. Stephen [Ano10t]. Stepwise [CGB^{+}14]. Stereology [Fot10b]. Steve [Che12a]. Stevens [Hor12a]. Stewart [Hor11f]. Steyerberg [And11]. Stochastic [BHGL19, BA18, CBC15, Hly12, Mye11c, OY17, PT14, SCH17, She17, Tan15b, Tan16, TQW13, WH19, ZLD16, Ano12p, Hly11, Hor10c]. Stockwell [Pen10]. Storm [NAEK15]. Strategic [Lip19f]. Strategies [DWD^{+}14, LC15, KJB11, PJW13]. Strategy [Bar11, GG12, HWL16, LAMO^{+}16, LMS15, XWL18]. Stratum [TG17]. Streams [RTA11, XWL18, YPS18, ZWZJ15]. Street [Ban10a]. Strength [VX19, VGS19, WLWZ16]. Strength- [VGS19]. String [ZHOY13]. Strong [LZ14]. Stroock [Ano11i]. Struchiner [Wie11]. Structure [Wad16, ZHOY13]. Structure-Activity [ZHOY13]. Structured [YPP19, ZCM^{+}18]. Structures [Ano10q, FOU13]. Studies [BMS10, Car14, CMAC12, Cob10, CRCH18, Fra19, LPVW13, Lip19c, WMEW13, Gol11b, Nat12, Wie11]. Study [Atk12, LD11, LCAc13, Ano11j]. Sturmfels [Hor11b]. Subgraphs [NHB^{+}13]. Subgrouped [BJFB14]. Subject [TH10]. Subset [WB11]. Subsurface [MMEDG14]. Subsystem [LCAc13]. Sudjianto [Han10]. Suess [Laz11]. Sufficient [CSY18, MZ15, ZM19]. Suhov [Hly10b]. Sullivant [Hor11b]. Sun [Ano11f]. Superimposed [ZTEM17]. Superposition [TL19]. Supersaturated [PCWW16]. Support [Mor15, Lip12a, Hor11f]. Surface [SRZ^{+}19, ZQ18, dCCT15]. Surfaces [BA18, JDTW15, KQ14]. Surrogate [CWW17, TWY14]. Surrogates [Mor12]. Surveillance [RWP15]. Survival [Ano10e, Bai12, ZLD16, Ano12q]. Susanne [Ano12h]. Sutradhar [Gol12c]. Swarm [LMW19, PCWW16]. Switzerland [Ahm19b]. Sylvan [Nag11]. Synthesis [Ano10s]. System [Ano13h, CMAC12, DSM13, GW13, Haz10, LCAc13, MKBN19, QX14, ZW17]. Systematic [ZCW19]. Systems [Ano12p, CMR10, DMW10, DJ15, GJ10, GACH10, HEM10a, HEM10b, Hor10c, JSMS15, LAMO^{+}16, Lip12b, O’C10, SF10, Sin10, VWGR11, WAC10, XHM^{+}17, ZZZ^{+}15]. Székely [HS17].

T [Mic10, Oli12b, You10, Yu12]. Tadashi [Ano10j]. Tail [MW19, RZ16]. Taillie [Nat12]. Tails [JWFW19]. Takane [Oli12c]. Takeuchi [Oli12c]. Tamaki [Gho10b]. Tango [Hor11d]. Taniguchi [Gho10b]. Tanur [Ano10t]. Tasks [WHQ18]. Taylor [Bar11]. Technique [Jos12a, LMW19, RMZ15]. Techniques [Ban10a, LC15, KRSA15, McC12c, PCWW16, Bha12]. Technologies [Sym11]. Technology [Ano10e]. Technometrics [Ano14f, Ano14g, Ano17a, Ano10a, Ano10b, Ano11a, Ano12w, Ano12a, Ano13h, Ano13g, Ano13a, Ano16h, Ano16b, Ano17g, Ano18d, Ano19b, Ano19a, Apl17, E19]. Telecommunications [BVW10, Han10]. Telman [Sen11]. Temperature [BWMG14, CG16]. Temperature-Emissivity [BWMG14]. Temporal [CG16, cCOP^{+}13, NKCB14, PSB^{+}13, SWF16, YPS18]. Tensor [FPG19, YPP19, ZL17]. Terje [Hor12b]. Term [PHD16]. Terrain [ZCM^{+}18]. Terry [Ano12l]. Test [LHDP14, PSR17, TEN14, XTP17, XKHY18]. Testing [FP17, HLP13, ILS^{+}17, LÇC12, LRL^{+}12b, LRL^{+}12a, SZ19, Ano11p]. Tests [LHTD18, Liu12, Neu12b, SG18, SGC14, TL16, ZC18, Wył10]. Textured [BA18, TEN14]. Thas [Gho11a]. Theil [Lip11c]. Their [Hly10b, Mal19, SWW^{+}18]. Theoretical [Qu11]. Theory [Ano10d, Ano11i, Hly10a, Par11, You10, Mye11c, Oli12b, Lip19b, Mye11c]. Thickness [LM15]. Thomas [Ano10q, Ano11k]. Thorn [Neu12c]. Thoughts [Jos12b, SJ12]. Three [DNV16, MZ13, VX19]. Three-Dimensional [DNV16]. Threshold [Wad16]. Thresholded [WMP18]. Thresholds [KRSW19]. Throughput [LSP13, Qu10]. Time [Ahm19c, Ano10g, Ano10m, CMR10, Dri10, FP17, HRC10, HM13, JDLH16, Mor12, NS15, PTZ19, QWKR14, TS16, TMAZ18, YPS18, ZSGM14, Qia10b, Sea11b]. Time-Between-Event [QWKR14]. Time-Dependent [HRC10]. Time-Lagged [TS16]. Time-Varying [Mor12]. Times [Fot10a]. Titanium [LMT14]. Tobias [Ano12e, Ban11]. Toine [Ano10r]. Tolerance [EW10, Hof10, You10, ZPM14]. Tomic [Hor11e]. Ton [Ano10r]. Tool [Ano11u, LM13]. Tools [Jar12b]. Topics [Qu11, Ano12j]. Tormod [Hor11e]. Torsten [Che11b, Pfa12]. Toshiro [Hor11d]. Total [HS19]. Tough [Bre14a]. Traffic [CM14, VSM13]. Transfer [FK14, ZZZ^{+}15]. Transformation [MZ15, ZY17]. Transient [WB18]. Transit [Haz10]. Translation [Hor11c]. Transportation [Car14]. Tree [PH16, ZHOY13]. Treed [KKS^{+}14]. Trend [LÇC12, XHM^{+}17]. Trend-Renewal [XHM^{+}17]. Trials [Ano10r, Ano11g, Che11a, Gol12b, Gol12a]. Trindade [Ano12e]. Trumbo [Laz11]. Truncated [ZSGM14]. Tukey [CRW18]. Tunable [PGRC13a, TQW13]. Tuning [HLBS19]. Turbines [PHD16]. Tutz [Ano10q]. Twitter [Tad13, XMM16]. Two [CGR^{+}15, Edw11, ES17, EJLN17, FQ18, Grö14, HL17, MSE17, MDS12, PCWW16, SWW^{+}18, SdCR12, SNM10, TBMM15, THS15, Van10, VX19, VGS19, WMT12, Zha11]. Two-Dimensional [CGR^{+}15]. Two-Factor [ES17, Grö14, WMT12]. Two-Level [Edw11, ES17, EJLN17, Grö14, HL17, MSE17, PCWW16, SdCR12, Van10, VX19, VGS19]. Two-Stage [MDS12, TBMM15]. Two-Way [THS15, Zha11]. Type [KMM11, WYJ10]. Types [SDR15].

U [Hly10a]. Ultrasonic [LMT14]. Unbalanced [Hof10]. Uncertainties [AGRI19]. Uncertainty [CWW17, CGB^{+}14, Fra19, Got10b, Got10a, MMEDG14, NAEK15, RKE10, Tan15b, TG10, WLWZ16]. Understanding [HS19, Lip12d]. Unequal [Zha11]. Unified [ZCM^{+}18]. Unit [KKS^{+}14]. Univariate [QL11, QX14, WLHB15]. Unknown [WSZ14]. Unmixing [SHL15]. Unreplicated [EDR16, Fra19, HL17]. Unrest [HLSR15]. Unreturned [YT16]. Up-and-Down [XTP17]. Updating [JDLH16, SWW^{+}16, And11]. Use [GJ10, HEM10a, HEM10b, HM10, O’C10, SF10, Sin10, WAC10]. Use-Rate [HM10]. Used [Bre14a]. Users [XMM16, Gol11c, Rut12, Sau11]. Using [Ano11o, Ano12o, AGV12, Bai12, CRW18, Che11b, EDR16, FPG19, GBH^{+}13, LC15, GW13, HRC10, HCT17, HGL^{+}13, JFJ^{+}19, JDLH16, JWFW19, Jos12a, JDTW15, JWG^{+}19, KMQ18, KRSA15, Ke10, Ke11b, KZS^{+}18, LVB13, LMT14, LQ14, LZ16, LMZ19, LD11, Lip19f, LCAc13, MMEDG14, MBR14, Ng11, OY17, OW17, Par14, Pfa12, PCWW16, PSR17, PHD16, Sar11, SCH17, She17, SYW19, SdCR12, Tan15a, TL19, WND^{+}18, XMM16, XTP17, XC10, YCZ17, ZHOY13, ZCM^{+}18, BLK19]. Utility [WHQ18]. Utilizing [LACR11].

V [Ano10s, Cob10]. Vaccine [Wie11]. Valentine [Ano10s]. Valid [Gol11a]. Validation [And11, LC15]. Value [Jar12a, Oli12c, Wad16]. Valued [BLK19, HNKF18, JAP15]. Values [HRV19]. Variability [HLBS19]. Variable [AGV12, BN12, BLK19, KSB16, MZ15, WSZ14, ZLD16]. Variables [HLLY16, Lip19a, Pen15, ZCW19]. Variance [MGJ14, Oli10a]. Variation [HS19, SAR16]. Variational [LAMO^{+}16]. Variations [Mic10]. Varmuza [Lip10b]. Varying [DSM13, GAZ15, KL17, Mor12, NSS11, STZW16]. Vector [Hor11f, NS15]. Vedel [Fot10b]. Verbeke [Gho11b]. Verdooren [AA12]. Vertov [Lip19c]. Very [NKCB14]. Via [HN18, YPP19, CGR^{+}15, CM14, Fra19, HTW17, MGSC18, PJR15, PT14, RKE10, YPS17, YPS18, ZLD16]. Vicente [Gho12a]. View [Ano11i]. Vine [XHX17]. Violence [Lip12d]. Virginia [Ano12h]. Visualising [Lip19c]. Visualization [Hor10b, Kuh10, PTZ19]. Visualizing [AGRI19]. Vladimir [Ano11n, Nag11]. Vol [Fot10b, Hor11b]. Volinsky [Han10]. Volume [Ano10z, Ano11w, Hly10b]. Voting [Bar11].

W [Ahm10, And11, Ano11i, Ano11k, Ano12n, Ano12c, Bai11, Jar12b, Jen11, McC10, Mye11b, Qu11, Van10]. Walker [Oli10b]. Wallenstein [Nag11]. Wallis [Ano12c]. Walter [Mic10]. Waltraud [Ano10e]. Wang [Qiu12, Mye12d]. Warning [XHX17]. Warping [ZSGM14]. Warranty [HM10, LCL12]. Wave [BWMG14, SFH^{+}13]. Wavelet [HNKF18, Mal19, BL11]. Way [THS15, Zha11]. Wayne [Hor10b]. Weak [LZ14]. Wear [GGP10]. Weathering [HDM^{+}15]. Wegkamp [Ano12m]. Weibull [SNM10]. Weighted [WZ19]. Weighting [JK11]. Weihs [Ano11u]. Weisberg [Gol11a]. Weisburd [Ano12r]. Well [DM12, HRV19, Lee12, OW12]. Wellek [Ano11p]. Wendy [Ano12g, Mar19a]. Weron [Jar12b]. Werthamer [Ahm19b]. West [Sea11b]. Westfall [Che11b]. Wicklin [Sea12]. Wide [VSM13]. Wilcox [Oli11a]. Wilkinson [Ano12p]. Wilks [Han10]. Willem [Ano12m]. William [Ano16k, Ano17i, Ano18g, Cob10, Ke11a, Bre13, Bre14b, Hla15, Syl12]. Williams [Sau12]. Wilrich [Jen11]. Wind [PHD16, TMAZ18]. Window [RZ16, ZYH14]. Window-Censored [RZ16]. Winkler [McC10]. Wise [Far14]. Within [Ke11b]. Wojbor [Ano11c]. Wojtek [Qia10a]. Wolfgang [Ano10j, McC12a]. Work [Lip19c]. Workplace [MBR14]. Works [Ano11n, Ano11l, Ano11m, Ano12k, Ano12l, Ano12m, Hor12c, Hor12d]. World [Lip11b]. Woyczynski [Ano11c]. Wu [Che11c, Kat11, Par11]. Wynn [Ano10f].

X [Nku10]. X.-K [Nku10]. Xiao [Che12a]. Xiao-Li [Che12a]. Xiaohui [Qiu12]. xx [Ahm19b]. xxv [Mar19a].

Yanai [Oli12c]. Ye [Ano11f]. Yuedong [Mye12d]. Yuliya [Mye11c]. Yuri [Hly10b].

Zhidong [Ano12n]. Zidong [Qiu12]. Ziegel [Ano10c, Ano11b, Ano13b, Ano14a, Ano16a, Ano16c, Ano17b]. Ziegler [Mye12a]. Zimmerman [Gho12a]. Zuur [Bai11, Oli10b]. Zwet [Ano12m]. Zwinderman [Ano10r].

References

Adams:2012:BRB

[AA12] Wayne Adams and Mark Anderson. Book review: Optimal Experimental Design with R by Dieter Rasch; Jurgen Pilz; L. R. Verdooren; Albrecht Gebhardt. Technometrics, 54(2):203, May 2012. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

            (electronic). URL http://www.jstor.org/stable/41714877.

Albrecht:2013:R

[AANC13] Mark C. Albrecht, Thomas A. Albrecht, Christopher J. Nachtsheim, and R. Dennis Cook. Rejoinder. Technometrics, 55(3): 292–295, 2013. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Anderson-Cook:2014:CTG

[ACH14] Christine M. Anderson-Cook and Michael S. Hamada. Comment: Toward guidelines for practitioners on screening designs and analysis. Technometrics, 56(1):16–19, 2014. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). See [DWD^{+}14].

Arnouts:2012:SLD

[AG12] Heidi Arnouts and Peter Goos. Staggered-level designs for experiments with more than one hard-to-change factor. Technometrics, 54(4):355–366, November 2012. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/41714920.

Alonso:2011:SDF

[AGMRS11] Andrés M. Alonso, Carolina García-Martos, Julio Rodríguez, and María Jesús Sánchez. Seasonal dynamic factor analysis and bootstrap inference: Application to electricity market forecasting. Technometrics, 53(2):137–151, May 2011. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://

www.jstor.org/stable/23209726.

Azzimonti:2019:PEV

[AGRI19] Dario Azzimonti, David Ginsbourger, Jérémy Rohmer, and Déborah Idier. Profile extrema for visualizing and quantifying uncertainties on excursion regions: Application to coastal flooding. Technometrics, 61(4):474–493, 2019. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Antoniadis:2012:VSA

[AGV12] Anestis Antoniadis, Irène Gijbels, and Anneleen Verhasselt. Variable selection in additive models using P -splines. Technometrics, 54(4):425–438, November 2012. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/41714926.

Ahmed:2010:BRB

[Ahm10] S. E. Ahmed. Book review: Encyclopedia of Statistics in Quality and Reliability by Fabrizio Ruggeri; Ron S. Kenett; Frederick W. Faltin. Technometrics, 52(2):263, May 2010. CODEN TCMTA2.

ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://

www.jstor.org/stable/27867239.

Ahmad:2019:BRS

[Ahm19a] M. Rauf Ahmad. Book review: Statistical Analysis of Noise Estimation in MRI: Modeling, Filtering and Estimation. Technometrics, 61(3):426–427, 2019. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Ahmed:2019:BRR

[Ahm19b] S. Ejaz Ahmed. Book review: Risk and Reward: The Science of Casino Blackjack (2nd ed.), by N. Richard Werthamer. Cham, Switzerland: Springer International Publishing AG, 2018, xx + 150 pp., ISBN: 978-3-319-91384-1, 978-3-319-91385-8 (ebook). DOI: 10.1007/978-3-319-91385-8. Technometrics, 61(4): 567, 2019. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Ahmed:2019:CTM

[Ahm19c] S. Ejaz Ahmed. Continuous time modeling in the behavioural and related sciences. Technometrics, 61(4):567–568, 2019. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Ahmed:2019:DNS

[Ahm19d] S. Ejaz Ahmed. Dynamic neuroscience statistic, modeling, and control. Technometrics, 61(4):568, 2019. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Ahmed:2019:PLS

[Ahm19e] S. Ejaz Ahmed. Partial least squares path modeling. Technometrics, 61(4):568–569, 2019. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Arnesen:2016:BDC

[AHS16] Petter Arnesen, Tracy Holsclaw, and Padhraic Smyth. Bayesian detection of changepoints in finite-state Markov chains for multiple sequences. Technometrics, 58(2):205–213, 2016. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Agostinelli:2014:REG

[AMY14] Claudio Agostinelli, Alfio Marazzi, and Victor J. Yohai. Robust estimators of the generalized log-gamma distribution. Technometrics, 56(1):92–101, 2014. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Albrecht:2013:EDE

[ANAC13] Mark C. Albrecht, Christopher J. Nachtsheim, Thomas A. Albrecht, and R. Dennis Cook. Experimental design for engineering dimensional analysis. Technometrics, 55(3):257–270, 2013. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Albrecht:2014:LE

[ANAC14] Marck C. Albrecht, Christopher J. Nachtsheim, Thomas A. Albrecht, and R. Dennis Cook. Letter to the Editor. Technometrics, 56(2):268, 2014. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Andrews:2011:BRB

[And11] Chris Andrews. Book review: Clinical Prediction Models: A Practical Approach to Development, Validation, and Updating by Ewout W. Steyerberg. Technometrics, 53(1):101–102, February 2011. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/40997302.

Anonymous:2010:TPa

[Ano10a] Anonymous. The 2008 Technometrics prizes. Technometrics, 52 (1):1–3, 2010. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Anonymous:2010:TPb

[Ano10b] Anonymous. The 2009 Technometrics prizes. Technometrics, 52 (4):373–375, 2010. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Anonymous:2010:AZP

[Ano10c] Anonymous. Announcement of ziegel prize. Technometrics, 52 (3):371, 2010. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Anonymous:2010:BRBl

[Ano10d] Anonymous. Book review: Advances in Data Analysis — Theory and Applications to Reliability and Inference, Data Mining, Bioinformatics, Lifetime Data, and Neural Networks by Christos H. Skiadas. Technometrics, 52(3):370, August 2010. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/27867276.

Anonymous:2010:BRBm

[Ano10e] Anonymous. Book review: Advances in Degradation Modeling: Applications to Reliability, Survival Analysis, and Finance (Statistics for Industry and Technology) by M. S. Nikulin; Nikolaos Limnios; N. Balakrishnan; Waltraud Kahle; Catherine Huber-Carol. Technometrics, 52(4):469, November 2010. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/40997266.

Anonymous:2010:BRBn

[Ano10f] Anonymous. Book review: Algebraic and Geometric Methods in Statistics by Paolo Gibilisco; Eva Riccomagno; Maria Piera Rogantin; Henry P. Wynn. Technometrics, 52(4):469, November 2010. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/40997267.

Anonymous:2010:BRBb

[Ano10g] Anonymous. Book review: Analysis of Integrated and Cointegrated Time Series With R (2nd ed.) by Bernhard Pfaff. Technometrics, 52(1):145, February 2010. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/40586696.

Anonymous:2010:BRBe

[Ano10h] Anonymous. Book review: Biological Data Mining by Jake Y. Chen; Stefano Lonardi. Technometrics, 52(2):264, May 2010. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/27867241.

Anonymous:2010:BRBo

[Ano10i] Anonymous. Book review: Causality — Models, Reasoning, and Inference (2nd ed.) by Judea Pearl. Technometrics, 52(4):469– 470, November 2010. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
40997268.

Anonymous:2010:BRBc

[Ano10j] Anonymous. Book review: Cooperation in Classification and Data Analysis by Akinori Okada; Tadashi Imaizumi; Hans-Hermann Bock; Wolfgang Gaul. Technometrics, 52(1):145, February 2010. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
40586697.

Anonymous:2010:BRBj

[Ano10k] Anonymous. Book review: Error and Inference: Recent Exchanges on Experimental Reasoning, Reliability, and the Objectivity and Rationality of Science by Deborah G. Mayo; Aris Spanos. Technometrics, 52(3):369, August 2010. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://

www.jstor.org/stable/27867274.

Anonymous:2010:BRBq

[Ano10l] Anonymous. Book review: Introduction to Machine Learning (2nd ed.) by Ethem Alpaydin. Technometrics, 52(4):470, November 2010. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/40997270.

Anonymous:2010:BRBp

[Ano10m] Anonymous. Book review: Price Indexes in Time and Space — Methods and Practice by Luigi Biggeri; Guido Ferrari. Technometrics, 52(4):470, November 2010. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/40997269.

Anonymous:2010:BRBf

[Ano10n] Anonymous. Book review: Simulation and the Monte Carlo Method (2nd ed.) by Rueven Y. Rubinstein; Dirk P. Kroese. Technometrics, 52(3):368, August 2010. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/27867270.

Anonymous:2010:BRBk

[Ano10o] Anonymous. Book review: Statistical Analysis of Management Data (2nd ed.) by Hubert Gatignon. Technometrics, 52(3):369– 370, August 2010. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
27867275.

Anonymous:2010:BRBh

[Ano10p] Anonymous. Book review: Statistical Methods for the Evaluation of Educational Services and Quality of Products by Matilde Bini; Paola Monari; Domenico Piccolo; Luigi Salmaso. Technometrics, 52(3):368–369, August 2010. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/27867272.

Anonymous:2010:BRBg

[Ano10q] Anonymous. Book review: Statistical Modelling and Regression Structures — Festschrift in Honour of Ludwig Fahrmeir by Thomas Kneib; Gerhard Tutz. Technometrics, 52(3):368, August 2010. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/27867271.

Anonymous:2010:BRBd

[Ano10r] Anonymous. Book review: Statistics Applied to Clinical Trials (4th ed.) by Ton J. Cleophas; Aeilko H. Zwinderman; Toine F. Cleophas; Eugene P. Cleophas. Technometrics, 52(2):264, May 2010. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/27867240.

Anonymous:2010:BRBa

[Ano10s] Anonymous. Book review: The Handbook of Research Synthesis and Meta-Analysis (2nd ed.) by Harris Cooper; Larry V. Hedges; Jeffrey C. Valentine. Technometrics, 52(1):144–145, February 2010. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/40586695.

Anonymous:2010:BRBi

[Ano10t] Anonymous. Book review: The Pleasure of Statistics — The Autobiography of Frederick Mosteller by Stephen E. Fienberg; David C. Hoaglin; Judith M. Tanur. Technometrics, 52(3):369, August 2010. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/27867273.

Anonymous:2010:BRa

[Ano10u] Anonymous. Book reviews. Technometrics, 52(1):138–145, 2010. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Anonymous:2010:BRc

[Ano10v] Anonymous. Book reviews. Technometrics, 52(2):258–264, 2010. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Anonymous:2010:BRd

[Ano10w] Anonymous. Book reviews. Technometrics, 52(3):362–370, 2010. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Anonymous:2010:BRe

[Ano10x] Anonymous. Book reviews. Technometrics, 52(4):461–470, 2010. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Anonymous:2010:EC

[Ano10y] Anonymous. Editorial collaborators. Technometrics, 52(4):471– 472, 2010. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Anonymous:2010:IV

[Ano10z] Anonymous. Index to volume 52 (2010). Technometrics, 52(4): 473–476, 2010. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Anonymous:2011:TP

[Ano11a] Anonymous. The 2010 Technometrics prizes. Technometrics, 53 (4):333–335, 2011. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Anonymous:2011:AZP

[Ano11b] Anonymous. Announcement of ziegel prize. Technometrics, 53 (3):332, 2011. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Anonymous:2011:BRBm

[Ano11c] Anonymous. Book review: A First Course in Statistics for Signal Analysis (2nd ed.) by Wojbor A. Woyczynski. Technometrics, 53 (4):447, November 2011. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/41714967.

Anonymous:2011:BRBa

[Ano11d] Anonymous. Book review: Applied Probability (2nd ed.) by Kenneth Lange. Technometrics, 53(1):107, February 2011. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/40997312.

Anonymous:2011:BRBi

[Ano11e] Anonymous. Book review: Data Analysis and Classification by Francesco Palumbo; Carlo Natale Lauro; Michael J. Greenacre. Technometrics, 53(3):331, August 2011. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://

www.jstor.org/stable/23210423.

Anonymous:2011:BRBf

[Ano11f] Anonymous. Book review: Frontiers of Statistical Decision Making and Bayesian Analysis — In Honor of James O. Berger by Ming-Hui Chen; Dipak K. Dey; Peter Muller; Dongchu Sun; Keying Ye. Technometrics, 53(2):222, May 2011. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://

www.jstor.org/stable/23209754.

Anonymous:2011:BRBg

[Ano11g] Anonymous. Book review: Fundamentals of Clinical Trials (4th ed.) by Lawrence M. Friedman; Curt D. Furberg; David L. DeMets. Technometrics, 53(2):222–223, May 2011. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/23209755.

Anonymous:2011:BRBd

[Ano11h] Anonymous. Book review: Numerical Analysis for Statisticians (2nd ed.) by Kenneth Lange. Technometrics, 53(2):222, May 2011. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/23209752.

Anonymous:2011:BRBl

[Ano11i] Anonymous. Book review: Probability Theory — An Analytic View (2nd ed.) by Daniel W. Stroock. Technometrics, 53(4):447, November 2011. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
41714966.

Anonymous:2011:BRBj

[Ano11j] Anonymous. Book review: Regression Estimators — A Comparative Study (2nd ed.) by Marvin H. J. Gruber. Technometrics, 53(4):446, November 2011. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/41714964.

Anonymous:2011:BRBn

[Ano11k] Anonymous. Book review: Robust Nonparametric Statistical Methods (2nd ed.) by Thomas P. Hettmansperger; Joseph W. McKean. Technometrics, 53(4):447, November 2011. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/41714968.

Anonymous:2011:BRBh

[Ano11l] Anonymous. Book review: Selected Works of C. C. Heyde by Ross Maller; Ishwar Basawa; Peter Hall; Eugene Seneta. Technometrics, 53(2):223, May 2011. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/23209756.

Anonymous:2011:BRBk

[Ano11m] Anonymous. Book review: Selected Works of Debabrata Basu by Anirban DasGupta. Technometrics, 53(4):446–447, November 2011. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/41714965.

Anonymous:2011:BRBe

[Ano11n] Anonymous. Book review: Selected Works of R. M. Dudley by Evarist Gine; Vladimir Koltchinskii; Rimas Norvaisa. Technometrics, 53(2):222, May 2011. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/23209753.

Anonymous:2011:BRBb

[Ano11o] Anonymous. Book review: Statistical Data Mining Using SAS Applications by George Fernandez. Technometrics, 53(1):107, February 2011. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
40997313.

Anonymous:2011:BRBc

[Ano11p] Anonymous. Book review: Testing Statistical Hypotheses of Equivalence and Noninferiority (2nd ed.) by Stefan Wellek. Technometrics, 53(1):107–108, February 2011. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://

www.jstor.org/stable/40997314.

Anonymous:2011:BRa

[Ano11q] Anonymous. Book reviews. Technometrics, 53(1):98–108, 2011. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Anonymous:2011:BRc

[Ano11r] Anonymous. Book reviews. Technometrics, 53(2):209–223, 2011. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Anonymous:2011:BRd

[Ano11s] Anonymous. Book reviews. Technometrics, 53(3):322–331, 2011. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Anonymous:2011:BRe

[Ano11t] Anonymous. Book reviews. Technometrics, 53(4):440–447, 2011. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Anonymous:2011:CTR

[Ano11u] Anonymous. Classification as a tool for research: Proceedings of the 11th IFCS Biennial Conference and 33rd Annual Conference of the Gesellschaft für … Data Analysis, and Knowledge Organization by Hermann Locarek-Junge; Claus Weihs. Technometrics, 53(3):330–331, August 2011. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/23210422.

Anonymous:2011:EC

[Ano11v] Anonymous. Editorial collaborators. Technometrics, 53(4):448– 449, 2011. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Anonymous:2011:IV

[Ano11w] Anonymous. Index to volume 53 (2011). Technometrics, 53(4): 450–453, 2011. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Anonymous:2012:TP

[Ano12a] Anonymous. The 2011 Technometrics prizes. Technometrics, 54 (4):334–336, 2012. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Anonymous:2012:EB

[Ano12b] Anonymous. 2012 editorial board. Technometrics, 54(4):451, 2012. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Anonymous:2012:BRBg

[Ano12c] Anonymous. Book review: A Beginner’s Guide to Discrete Mathematics, 2nd Edition by W. D. Wallis. Technometrics, 54(2):204, May 2012. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/41714881.

Anonymous:2012:BRBj

[Ano12d] Anonymous. Book review: Analysis of Phylogenetic and Evolution With R (2nd ed.) by Emmanuel Paradis. Technometrics, 54 (3):330–331, August 2012. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/41714914.

Anonymous:2012:BRBe

[Ano12e] Anonymous. Book review: Applied Reliability (3rd ed.) by Paul A. Tobias; David C. Trindade. Technometrics, 54(1):106, February 2012. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/41714852.

Anonymous:2012:BRBc

[Ano12f] Anonymous. Book review: Complex Data Modeling and Computationally Intensive Statistical Methods by Pietro Mantovan; Piercesare Secch. Technometrics, 54(1):105–106, February 2012. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/41714850.

Anonymous:2012:BRBh

[Ano12g] Anonymous. Book review: Exploratory Data Analysis With MATLAB (Second Edition) by Wendy L. Martinez; Angel R. Martinez; Jeffrey L. Solka. Technometrics, 54(3):330, August 2012. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/41714912.

Anonymous:2012:BRBa

[Ano12h] Anonymous. Book review: Practical Multivariate Analysis (5th ed.) by Abdelmomem Afifi; Susanne May; Virginia A. Clark. Technometrics, 54(1):105, February 2012. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://

www.jstor.org/stable/41714848.

Anonymous:2012:BRBf

[Ano12i] Anonymous. Book review: Probability with Statistical Applications, 2nd Edition by Rinaldo B. Schinazi. Technometrics, 54 (2):204, May 2012. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
41714880.

Anonymous:2012:BRBd

[Ano12j] Anonymous. Book review: Recent Advances in Functional Data Analysis and Related Topics by Frederic Ferraty. Technometrics, 54(1):106, February 2012. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/41714851.

Anonymous:2012:BRBi

[Ano12k] Anonymous. Book review: Selected Works of E. L. Lehmann by Javier Rojo. Technometrics, 54(3):330, August 2012. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/41714913.

Anonymous:2012:BRBn

[Ano12l] Anonymous. Book review: Selected Works of Terry Speed by Sandrine Dudoit. Technometrics, 54(4):446, November 2012. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/41714940.

Anonymous:2012:BRBp

[Ano12m] Anonymous. Book review: Selected Works of Willem van Zwet by Sara van de Geer; Marten Wegkamp. Technometrics, 54(4):446, November 2012. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
41714942.

Anonymous:2012:BRBb

[Ano12n] Anonymous. Book review: Spectral Analysis of Large Dimensional Random Matrices (2nd ed.) by Zhidong Bai; Jack W.

Silverstein. Technometrics, 54(1):105, February 2012. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/41714849.

Anonymous:2012:BRBm

[Ano12o] Anonymous. Book review: Statistics and Data Analysis for Microarrays Using R and Bioconductor, 2nd Edition by Sorin Dräghici. Technometrics, 54(4):445–446, November 2012. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/41714939.

Anonymous:2012:BRBl

[Ano12p] Anonymous. Book review: Stochastic Modelling for Systems Biology, 2nd Edition by Darren J. Wilkinson. Technometrics, 54 (4):445, November 2012. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/41714938.

Anonymous:2012:BRBk

[Ano12q] Anonymous. Book review: Survival Analysis (3rd ed.) by David G. Kleinbaum; Mitchel Klein. Technometrics, 54(3):331, August 2012. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/41714915.

Anonymous:2012:BRBo

[Ano12r] Anonymous. Book review: The Handbook of Quantitative Criminology by Alex R. Piquero; David Weisburd. Technometrics, 54 (4):446, November 2012. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/41714941.

Anonymous:2012:BRa

[Ano12s] Anonymous. Book reviews. Technometrics, 54(1):97–106, 2012. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Anonymous:2012:BRc

[Ano12t] Anonymous. Book reviews. Technometrics, 54(2):198–205, 2012. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Anonymous:2012:BRd

[Ano12u] Anonymous. Book reviews. Technometrics, 54(3):319–331, 2012. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Anonymous:2012:BRe

[Ano12v] Anonymous. Book reviews. Technometrics, 54(4):439–446, 2012. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Anonymous:2012:TEC

[Ano12w] Anonymous. Technometrics editorial collaborators. Technometrics, 54(4):448–449, 2012. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Anonymous:2012:E

[Ano12x] Anonymous. Erratum. Technometrics, 54(4):450, 2012. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Anonymous:2013:TP

[Ano13a] Anonymous. The 2012 Technometrics prizes. Technometrics, 55 (4):383–385, 2013. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Anonymous:2013:ZAA

[Ano13b] Anonymous. 2012 ziegel award announcement. Technometrics, 55(4):573, 2013. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Anonymous:2013:BRa

[Ano13c] Anonymous. Book reviews. Technometrics, 55(1):108–117, 2013. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Anonymous:2013:BRb

[Ano13d] Anonymous. Book reviews. Technometrics, 55(2):247–255, 2013. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Anonymous:2013:BRc

[Ano13e] Anonymous. Book reviews. Technometrics, 55(3):374–382, 2013. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Anonymous:2013:BRd

[Ano13f] Anonymous. Book reviews. Technometrics, 55(4):562–572, 2013. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Anonymous:2013:TEC

[Ano13g] Anonymous. Technometrics editorial collaborators. Technometrics, 55(4):574–575, 2013. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Anonymous:2013:CPS

[Ano13h] Anonymous. CALL FOR PAPERS system informatics — a special issue of Technometrics. Technometrics, 55(1):1, 2013. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Anonymous:2013:C

[Ano13i] Anonymous. Corrigendum. Technometrics, 55(3):373, 2013. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Anonymous:2013:EEB

[Ano13j] Anonymous. EOV Ed board. Technometrics, 55(4):ebi, 2013. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Anonymous:2014:ZAA

[Ano14a] Anonymous. 2013 ziegel award announcement. Technometrics, 56(3):416, 2014. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Anonymous:2014:BRa

[Ano14b] Anonymous. Book reviews. Technometrics, 56(1):123–130, 2014. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Anonymous:2014:BRb

[Ano14c] Anonymous. Book reviews. Technometrics, 56(2):261–267, 2014. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Anonymous:2014:BRc

[Ano14d] Anonymous. Book reviews. Technometrics, 56(3):407–415, 2014. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Anonymous:2014:BRd

[Ano14e] Anonymous. Book reviews. Technometrics, 56(4):551–557, 2014. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Anonymous:2014:TP

[Ano14f] Anonymous. The 2013 Technometrics Prizes. Technometrics, 56 (4):419–421, 2014. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Anonymous:2014:TEC

[Ano14g] Anonymous. Technometrics editorial collaborators. Technometrics, 56(4):558–559, 2014. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Anonymous:2014:EEB

[Ano14h] Anonymous. EOV Ed board. Technometrics, 56(4):ebi, 2014. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Anonymous:2015:BRa

[Ano15a] Anonymous. Book reviews. Technometrics, 57(1):138–143, 2015. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Anonymous:2015:BRb

[Ano15b] Anonymous. Book reviews. Technometrics, 57(2):292–300, 2015. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Anonymous:2015:BRc

[Ano15c] Anonymous. Book reviews. Technometrics, 57(3):440–443, 2015. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00401706.2015.1070657.

Anonymous:2015:BR

[Ano15d] Anonymous. Book reviews. Technometrics, 57(4):586–593, 2015. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00401706.2015.1106893.

Anonymous:2015:EBE

[Ano15e] Anonymous. Editorial board EOV. Technometrics, 57(4):594, 2015. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00401706.2015.1103551.

Anonymous:2016:ZAAa

[Ano16a] Anonymous. 2014 Ziegel Award announcement. Technometrics, 58(1):152–153, 2016. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00401706.2015.1105697.

Anonymous:2016:TP

[Ano16b] Anonymous. The 2015 Technometrics prizes. Technometrics, 58 (4):532–534, 2016. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Anonymous:2016:ZAAb

[Ano16c] Anonymous. 2015 Ziegel Award announcement. Technometrics, 58(4):535, 2016. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Anonymous:2016:BRa

[Ano16d] Anonymous. Book reviews. Technometrics, 58(1):148–151, 2016. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00401706.2016.1142315.

Anonymous:2016:BRb

[Ano16e] Anonymous. Book reviews. Technometrics, 58(2):277–281, 2016. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Anonymous:2016:BRc

[Ano16f] Anonymous. Book reviews. Technometrics, 58(3):404–??, 2016. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Anonymous:2016:BRd

[Ano16g] Anonymous. Book reviews. Technometrics, 58(4):524–531, 2016. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Anonymous:2016:TEC

[Ano16h] Anonymous. Technometrics editorial collaborators. Technometrics, 58(4):536–537, 2016. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Anonymous:2016:EBE

[Ano16i] Anonymous. Editorial board EOV. Technometrics, 58(4):ebi, 2016. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Anonymous:2016:E

[Ano16j] Anonymous. Erratum. Technometrics, 58(3):411–??, 2016. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Anonymous:2016:NSW

[Ano16k] Anonymous. Nominations sought for 2016 William G. Hunter Award. Technometrics, 58(2):282, 2016. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Anonymous:2017:TP

[Ano17a] Anonymous. The 2016 Technometrics Prizes. Technometrics, 59 (4):552–554, 2017. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Anonymous:2017:ZAA

[Ano17b] Anonymous. 2016 Ziegel Award announcement. Technometrics, 59(4):555, 2017. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Anonymous:2017:BRa

[Ano17c] Anonymous. Book reviews. Technometrics, 59(1):126–133, 2017. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Anonymous:2017:BRb

[Ano17d] Anonymous. Book reviews. Technometrics, 59(2):271–280, 2017. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Anonymous:2017:BRc

[Ano17e] Anonymous. Book reviews. Technometrics, 59(3):404–412, 2017. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Anonymous:2017:BR

[Ano17f] Anonymous. Book reviews. Technometrics, 59(4):542–551, 2017. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Anonymous:2017:TEC

[Ano17g] Anonymous. Technometrics editorial collaborators. Technometrics, 59(4):556–557, 2017. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Anonymous:2017:EBE

[Ano17h] Anonymous. Editorial board EOV. Technometrics, 59(4):ebi, 2017. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Anonymous:2017:NSW

[Ano17i] Anonymous. Nominations sought for 2017 William G. Hunter Award. Technometrics, 59(2):281, 2017. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Anonymous:2018:BRa

[Ano18a] Anonymous. Book reviews. Technometrics, 60(1):124–132, 2018. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Anonymous:2018:BRb

[Ano18b] Anonymous. Book reviews. Technometrics, 60(2):263–269, 2018. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Anonymous:2018:BRc

[Ano18c] Anonymous. Book reviews. Technometrics, 60(3):408–414, 2018. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Anonymous:2018:TEC

[Ano18d] Anonymous. Technometrics editorial collaborators. Technometrics, 60(4):557–558, 2018. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Anonymous:2018:Cb

[Ano18e] Anonymous. Clarification. Technometrics, 60(4):559, 2018. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Anonymous:2018:Ca

[Ano18f] Anonymous. Corrigendum. Technometrics, 60(1):133, 2018. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Anonymous:2018:NSW

[Ano18g] Anonymous. Nominations sought for 2018 William G. Hunter Award. Technometrics, 60(2):271, 2018. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Anonymous:2018:SR

[Ano18h] Anonymous. Statement of removal. Technometrics, 60(2):270, 2018. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Anonymous:2019:TP

[Ano19a] Anonymous. The 2017 Technometrics prizes. Technometrics, 61 (1):136–138, 2019. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Anonymous:2019:TEC

[Ano19b] Anonymous. Technometrics editorial collaborators. Technometrics, 61(4):i–ii, 2019. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Anonymous:2019:C

[Ano19c] Anonymous. Correction. Technometrics, 61(2):281, 2019. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Anonymous:2019:EA

[Ano19d] Anonymous. Editorial announcement. Technometrics, 61(1):1, 2019. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Apley:2010:NPM

[Apl10] Daniel W. Apley. [Nonparametric Profile Monitoring by Mixed Effects Modeling]: Comment. Technometrics, 52(3):277–280, August 2010. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/27867246.

Apley:2012:PDC

[Apl12] Daniel W. Apley. Posterior distribution charts: A Bayesian approach for graphically exploring a process mean. Technometrics, 54(3):279–293, August 2012. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/41714896.

Editor:2017:TER

[Apl17] Daniel W. Apley. Technometrics 2017 Editor’s report. Technometrics, 59(4):413–415, 2017. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Avron:2016:HPK

[AS16] Haim Avron and Vikas Sindhwani. High-performance kernel machines with implicit distributed optimization and randomization. Technometrics, 58(3):341–??, 2016. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Amo-Salas:2016:OED

[ASDMLF16] Mariano Amo-Salas, Elvira Delgado-Márquez, and Jesús López-Fidalgo. Optimal experimental designs in the flow rate of particles. Technometrics, 58(2):269–276, 2016. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Atkinson:2012:BRB

[Atk12] A. C. Atkinson. Book review: Optimal Design of Experiments: A Case Study Approach by Peter Goos; Bradley Jones. Technometrics, 54(3):328–329, August 2012. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/41714910.

Bui:2018:MDA

[BA18] Anh Tuan Bui and Daniel W. Apley. A monitoring and diagnostic approach for stochastic textured surfaces. Technometrics, 60(1): 1–13, 2018. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Bagchi:2010:MEP

[Bag10] Sunanda Bagchi. Main-effect plans orthogonal through the block factor. Technometrics, 52(2):243–249, May 2010. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/27867228. With supple-

mentary material available online.

Bailey:2011:BRB

[Bai11] Mark Bailey. Book review: A Beginner’s Guide to R by Alain F. Zuur; Elena N. Ieno; Erik H. W. G. Meesters. Technometrics, 53(1):105, February 2011. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/40997308.

Bailey:2012:BRB

[Bai12] Mark Bailey. Book review: Modeling Survival Data Using Frailty Models by David D. Hanagal. Technometrics, 54(3):324, August 2012. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/41714904.

Bandyopadhyay:2010:BRBb

[Ban10a] Dipankar Bandyopadhyay. Book review: Measuring Efficiency in Health Care: Analytic Techniques and Health Policy by Rowena Jacobs; Peter C. Smith; Andrew Street. Technometrics, 52(4): 461–462, November 2010. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/40997253.

Bandyopadhyay:2010:BRBa

[Ban10b] Dipankar Bandyopadhyay. Book review: Multiscale Modeling: A Bayesian Perspective by Marco A. R. Ferreira; Herbert K. H. Lee. Technometrics, 52(2):259–260, May 2010. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://

www.jstor.org/stable/27867233.

Bandyopadhyay:2011:BRB

[Ban11] Dipankar Bandyopadhyay. Book review: Bayesian Econometric Methods by Gary Koop; Dale J. Poirier; Justin L. Tobias. Technometrics, 53(2):210, May 2011. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/23209732.

Barker:2010:BRB

[Bar10] Lawrence Barker. Book review: Bayesian Evaluation of Informative Hypotheses by Herbert Hoijtink; Irene Klugkist; Paul A. Boelen. Technometrics, 52(3):365–366, August 2010. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/27867265.

Barrett:2011:BRB

[Bar11] J. Douglas Barrett. Book review: Mathematics and Politics: Strategy, Voting, Power, and Proof by Alan Taylor; Allison Pacelli. Technometrics, 53(1):100–101, February 2011. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/40997300.

Brown:2015:GBM

[BDB15] L. Brown, A. N. Donev, and A. C. Bissett. General blending models for data from mixture experiments. Technometrics, 57(4):449– 456, 2015. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00401706.2014.947003.

Bardwell:2019:MRC

[BFE^{+}19] Lawrence Bardwell, Paul Fearnhead, Idris A. Eckley, Simon Smith, and Martin Spott. Most recent changepoint detection in panel data. Technometrics, 61(1):88–98, 2019. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Braverman:2012:MDA

[BFK^{+}12] Amy J. Braverman, Eric J. Fetzer, Brian H. Kahn, Evan M. Manning, Robert B. Oliphant, and Joao P. Teixeira. Massive dataset analysis for NASA’s Atmospheric Infrared Sounder. Technometrics, 54(1):1–15, February 2012. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/41714827.

Boukouvalas:2014:ESM

[BGMA14] Alexis Boukouvalas, John Paul Gosling, and Hugo Maruri-Aguilar. An efficient screening method for computer experiments. Technometrics, 56(4):422–431, 2014. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Burr:2013:MNS

[BH13] Tom Burr and Michael S. Hamada. Moving neutron source detection in radiation portal monitoring. Technometrics, 55(3):296– 308, 2013. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Bhattacharjya:2012:BRB

[Bha12] Debarun Bhattacharjya. Book review: Probabilistic Graphical Models: Principles and Techniques by Daphne Koller; Nir Friedman. Technometrics, 54(1):99, February 2012. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/41714837.

Binois:2019:RES

[BHGL19] Mickaël Binois, Jiangeng Huang, Robert B. Gramacy, and Mike Ludkovski. Replication or exploration? Sequential design for stochastic simulation experiments. Technometrics, 61(1):7–23, 2019. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Bell:2014:DFM

[BJFB14] Richard C. Bell Jr., L. Allison Jones-Farmer, and Nedret Billor. A distribution-free multivariate Phase I location control chart for subgrouped data from elliptical distributions. Technometrics, 56 (4):528–538, 2014. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Burr:2011:BRBa

[BL11] Tom Burr and Claire Longo. Book review: Wavelet Methods in Statistics With R by G. P. Nason. Technometrics, 53(2):212–213, May 2011. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/23209736.

Bueno-Larraz:2019:VSP

[BLK19] Beatriz Bueno-Larraz and Johannes Klepsch. Variable selection for the prediction of C[0, 1]-valued autoregressive processes using reproducing kernel Hilbert spaces. Technometrics, 61(2):139–153, 2019. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Ba:2015:OSL

[BMB15] Shan Ba, William R. Myers, and William A. Brenneman. Optimal sliced Latin hypercube designs. Technometrics, 57(4):479–487, 2015. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00401706.2014.957867.

Browne:2010:LGR

[BMS10] Ryan Browne, Jock MacKay, and Stefan Steiner. Leveraged gauge R&R studies. Technometrics, 52(3):294–302, August 2010. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/27867251.

Bingham:2012:NVS

[BN12] Derek Bingham and Vijayan N. Nair. Noise variable settings in robust design experiments. Technometrics, 54(4):388–397, November 2012. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/41714923.

Booth:2010:BRBb

[Boo10a] David Booth. Book review: Bayesian Biostatistics and Diagnostic Medicine by Lyle D. Broemeling. Technometrics, 52(3):367– 368, August 2010. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
27867269.

Booth:2010:BRBa

[Boo10b] David Booth. Book review: Computational Statistics: An Introduction to R by Gunther Sawitzki. Technometrics, 52(3):364, August 2010. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/27867261.

Bornkamp:2012:CCI

[Bor12] Björn Bornkamp. Comment: “Comparison With Iterated Laplace Approximation” [MR2967968]. Technometrics, 54(3):225–227, August 2012. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
41714887.

Brenneman:2013:NSW

[Bre13] William A. Brenneman. Nominations sought for William G. Hunter Award. Technometrics, 55(2):121, 2013. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Brenneman:2014:CSU

[Bre14a] William A. Brenneman. Comment: Simulation used to solve tough practical problems. Technometrics, 56(1):19–20, 2014. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Brenneman:2014:NSW

[Bre14b] William A. Brenneman. Nominations sought for William G. Hunter Award. Technometrics, 56(2):131, 2014. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Burr:2010:BRB

[Bur10] Tom Burr. Book review: Functional Data Analysis With R and MATLAB by J. Ramsay; G. Hooker; S. Graves. Technometrics, 52(4):466, November 2010. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/40997262.

Burr:2011:BRBb

[Bur11] Tom Burr. Book review: Statistical Inference — An Integrated Bayesian/Likelihood Approach by Murray Aitken. Technometrics, 53(4):443–444, November 2011. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/41714961.

Burr:2012:BRB

[Bur12] Tom Burr. Book review: Handbook of Spatial Statistics by Alan Gelfand; Peter Diggle; Montserrat Fuentes; Peter Guttorp. Technometrics, 54(2):202, May 2012. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/41714875.

Becker:2010:FDT

[BVW10] Richard A. Becker, Chris Volinsky, and Allan R. Wilks. Fraud detection in telecommunications: History and lessons learned. Technometrics, 52(1):20–33, February 2010. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://

www.jstor.org/stable/40586677. See discussion [Han10].

Berrett:2014:BNM

[BWMG14] Candace Berrett, Gustavious Paul Williams, Todd Moon, and Jacob Gunther. A Bayesian nonparametric model for temperatureemissivity separation of long-wave hyperspectral images. Technometrics, 56(2):200–211, 2014. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Bzik:2011:BRB

[Bzi11] Thomas Bzik. Book review: Just Enough SAS(R): A Quick-Start Guide to SAS(R) for Engineers by Robert A. Rutledge. Technometrics, 53(2):216–217, May 2011. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://

www.jstor.org/stable/23209742.

Carvalho:2014:BSA

[Car14] Luis Carvalho. A Bayesian statistical approach for inference on static origin-destination matrices in transportation studies. Technometrics, 56(2):225–237, 2014. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Cao:2018:MCC

[CBBJ18] Fang Cao, Shan Ba, William A. Brenneman, and V. Roshan Joseph. Model calibration with censored data. Technometrics, 60 (2):255–262, 2018. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Choe:2015:ISR

[CBC15] Youngjun Choe, Eunshin Byon, and Nan Chen. Importance sampling for reliability evaluation with stochastic simulation models. Technometrics, 57(3):351–361, 2015. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/00401706.2014.1001523.

Chakraborty:2017:ENM

[CBD^{+}17] Avishek Chakraborty, Derek Bingham, Soma S. Dhavala, Carolyn C. Kuranz, R. Paul Drake, Michael J. Grosskopf, Erica M. Rutter, Ben R. Torralva, James P. Holloway, Ryan G. McClarren, and Bani K. Mallick. Emulation of numerical models with overspecified basis functions. Technometrics, 59(2):153–164, 2017. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Chair:2013:SIC

[CCCC13] Kary Myers Coda Chair, Earl Lawrence Coda Co-Chair, and Hugh A. Chipman. Special issue: Conference on data analysis (CoDA)-Exploring data-focused research across the department of energy. Technometrics, 55(4):389, 2013. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Chen:2013:PEA

[cCOP^{+}13] Wei chen Chen, George Ostrouchov, David Pugmire, Prabhat, and Michael Wehner. A parallel EM algorithm for model-based clustering applied to the exploration of large spatio-temporal data. Technometrics, 55(4):513–523, 2013. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Croux:2013:RSP

[CFF13] Christophe Croux, Peter Filzmoser, and Heinrich Fritz. Robust sparse principal component analysis. Technometrics, 55(2):202– 214, 2013. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Castruccio:2016:CES

[CG16] Stefano Castruccio and Marc G. Genton. Compressing an ensemble with statistical models: An algorithm for global 3D spatiotemporal temperature. Technometrics, 58(3):319–??, 2016. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Chevalier:2014:FPK

[CGB^{+}14] Clément Chevalier, David Ginsbourger, Julien Bect, Emmanuel Vazquez, Victor Picheny, and Yann Richet. Fast parallel krigingbased stepwise uncertainty reduction with application to the identification of an excursion set. Technometrics, 56(4):455–465, 2014. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Chakrabarty:2015:BED

[CGR^{+}15] Dalia Chakrabarty, Nare Gabrielyan, Fabio Rigat, Richard Beanland, and Shashi Paul. Bayesian estimation of density via multiple sequential inversions of two-dimensional images with application to electron microscopy. Technometrics, 57(2):217–233, 2015. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Chatterjee:2011:BRB

[Cha11] Snigdhansu Chatterjee. Book review: Random Networks for Communication by Massimo Franceschetti; Ronald Meester. Technometrics, 53(4):440, November 2011. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://

www.jstor.org/stable/41714956.

Chatterjee:2012:BRB

[Cha12] Snigdhansu Chatterjee. Book review: Evolutionary Statistical Procedures by Roberto Baragona; Francesco Battaglia; Irene Poli. Technometrics, 54(4):440, November 2012. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://

www.jstor.org/stable/41714928.

Chernick:2011:BRBb

[Che11a] Michael R. Chernick. Book review: Bayesian Adaptive Methods for Clinical Trials by Scott M. Berry; Bradley P. Carlin; J. Jack Lee; Peter Müller. Technometrics, 53(3):322–323, August 2011. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/23210407.

Chernick:2011:BRBc

[Che11b] Michael R. Chernick. Book review: Multiple Comparisons Using R by Frank Bretz; Torsten Hothorn; Peter Westfall. Technometrics, 53(3):327–328, August 2011. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/23210417.

Chernick:2011:BRBa

[Che11c] Michael R. Chernick. Book review: Statistical and Computational Pharmacogenomics by Rongling Wu; Min Lin. Technometrics, 53 (1):100, February 2011. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/40997299.

Chernick:2012:BRBa

[Che12a] Michael R. Chernick. Book review: Handbook of Markov Chain Monte Carlo by Steve Brooks; Andrew Gelman; Galin Jones; Xiao-Li Meng. Technometrics, 54(3):319–320, August 2012. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/41714899.

Chernick:2012:BRBb

[Che12b] Michael R. Chernick. Book review: Statistical Learning for Biomedical Data by James D. Malley; Karen G. Malley; Sinisa Pajevic. Technometrics, 54(4):441–442, November 2012. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/41714931.

Chipman:2011:ER

[Chi11] Hugh A. Chipman. Editor’s report. Technometrics, 53(4):336– 338, November 2011. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
41714947.

Chipman:2012:EA

[Chi12a] Hugh A. Chipman. Editorial announcement. Technometrics, 54 (4):333, 2012. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Chipman:2012:ER

[Chi12b] Hugh A. Chipman. Editor’s report. Technometrics, 54(4):337– 339, November 2012. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
41714918.

Chipman:2012:MHS

[Chi12c] Hugh A. Chipman. In memoriam: Harry Smith, Jr. Technometrics, 54(2):206, May 2012. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/41714882.

Chipman:2013:ER

[Chi13] Hugh A. Chipman. Editor’s report. Technometrics, 55(4):386– 388, 2013. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Cui:2010:PDF

[CHKC10] Yue Cui, James S. Hodges, Xiaoxiao Kong, and Bradley P. Carlin. Partitioning degrees of freedom in hierarchical and other richly parameterized models. Technometrics, 52(1):124–136, February 2010. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/40586685.

Chokshi:2011:AHA

[Cho11] Daksha Chokshi. Announcement of 2011 Hunter award. Technometrics, 53(2):111, 2011. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Chokshi:2012:AHA

[Cho12] Daksha Chokshi. Announcement of 2012 Hunter award. Technometrics, 54(2):107, 2012. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Charnigo:2011:GCD

[CHS11] Richard Charnigo, Benjamin Hall, and Cidambi Srinivasan. A generalized C_{p} criterion for derivative estimation. Technometrics, 53(3):238–253, August 2011. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/23210400.

Clemmensen:2011:SDA

[CHWE11] Line Clemmensen, Trevor Hastie, Daniela Witten, and Bjarne Ersbøll. Sparse discriminant analysis. Technometrics, 53(4):406– 413, November 2011. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
41714953.

Cinar:2011:BRB

[Cin11] Ali Cinar. Book review: Random Phenomena: Fundamentals of Probability and Statistics for Engineers by Babatunde A. Ogunnaike. Technometrics, 53(3):325–326, August 2011. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/23210412.

Chang:2014:MCT

[CJ14] Chia-Jung Chang and V. Roshan Joseph. Model calibration through minimal adjustments. Technometrics, 56(4):474–482, 2014. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Chen:2015:E

[CJDN15] Yong Chen, V. Roshan Joseph, Yu Ding, and Vijay Nair. Editorial. Technometrics, 57(3):303–304, 2015. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1080/00401706.2015.1040928.

Carpio:2014:MES

[CJL14] Jaime Carpio, Jesús Juan, and Damián López. Multivariate exponential smoothing and dynamic factor model applied to hourly electricity price analysis. Technometrics, 56(4):494–503, 2014. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Cheng:2016:CM

[CL16] Yichen Cheng and Faming Liang. Comment: ‘Modeling an Augmented Lagrangian for Blackbox Constrained Optimization’ by Gramacy et al. Technometrics, 58(1):15–17, 2016. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00401706.2015.1040927. See [GGL^{+}16a].

Capizzi:2011:LAR

[CM11] Giovanna Capizzi and Guido Masarotto. A least angle regression control chart for multidimensional data. Technometrics, 53 (3):285–296, August 2011. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/23210404.

Chen:2014:MCD

[CM14] Kehui Chen and Hans-Georg Müller. Modeling conditional distributions for functional responses, with application to traffic monitoring via GPS-enabled mobile phones. Technometrics, 56(3): 347–358, 2014. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Capizzi:2017:PDF

[CM17] Giovanna Capizzi and Guido Masarotto. Phase I distributionfree analysis of multivariate data. Technometrics, 59(4):484–495, 2017. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Chapman:2012:CEC

[CMAC12] Jessica L. Chapman, Max D. Morris, and Christine M. Anderson-Cook. Computationally efficient comparison of experimental designs for system reliability studies with binomial data. Technometrics, 54(4):410–424, November 2012. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://

www.jstor.org/stable/41714925.

Cano:2010:BRA

[CMR10] Javier Cano, Javier M. Moguerza, and David Ríos Insua. Bayesian reliability, availability, and maintainability analysis for hardware systems described through continuous time Markov chains. Technometrics, 52(3):324–334, August 2010. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/27867254.

Chipman:2010:NPM

[CMS10] Hugh A. Chipman, R. Jock MacKay, and Stefan H. Steiner. [Nonparametric Profile Monitoring by Mixed Effects Modeling]: Comment. Technometrics, 52(3):280–283, August 2010. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/27867247.

Coburn:2010:BRB

[Cob10] Timothy C. Coburn. Book review: Practical Geostatistics 2000 Series — Case Studies 2009 by Isobel Clark; William V. Harper. Technometrics, 52(3):367, August 2010. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://

www.jstor.org/stable/27867268.

Conover:2010:LE

[Con10] W. J. Conover. Letter to the Editor. Technometrics, 52(2):256– 257, May 2010. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
27867230.

Culp:2019:DIP

[CRBM19] Mark V. Culp, Kenneth J. Ryan, Prithish Banerjee, and Michael Morehead. On data integration problems with manifolds. Technometrics, 61(2):165–175, 2019. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Culp:2018:ARR

[CRCH18] Stacey L. Culp, Kenneth J. Ryan, Juan Chen, and Michael S. Hamada. Analysis of repeatability and reproducibility studies with ordinal measurements. Technometrics, 60(4):545–556, 2018. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Chang:2018:RLR

[CRW18] Le Chang, Steven Roberts, and Alan Welsh. Robust lasso regression using Tukey’s biweight criterion. Technometrics, 60(1): 36–47, 2018. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Crainiceanu:2012:CRC

[CS12] Ciprian M. Crainiceanu and Ana-Maria Staicu. [Clustering Random Curves Under Spatial Interdependence With Application to Service Accessibility]: Comment. Technometrics, 54(2):120– 122, May 2012. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/

41714856. See [JS12a].

Cook:2016:SPE

[CS16] R. Dennis Cook and Zhihua Su. Scaled predictor envelopes and partial least-squares regression. Technometrics, 58(2):155–165, 2016. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Chen:2018:ESE

[CSY18] Xin Chen, Wenhui Sheng, and Xiangrong Yin. Efficient sparse estimate of sufficient dimension reduction in high dimension. Technometrics, 60(2):161–168, 2018. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Chen:2016:CEI

[CW16] Hao Chen and William J. Welch. Comment: Expected improvement for efficient blackbox constrained optimization. Technometrics, 58(1):12–15, 2016. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/00401706.2015.1044119.

See [GGL^{+}16a].

Chen:2017:SDB

[CWW17] Ray-Bing Chen, Weichung Wang, and C. F. Jeff Wu. Sequential designs based on Bayesian uncertainty quantification in sparse representation surrogate modeling. Technometrics, 59(2):139– 152, 2017. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Chen:2017:EFR

[CY17] Piao Chen and Zhi-Sheng Ye. Estimation of field reliability based on aggregate lifetime data. Technometrics, 59(1):115–125, 2017. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Chen:2017:MDE

[CYZ17] Piao Chen, Zhi-Sheng Ye, and Xingqiu Zhao. Minimum distance estimation for the generalized Pareto distribution. Technometrics, 59(4):528–541, 2017. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Chen:2015:CSM

[CZ15a] Nan Chen and Shiyu Zhou. CUSUM statistical monitoring of M/M/1 queues and extensions. Technometrics, 57(2):245–256, 2015. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Cook:2015:SEM

[CZ15b] R. Dennis Cook and Xin Zhang. Simultaneous envelopes for multivariate linear regression. Technometrics, 57(1):11–25, 2015. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Chen:2016:DFM

[CZZ16] Nan Chen, Xuemin Zi, and Changliang Zou. A distributionfree multivariate control chart. Technometrics, 58(4):448–459, 2016. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Duan:2017:SFF

[DASS17] Weitao Duan, Bruce E. Ankenman, Susan M. Sanchez, and Paul J. Sanchez. Sliced full factorial-based latin hypercube designs as a framework for a batch sequential design algorithm. Technometrics, 59(1):11–22, 2017. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Davis:2013:CDA

[Dav13] Tim Davis. Comment: Dimensional analysis in statistical engineering. Technometrics, 55(3):271–274, 2013. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

delCastillo:2011:SSA

[dCC11] Enrique del Castillo and Bianca M. Colosimo. Statistical shape analysis of experiments for manufacturing processes. Technometrics, 53(1):1–15, February 2011. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/40997288.

delCastillo:2015:GGP

[dCCT15] Enrique del Castillo, Bianca M. Colosimo, and Sam Davanloo Tajbakhsh. Geodesic Gaussian processes for the parametric reconstruction of a free-form surface. Technometrics, 57(1):87–99, 2015. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Dean:2012:CRC

[DF12] C. B. Dean and Cindy X. Feng. [Clustering Random Curves Under Spatial Interdependence With Application to Service Accessibility]: Comment. Technometrics, 54(2):122–123, May 2012. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/41714857. See [JS12a].

Dias:2013:HMA

[DGS13] Ronaldo Dias, Nancy L. Garcia, and Alexandra M. Schmidt. A hierarchical model for aggregated functional data. Technometrics, 55(3):321–334, 2013. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Deng:2015:QMJ

[DJ15] Xinwei Deng and Ran Jin. QQ models: Joint modeling for quantitative and qualitative quality responses in manufacturing systems. Technometrics, 57(3):320–331, 2015. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic).  URL http://www.tandfonline.com/doi/abs/10.
1080/00401706.2015.1029079.

Deng:2017:AGP

[DLLR17] X. Deng, C. Devon Lin, K.-W. Liu, and R. K. Rowe. Additive Gaussian process for computer models with qualitative and quantitative factors. Technometrics, 59(3):283–292, 2017. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Dong:2017:QNM

[DLY^{+}17] Ling Dong, Xiaodong Li, Dan Yu, Hui Zhang, Zhong Zhang, Yanjun Qian, and Yu Ding. Quantifying nanoparticle mixing state to account for both location and size effects. Technometrics, 59(3):391–403, 2017. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Dasgupta:2012:CDD

[DM12] Tirthankar Dasgupta and Xiao-Li Meng. Comment: “Dolt and Do It Well” [MR2967968]. Technometrics, 54(3):227–231, August 2012. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/41714888.

deMast:2010:MER

[dMvW10] Jeroen de Mast and Wessel N. van Wieringen. Modeling and evaluating repeatability and reproducibility of ordinal classifications. Technometrics, 52(1):94–106, February 2010. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://

www.jstor.org/stable/40586683.

Dasgupta:2010:RDM

[DMW10] Tirthankar Dasgupta, Arden Miller, and C. F. Jeff Wu. Robust design of measurement systems. Technometrics, 52(1):80– 93, February 2010. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
40586682.

Du:2016:BIN

[DNV16] Chuanlong Du, Daniel J. Nordman, and Stephen B. Vardeman. Bayesian inference for a new class of distributions on equivalence classes of three-dimensional orientations with applications to materials science. Technometrics, 58(2):214–224, 2016. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Dette:2010:GLH

[DP10] Holger Dette and Andrey Pepelyshev. Generalized Latin hypercube design for computer experiments. Technometrics, 52(4): 421–429, November 2010. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/40997249.

Drignei:2010:FAC

[Dri10] Dorin Drignei. Functional ANOVA in computer models with time series output. Technometrics, 52(4):430–437, November 2010. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/40997250.

Damle:2017:GAA

[DS17] Anil Damle and Yuekai Sun. A geometric approach to archetypal analysis and nonnegative matrix factorization. Technometrics, 59 (3):361–370, 2017. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Draguljic:2012:NSF

[DSD12] Danel Draguljić, Thomas J. Santner, and Angela M. Dean. Noncollapsing space-filling designs for bounded nonrectangular regions. Technometrics, 54(2):169–178, May 2012. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/41714865.

Danila:2013:ABM

[DSM13] Oana Danila, Stefan H. Steiner, and R. Jock MacKay. Assessing a binary measurement system with varying misclassification rates when a gold standard is available. Technometrics, 55(3):335–345, 2013. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Draguljic:2014:SSP

[DWD^{+}14] Danel Draguljić, David C. Woods, Angela M. Dean, Susan M. Lewis, and Anna-Jane E. Vine. Screening strategies in the presence of interactions. Technometrics, 56(1):1–16, 2014. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). See comment [ACH14].

Editor:2019:TER

[E19] Daniel Apley (Technometrics Editor and 2017-2019). Technometrics 2018 Editor’s report. Technometrics, 61(1):2–6, 2019. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Editor:2015:EA

[Edi15] Peihua Qiu Editor. Editorial announcement. Technometrics, 57(4):445, 2015. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/00401706.2015.1096586.

Espinosa:2016:BPA

[EDR16] Valeria Espinosa, Tirthankar Dasgupta, and Donald B. Rubin. A Bayesian perspective on the analysis of unreplicated factorial experiments using potential outcomes. Technometrics, 58(1):62– 73, 2016. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00401706.2015.1006337.

Edwards:2011:OSP

[Edw11] David J. Edwards. Optimal semifoldover plans for two-level orthogonal designs. Technometrics, 53(3):274–284, August 2011. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/23210403.

Einmahl:2012:HHP

[EG12] John H. J. Einmahl and Maria Gantner. The half-half plot. Technometrics, 54(2):138–146, May 2012. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://

www.jstor.org/stable/41714862.

Errore:2017:BFC

[EJLN17] Anna Errore, Bradley Jones, William Li, and Christopher J. Nachtsheim. Benefits and fast construction of efficient two-level foldover designs. Technometrics, 59(1):48–57, 2017. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Ezzat:2018:SDF

[EPD18] Ahmed Aziz Ezzat, Arash Pourhabib, and Yu Ding. Sequential design for functional calibration of computer models. Technometrics, 60(3):286–296, 2018. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Eendebak:2017:TLD

[ES17] Pieter T. Eendebak and Eric D. Schoen. Two-level designs to estimate all main effects and two-factor interactions. Technometrics, 59(1):69–79, 2017. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Espejo:2011:BRB

[Esp11] Mariano Ruiz Espejo. Book review: Indirect Sampling by Pierre Lavallée. Technometrics, 53(3):327, August 2011. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/23210416.

Emura:2010:ATL

[EW10] Takeshi Emura and Hsiuying Wang. Approximate tolerance limits under log-location-scale regression models in the presence of censoring. Technometrics, 52(3):313–323, August 2010. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/27867253.

Farcomeni:2014:RCC

[Far14] Alessio Farcomeni. Robust constrained clustering in presence of entry-wise outliers. Technometrics, 56(1):102–111, 2014. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Flaherty:2015:ROB

[FD15] Patrick Flaherty and Ronald W. Davis. Robust optimization of biological protocols. Technometrics, 57(2):234–244, 2015. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Farah:2014:BIS

[FK14] Marian Farah and Athanasios Kottas. Bayesian inference for sensitivity analysis of computer simulators, with an application to radiative transfer models. Technometrics, 56(2):159–173, 2014. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Fielding:2011:EMS

[FNL11] Mark Fielding, David J. Nott, and Shie-Yui Liong. Efficient MCMC schemes for computationally expensive posterior distributions. Technometrics, 53(1):16–28, February 2011. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/40997289.

Forrester:2013:CPP

[For13] Alexander I. J. Forrester. Comment: Properties and practicalities of the expected quantile improvement. Technometrics, 55(1):13– 18, 2013. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Fotopoulos:2010:BRBb

[Fot10a] Stergios B. Fotopoulos. Book review: Markov Processes, Gaussian Processes and Local Times by Michael B. Marcus; Jay Rosen. Technometrics, 52(3):366, August 2010. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://

www.jstor.org/stable/27867266.

Fotopoulos:2010:BRBa

[Fot10b] Stergios B. Fotopoulos. Book review: Stereology for Statisticians. Monograph on Statistics and Applied Probability, Vol. 103 by Adrian Baddeley; Eva B. Vedel Jensen. Technometrics, 52 (3):362–363, August 2010. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/27867258.

Fricker:2013:MGP

[FOU13] Thomas E. Fricker, Jeremy E. Oakley, and Nathan M. Urban. Multivariate Gaussian process emulators with nonseparable covariance structures. Technometrics, 55(1):47–56, 2013. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Fokianos:2017:CTP

[FP17] K. Fokianos and M. Pitsillou. Consistent testing for pairwise dependence in time series. Technometrics, 59(2):262–270, 2017. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Fang:2019:IBP

[FPG19] Xiaolei Fang, Kamran Paynabar, and Nagi Gebraeel. Imagebased prognostics using penalized tensor regression. Technometrics, 61(3):369–384, 2019. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Feng:2018:DDB

[FQ18] Long Feng and Peihua Qiu. Difference detection between two images for image monitoring. Technometrics, 60(3):345–359, 2018. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Franck:2019:DHA

[Fra19] Christopher T. Franck. Detection of hidden additivity and inference under model uncertainty for unreplicated factorial studies via Bayesian model selection and averaging. Technometrics, 61 (3):283–296, 2019. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Frey:2012:BRB

[Fre12] Michael Frey. Book review: Statistical Analysis of Network Data by Eric D. Kolaczyk. Technometrics, 54(1):97–98, February 2012. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/41714835.

Frey:2013:CDA

[Fre13] Daniel D. Frey. Comment: Dimensional analysis and experimentation as a catalyst to learning from data. Technometrics, 55 (3):275–278, 2013. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Graves:2010:RMA

[GACH10] Todd L. Graves, Christine M. Anderson-Cook, and Michael S. Hamada. Reliability models for almost-series and almost-parallel systems. Technometrics, 52(2):160–171, May 2010. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/27867221.

Gao:2015:CDM

[GAZ15] Qibing Gao, Mihye Ahn, and Hongtu Zhu. Cook’s distance measures for varying coefficient models with functional responses. Technometrics, 57(2):268–280, 2015. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Guhaniyogi:2018:MKS

[GB18] Rajarshi Guhaniyogi and Sudipto Banerjee. Meta-kriging: Scalable Bayesian modeling and inference for massive spatial datasets. Technometrics, 60(4):430–444, 2018. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Goh:2013:PCM

[GBH^{+}13] Joslin Goh, Derek Bingham, James Paul Holloway, Michael J. Grosskopf, Carolyn C. Kuranz, and Erica Rutter. Prediction and computer model calibration using outputs from multifidelity simulators. Technometrics, 55(4):501–512, 2013. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Gervini:2015:DRR

[Ger15] Daniel Gervini. Dynamic retrospective regression for functional data. Technometrics, 57(1):26–34, 2015. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Gu:2018:AHD

[GFK^{+}18] Yuwen Gu, Jun Fan, Lingchen Kong, Shiqian Ma, and Hui Zou. ADMM for high-dimensional sparse penalized quantile regression. Technometrics, 60(3):319–331, 2018. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Goos:2012:GSA

[GG12] Peter Goos and Steven G. Gilmour. A general strategy for analyzing data from split-plot and multistratum experimental designs. Technometrics, 54(4):340–354, November 2012. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/41714919.

Gramacy:2016:MAL

[GGL^{+}16a] Robert B. Gramacy, Genetha A. Gray, Sébastien Le Digabel, Herbert K. H. Lee, Pritam Ranjan, Garth Wells, and Stefan M. Wild. Modeling an augmented Lagrangian for blackbox constrained optimization. Technometrics, 58(1):1–11, 2016. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00401706.2015.1014065. See comments [CW16, CL16,

PGK16, HWL16] and rejoinder [GGL^{+}16b].

Gramacy:2016:R

[GGL^{+}16b] Robert B. Gramacy, Genetha A. Gray, Sébastien Le Digabel, Herbert K. H. Lee, Pritam Ranjan, Garth Wells, and Stefan M. Wild. Rejoinder. Technometrics, 58(1):26–29, 2016. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00401706.2015.1106979. See [GGL^{+}16a].

Giorgio:2010:SDW

[GGP10] Massimiliano Giorgio, Maurizio Guida, and Gianpaolo Pulcini. A state-dependent wear model with an application to marine engine cylinder liners. Technometrics, 52(2):172–187, May 2010. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/27867222.

Gramacy:2016:SNS

[GH16] Robert B. Gramacy and Benjamin Haaland. Speeding up neighborhood search in local Gaussian process prediction. Technometrics, 58(3):294–??, 2016. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Ghosh:2010:BRBb

[Gho10a] Subir Ghosh. Book review: Analyzing Receiver Operating Characteristic Curves With SAS by Mithat Gonena. Technometrics, 52(3):365, August 2010. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/27867264.

Ghosh:2010:BRBa

[Gho10b] Subir Ghosh. Book review: Optimal Statistical Inference in Financial Engineering by Masanobu Taniguchi; Junichi Hirukawa; Kenichiro Tamaki. Technometrics, 52(2):261, May 2010. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/27867235.

Ghosh:2011:BRBa

[Gho11a] Subir Ghosh. Book review: Comparing Distributions by Olivier Thas. Technometrics, 53(1):99, February 2011. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/40997297.

Ghosh:2011:BRBb

[Gho11b] Subir Ghosh. Book review: Linear Mixed Models for Longitudinal Data by Geert Verbeke; Geert Molenberghs. Technometrics, 53(3):324, August 2011. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/23210410.

Ghosh:2012:BRBa

[Gho12a] Subir Ghosh. Book review: Antedependence Models for Longitudinal Data by Dale L. Zimmerman; Vicente A. Núñez-Antón. Technometrics, 54(2):198–199, May 2012. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://

www.jstor.org/stable/41714868.

Ghosh:2012:BRBb

[Gho12b] Subir Ghosh. Book review: Expansions and Asymptotics for Statistics by Christopher G. Small. Technometrics, 54(2):199, May 2012. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/41714869.

Ghosh:2019:BRL

[Gho19] Subir Ghosh. Book review: Lectures on Categorical Data Analysis. Technometrics, 61(4):560, 2019. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Gaver:2010:RGM

[GJ10] Donald P. Gaver and Patricia A. Jacobs. [Reliability Growth Management Metrics and Statistical Methods for Discrete-Use Systems]: Comment. Technometrics, 52(4):389–391, November 2010. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/40997242.

Goos:2019:OED

[GJ19] Peter Goos and Bradley Jones. Optimal experimental design in the presence of nested factors. Technometrics, 61(4):533–544, 2019. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Galeano:2015:MDF

[GJL15] Pedro Galeano, Esdras Joseph, and Rosa E. Lillo. The Mahalanobis distance for functional data with applications to classification. Technometrics, 57(2):281–291, 2015. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Gramacy:2012:GPS

[GL12] Robert B. Gramacy and Heng Lian. Gaussian process singleindex models as emulators for computer experiments. Technometrics, 54(1):30–41, February 2012. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/41714829.

Grego:2014:PBP

[GLLS14] John Grego, Shuang Li, James Lynch, and Jayaram Sethuraman. Partition-based priors and multiple event censoring: An analysis of Rosen’s fibrous composite experiment. Technometrics, 56(3): 359–371, 2014. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Goldstein:2011:BRBb

[Gol11a] Richard Goldstein. Book review: Bias and Causation: Models and Judgment for Valid Comparisons by Herbert I. Weisberg. Technometrics, 53(3):323–324, August 2011. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://

www.jstor.org/stable/23210409.

Goldstein:2011:BRBa

[Gol11b] Richard Goldstein. Book review: Design of Observational Studies by Paul R. Rosenbaum. Technometrics, 53(2):218–219, May 2011. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/23209745.

Goldstein:2011:BSU

[Gol11c] Richard Goldstein. R for Stata Users by Robert A. Muenchen; Joseph M. Hilbe. Technometrics, 53(4):442, November 2011. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/41714959.

Goldstein:2012:BRBc

[Gol12a] Richard Goldstein. Book review: Adaptive and Flexible Clinical Trials by Richard Chin. Technometrics, 54(4):443, November 2012. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/41714933.

Goldstein:2012:BRBb

[Gol12b] Richard Goldstein. Book review: Controversial Statistical Issues in Clinical Trials by Shein-Chung Chow. Technometrics, 54 (3):322–323, August 2012. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/41714902.

Goldstein:2012:BRBa

[Gol12c] Richard Goldstein. Book review: Dynamic Mixed Models for Familial Longitudinal Data by Brajendra C. Sutradhar. Technometrics, 54(2):199–200, May 2012. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/41714871.

Gotwalt:2010:ASC

[Got10a] Christopher M. Gotwalt. Addendum to “Fast Computation of Designs Robust to Parameter Uncertainty for Nonlinear Settings”.

Technometrics, 52(1):137, 2010. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Gotwalt:2010:AFC

[Got10b] Christopher M. Gotwalt. Addendum to “Fast Computation of Designs Robust to Parameter Uncertainty for Nonlinear Settings” [MR2664866]. Technometrics, 52(1):137, February 2010. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/40586686.

Gramacy:2013:CAE

[Gra13] Robert B. Gramacy. Comment: On advances in expected improvement. Technometrics, 55(1):19–20, 2013. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Gromping:2014:NDF

[Grö14] Ulrike Grömping. A note on dominating fractional factorial twolevel designs with clear two-factor interactions. Technometrics, 56(1):42–45, 2014. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Gruber:2011:BRB

[Gru11] Marvin H. J. Gruber. Book review: A Comparison of the Bayesian and Frequentist Approaches to Estimation by Francisco J. Samaniego. Technometrics, 53(3):326–327, August 2011. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/23210414.

Guinness:2018:PGM

[Gui18] Joseph Guinness. Permutation and grouping methods for sharpening Gaussian process approximations. Technometrics, 60(4): 415–429, 2018. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Guo:2019:BRB

[Guo19] Jian Guo. Book review: Bayesian Inference-Data Evaluation and Decisions (2nd ed.). Technometrics, 61(4):561–562, 2019. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Guo:2013:BME

[GW13] Jiqiang Guo and Alyson G. Wilson. Bayesian methods for estimating system reliability using heterogeneous multilevel information. Technometrics, 55(4):461–472, 2013. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Guo:2013:BNM

[GWN13] Jiqiang Guo, Alyson G. Wilson, and Daniel J. Nordman. Bayesian nonparametric models for community detection. Technometrics, 55(4):390–402, 2013. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Hand:2010:FDT

[Han10] David J. Hand. Fraud detection in telecommunications and banking: Discussion of Becker, Volinsky, and Wilks (2010) and Sudjianto et al. (2010) [MR2752106; MR2752105]. Technometrics, 52 (1):34–38, February 2010. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/ stable/40586678. See [BVW10, SNY^{+}10].

Hazelton:2010:SIT

[Haz10] Martin L. Hazelton. Statistical inference for transit system origindestination matrices. Technometrics, 52(2):221–230, May 2010. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/27867226.

Heaton:2017:NGP

[HCT17] Matthew J. Heaton, William F. Christensen, and Maria A. Terres. Nonstationary Gaussian process models using spatial hierarchical clustering from finite differences. Technometrics, 59(1): 93–101, 2017. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Hong:2015:SMD

[HDM^{+}15] Yili Hong, Yuanyuan Duan, William Q. Meeker, Deborah L. Stanley, and Xiaohong Gu. Statistical methods for degradation data with dynamic covariates information and an application to outdoor weathering data. Technometrics, 57(2):180–193, 2015. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

He:2019:SRS

[He19] Xu He. Sliced rotated sphere packing designs. Technometrics, 61(1):66–76, 2019. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Heckler:2010:BRB

[Hec10] Charles Heckler. Book review: Medical Applications of Finite Mixture Models by Peter Schlattmann. Technometrics, 52(4):463– 464, November 2010. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
40997257.

Heckler:2011:BRB

[Hec11] Charles E. Heckler. Book review: SAS and R: Data Management, Statistical Analysis and Graphics by Ken Kleinman; Nicholas J. Horton. Technometrics, 53(2):221, May 2011. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://

www.jstor.org/stable/23209749.

Hall:2010:RGMa

[HEM10a] J. Brian Hall, Paul M. Ellner, and Ali Mosleh. Reliability growth management metrics and statistical methods for discreteuse systems. Technometrics, 52(4):379–389, November 2010. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/40997241. With supple-

mentary material available online.

Hall:2010:RGMb

[HEM10b] J. Brian Hall, Paul M. Ellner, and Ali Mosleh. [Reliability Growth Management Metrics and Statistical Methods for Discrete-Use Systems]: Rejoinder. Technometrics, 52(4):401–408, November 2010. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/40997247.

Heydorn:2012:BRB

[Hey12] Richard P. Heydorn. Book review: A First Course in Bayesian Statistical Methods by Peter D. Hoff. Technometrics, 54(1):98– 99, February 2012. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
41714836.

Hering:2011:CSP

[HG11] Amanda S. Hering and Marc G. Genton. Comparing spatial predictions. Technometrics, 53(4):414–425, November 2011. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/41714954.

Higdon:2013:CMC

[HGL^{+}13] Dave Higdon, Jim Gattiker, Earl Lawrence, Charles Jackson, Michael Tobis, Matt Pratola, Salman Habib, Katrin Heitmann, and Steve Price. Computer model calibration using the ensemble Kalman filter. Technometrics, 55(4):488–500, 2013. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Hinich:2011:BRB

[Hin11] Melvin Hinich. Book review: Computational Statistics by James Gentle. Technometrics, 53(1):104, February 2011. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/40997306.

Hung:2015:ACE

[HJM15] Ying Hung, V. Roshan Joseph, and Shreyes N. Melkote. Analysis of computer experiments with functional response. Technometrics, 57(1):35–44, 2015. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Henrey:2017:JIL

[HL17] Andrew J. Henrey and Thomas M. Loughin. Joint identification of location and dispersion effects in unreplicated two-level factorials. Technometrics, 59(1):23–35, 2017. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Hlavacek:2015:NSW

[Hla15] Douglas Hlavacek. Nominations sought for William G. Hunter Award. Technometrics, 57(2):301, 2015. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Hu:2019:ATP

[HLBS19] Wenhao Hu, Eric B. Laber, Clay Barker, and Leonard A. Stefanski. Assessing tuning parameter selection variability in penalized regression. Technometrics, 61(2):154–164, 2019. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Huang:2016:CEB

[HLLY16] Hengzhen Huang, Dennis K. J. Lin, Min-Qian Liu, and Jian-Feng Yang. Computer experiments with both qualitative and quantitative variables. Technometrics, 58(4):495–507, 2016. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Hall:2013:NAF

[HLP13] Peter Hall, Fred Lombard, and Cornelis J. Potgieter. A new approach to function-based hypothesis testing in location-scale families. Technometrics, 55(2):215–223, 2013. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Hoegh:2015:BMF

[HLSR15] Andrew Hoegh, Scotland Leman, Parang Saraf, and Naren Ramakrishnan. Bayesian model fusion for forecasting civil unrest. Technometrics, 57(3):332–340, 2015. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/00401706.2014.1001522.

Hlynka:2010:BRBa

[Hly10a] Myron Hlynka. Book review: An Introduction to Queuing Theory Modeling and Analysis in Applications by U. Narayan Bhat. Technometrics, 52(1):138–139, February 2010. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://

www.jstor.org/stable/40586687.

Hlynka:2010:BRBb

[Hly10b] Myron Hlynka. Book review: Probability and Statistics by Example. Volume 2: Markov Chains: A Primer in Random Processes and Their Applications by Yuri Suhov; Mark Kelbert. Technometrics, 52(2):263, May 2010. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/27867238.

Hlynka:2011:BRB

[Hly11] Myron Hlynka. Book review: Stochastic Processes (3rd ed.) by J. Medhi. Technometrics, 53(2):221–222, May 2011. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/23209751.

Hlynka:2012:BRB

[Hly12] Myron Hlynka. Book review: Reversibility and Stochastic Networks by F. P. Kelly. Technometrics, 54(2):204, May 2012. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/41714879.

Hong:2010:FFW

[HM10] Yili Hong and William Q. Meeker. Field-failure and warranty prediction based on auxiliary use-rate information. Technometrics, 52(2):148–159, May 2010. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/27867220.

Hong:2013:FFP

[HM13] Yili Hong and William Q. Meeker. Field-failure predictions based on failure-time data with dynamic covariate information. Technometrics, 55(2):135–149, 2013. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Hannart:2012:IBI

[HN12] Alexis Hannart and Philippe Naveau. An improved Bayesian information criterion for multiple change-point models. Technometrics, 54(3):256–268, August 2012. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/41714894.

Hernandez:2018:FCE

[HN18] Lucia N. Hernandez and Christopher J. Nachtsheim. Fast computation of exact G-optimal designs via I_{λ}-optimality. Technometrics, 60(3):297–305, 2018. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Hamilton:2018:CVW

[HNKF18] Jean Hamilton, Matthew A. Nunes, Marina I. Knight, and Piotr Fryzlewicz. Complex-valued wavelet lifting and applications. Technometrics, 60(1):48–60, 2018. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Hoffman:2010:OST

[Hof10] David Hoffman. One-sided tolerance limits for balanced and unbalanced random effects models. Technometrics, 52(3):303– 312, August 2010. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
27867252.

Hornikova:2010:BRBd

[Hor10a] Adriana Horníková. Book review: Analysis, Geometry, and Modeling in Finance — Advanced Methods in Option Pricing by Pierre Henry-Labordère. Technometrics, 52(3):366–367, August 2010. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/27867267.

Hornikova:2010:BRBa

[Hor10b] Adriana Horníková. Book review: Making Sense of Data II: A Practical Guide to Data Visualization, Advanced Data Mining Methods, and Applications by Glenn J. Myatt; Wayne P. Johnson. Technometrics, 52(1):141–142, February 2010. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/40586691.

Hornikova:2010:BRBc

[Hor10c] Adriana Horníková. Book review: Stochastic Control of Hereditary Systems and Applications by Mou-Hsiung Chang. Technometrics, 52(2):260–261, May 2010. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/27867234.

Hornikova:2010:BRBb

[Hor10d] Adriana Horníková. Book review: The Methodology and Practice of Econometrics, A Festschrift in Honour of David F. Hendry by Jennifer L. Castle; Neil Shephard. Technometrics, 52(1):142– 143, February 2010. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
40586693.

Hornikova:2011:BRBe

[Hor11a] Adriana Horníková. Book review: Analysing Seasonal Health Data (Series for Biology and Health) by Adrian G. Barnett; Annette J. Dobson. Technometrics, 53(4):441–442, November 2011. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/41714958.

Hornikova:2011:BRBc

[Hor11b] Adriana Horníková. Book review: Lectures on Algebraic Statistics. Oberwolfach Seminars, Vol. 39 by Mathias Drton; Bernd Sturmfels; Seth Sullivant. Technometrics, 53(3):323, August 2011. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/23210408.

Hornikova:2011:BRBd

[Hor11c] Adriana Horníková. Book review: Statistical Machine Translation by Philipp Koehn. Technometrics, 53(3):328–329, August 2011. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/23210419.

Hornikova:2011:BRBb

[Hor11d] Adriana Hornikova. Book review: Statistical Methods for Disease Clustering by Toshiro Tango. Technometrics, 53(2):215–216, May 2011. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/23209740.

Hornikova:2011:BRBf

[Hor11e] Adriana Horníková. Book review: Statistics for Sensory and Consumer Science by Tormod Naes; Per B. Brockhoff; Oliver Tomic. Technometrics, 53(4):445–446, November 2011. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/41714963.

Hornikova:2011:BRBa

[Hor11f] Adriana Horníková. Book review: Support Vector Machines by Ingo Stewart; Andreas Christman. Technometrics, 53(2):210–211, May 2011. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/23209734.

Hornikova:2012:BRBd

[Hor12a] Adriana Horníková. Book review: A Primer of Ecology With R by Martin Henry H. Stevens. Technometrics, 54(4):439–440, November 2012. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
41714927.

Hornikova:2012:BRBb

[Hor12b] Adriana Horníková. Book review: Quantitative Risk Assessment (The Scientific Platform) by Terje Aven. Technometrics, 54(1): 102–103, February 2012. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/41714844.

Hornikova:2012:BRBa

[Hor12c] Adriana Horníková. Book review: Selected Works in Probability and Statistics — Works of Donald L. Burkholder by Burgess Davis; Renming Song. Technometrics, 54(1):102, February 2012. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/41714843.

Hornikova:2012:BRBc

[Hor12d] Adriana Horníková. Book review: Selected Works of M. Rosenblatt by Richard A. Davis; Keh-Shin Lii; Dimitris N. Politis. Technometrics, 54(1):104–105, February 2012. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://

www.jstor.org/stable/41714847.

Holmstrom:2012:BSS

[HP12] Lasse Holmström and Leena Pasanen. Bayesian scale space analysis of differences in images. Technometrics, 54(1):16–29, February 2012. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/41714828.

Hwang:2016:SOA

[HQH16] Youngdeok Hwang, Peter Z. G. Qian, and Xu He. Sliced orthogonal array-based Latin hypercube designs. Technometrics, 58 (1):50–61, 2016. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/00401706.2014.993092.

Heaton:2010:ITD

[HRC10] Matthew J. Heaton, C. Shane Reese, and William F. Christensen. Incorporating time-dependent source profiles using the Dirichlet distribution in multivariate receptor models. Technometrics, 52 (1):67–79, February 2010. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/40586681.

Hubert:2016:SPH

[HRSV16] Mia Hubert, Tom Reynkens, Eric Schmitt, and Tim Verdonck. Sparse PCA for high-dimensional data with outliers. Technometrics, 58(4):424–434, 2016. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Hubert:2019:MAO

[HRV19] Mia Hubert, Peter J. Rousseeuw, and Wannes Van den Bossche. MacroPCA: An all-in-one PCA method allowing for missing values as well as cellwise and rowwise outliers. Technometrics, 61 (4):459–473, 2019. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Harari:2014:CCG

[HS14] Ofir Harari and David M. Steinberg. Convex combination of Gaussian processes for Bayesian analysis of deterministic computer experiments. Technometrics, 56(4):443–454, 2014. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Huo:2016:FCD

[HS16] Xiaoming Huo and Gábor J. Székely. Fast computing for distance covariance. Technometrics, 58(4):435–447, 2016. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). See clarifications [HS17].

Huo:2017:CFC

[HS17] Xiaoming Huo and Gábor J. Székely. Clarifications for “Fast Computing for Distance Covariance” by Xiaoming Huo and Gábor J. Székely. Technometrics, 59(1):134–135, 2017. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). See [HS16].

Huang:2019:DTV

[HS19] Huang Huang and Ying Sun. A decomposition of total variation depth for understanding functional outliers. Technometrics, 61 (4):445–458, 2019. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Holsclaw:2013:GPM

[HSL^{+}13] Tracy Holsclaw, Bruno Sansó, Herbert K. H. Lee, Katrin Heitmann, Salman Habib, David Higdon, and Ujjaini Alam. Gaussian process modeling of derivative curves. Technometrics, 55(1):57– 67, 2013. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

He:2017:OMF

[HTW17] Xu He, Rui Tuo, and C. F. Jeff Wu. Optimization of multi-fidelity computer experiments via the EQIE criterion. Technometrics, 59 (1):58–68, 2017. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Hunter:2013:MGP

[Hun13] J. Stuart Hunter. In memoriam: George E. P. Box. Technometrics, 55(2):119–120, 2013. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Huzurbazar:2010:ICF

[HW10] Aparna V. Huzurbazar and Brian J. Williams. Incorporating covariates in flowgraph models: Applications to recurrent event data. Technometrics, 52(2):198–208, May 2010. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/27867224.

Hare:2016:CNS

[HWL16] Warren Hare, Brian J. Williams, and Jason L. Loeppky. Comment: The NoMax strategy and correlated outputs. Technometrics, 58(1):22–25, 2016. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/00401706.2015.1077163.

See [GGL^{+}16a].

Ing:2017:MTR

[ILS^{+}17] Ching-Kang Ing, Tze Leung Lai, Milan Shen, KaWai Tsang, and Shu-Hui Yu. Multiple testing in regression models with applications to fault diagnosis in the big data era. Technometrics, 59 (3):351–360, 2017. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Jeon:2015:NKA

[JAP15] Yongho Jeon, Jeongyoun Ahn, and Cheolwoo Park. A nonparametric kernel approach to interval-valued data analysis. Technometrics, 57(4):566–575, 2015. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/00401706.2014.965346.

Jarrett:2012:BRBb

[Jar12a] Jeffrey E. Jarrett. Book review: Extreme Value Methods With Applications to Finance by Serguei Y. Novak. Technometrics, 54(3):327, August 2012. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/41714908.

Jarrett:2012:BRBa

[Jar12b] Jeffrey E. Jarrett. Book review: Statistical Tools for Finance and Insurance, (2nd ed.) by P. Cizek; W. K. Hardle; R. Weron. Technometrics, 54(1):102, February 2012. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://

www.jstor.org/stable/41714842.

Jiang:2016:OUC

[JDLH16] Huijing Jiang, Xinwei Deng, Vanessa López, and Hendrik F. Hamann. Online updating of computer model output using realtime sensor data. Technometrics, 58(4):472–482, 2016. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Joseph:2015:SEC

[JDTW15] V. Roshan Joseph, Tirthankar Dasgupta, Rui Tuo, and C. F. Jeff Wu. Sequential exploration of complex surfaces using minimum energy designs. Technometrics, 57(1):64–74, 2015. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Jensen:2010:BRB

[Jen10] Willis A. Jensen. Book review: Basic Concepts of Probability and Statistics in the Law by Michael O. Finkelstein. Technometrics, 52(4):465, November 2010. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/40997260.

Jensen:2011:BRB

[Jen11] Willis A. Jensen. Book review: Frontiers in Statistical Quality Control #9 by H. J. Lenz; P. H. Wilrich; W. Schmid. Technometrics, 53(2):220, May 2011. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/23209748.

Jones-Farmer:2014:ACC

[JFEH14] L. Allison Jones-Farmer, Jeremy D. Ezell, and Benjamin T. Hazen. Applying control chart methods to enhance data quality. Technometrics, 56(1):29–41, 2014. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Jhuang:2019:SSD

[JFJ^{+}19] An-Ting Jhuang, Montserrat Fuentes, Jacob L. Jones, Giovanni Esteves, Chris M. Fancher, Marschall Furman, and Brian J. Reich. Spatial signal detection using continuous shrinkage priors. Technometrics, 61(4):494–506, 2019. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Joseph:2019:SFD

[JGBM19] V. Roshan Joseph, Li Gu, Shan Ba, and William R. Myers. Spacefilling designs for robustness experiments. Technometrics, 61(1): 24–37, 2019. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Joseph:2011:RBI

[JK11] V. Roshan Joseph and Lulu Kang. Regression-based inverse distance weighting with applications to computer experiments. Technometrics, 53(3):254–265, August 2011. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://

www.jstor.org/stable/23210401.

Jala:2016:SDC

[JLLM^{+}16] Marjorie Jala, Céline Levy-Leduc, Éric Moulines, Emmanuelle Conil, and Joe Wiart. Sequential design of computer experiments for the assessment of fetus exposure to electromagnetic fields. Technometrics, 58(1):30–42, 2016. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/00401706.2014.979951.

Jones:2011:EDM

[JN11] Bradley Jones and Christopher J. Nachtsheim. Efficient designs with minimal aliasing. Technometrics, 53(1):62–71, February 2011. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/40997293.

Jones:2016:BSD

[JN16] Bradley Jones and Christopher J. Nachtsheim. Blocking schemes for definitive screening designs. Technometrics, 58(1):74–83, 2016. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00401706.2015.1013777.

Jones:2017:EDB

[JN17] Bradley Jones and Christopher J. Nachtsheim. Effective designbased model selection for definitive screening designs. Technometrics, 59(3):319–329, 2017. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Jones:2013:CES

[Jon13] Bradley Jones. Comment: Enhancing the search for compromise designs. Technometrics, 55(3):278–280, 2013. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Joseph:2012:BCU

[Jos12a] V. Roshan Joseph. Bayesian computation using design of experiments-based interpolation technique. Technometrics, 54 (3):209–225, August 2012. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/41714886.

Joseph:2012:CDS

[Jos12b] V. Roshan Joseph. [Comment: DoIt — Some Thoughts on How to Do It]: Rejoinder. Technometrics, 54(3):239–242, August 2012. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/41714892.

Joseph:2013:NND

[Jos13] V. Roshan Joseph. A note on nonnegative DoIt approximation. Technometrics, 55(1):103–107, 2013. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Jun:2013:MRM

[JP13] Mikyoung Jun and Eun Sug Park. Multivariate receptor models for spatially correlated multipollutant data. Technometrics, 55 (3):309–320, 2013. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Jiang:2012:CRCa

[JS12a] Huijing Jiang and Nicoleta Serban. Clustering random curves under spatial interdependence with application to service accessibility. Technometrics, 54(2):108–119, May 2012. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http:/

/www.jstor.org/stable/41714855. See [CS12, DF12, JSQ12,

LW12, WSZ12] and rejoinder [JS12b].

Jiang:2012:CRCb

[JS12b] Huijing Jiang and Nicoleta Serban. [Clustering Random Curves Under Spatial Interdependence With Application to Service Accessibility]: Rejoinder. Technometrics, 54(2):134–137, May 2012. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/41714861. See

[JS12a].

Jones:2015:BDM

[JSMS15] Bradley Jones, Rachel T. Silvestrini, Douglas C. Montgomery, and David M. Steinberg. Bridge designs for modeling systems with low noise. Technometrics, 57(2):155–163, 2015. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

James:2012:CRC

[JSQ12] Gareth M. James, Wenguang Sun, and Xinghao Qiao. [Clustering Random Curves Under Spatial Interdependence With Application to Service Accessibility]: Comment. Technometrics, 54(2): 123–126, May 2012. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/

41714858. See [JS12a].

Jiang:2019:REU

[JWFW19] Yunlu Jiang, You-Gan Wang, Liya Fu, and Xueqin Wang. Robust estimation using modified Huber’s functions with new tails. Technometrics, 61(1):111–122, 2019. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Joseph:2019:DSE

[JWG^{+}19] V. Roshan Joseph, Dianpeng Wang, Li Gu, Shiji Lyu, and Rui Tuo. Deterministic sampling of expensive posteriors using minimum energy designs. Technometrics, 61(3):297–308, 2019. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Joseph:2015:EDS

[JY15] V. Roshan Joseph and Huan Yan. Engineering-driven statistical adjustment and calibration. Technometrics, 57(2):257–267, 2015. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Katsaounis:2011:BRB

[Kat11] Tena I. Katsaounis. Book review: Mixed Effects Model for Complex Data by Lang Wu. Technometrics, 53(4):444–445, November 2011. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/41714962.

Kong:2018:DSF

[KAT18] Xiangshun Kong, Mingyao Ai, and Kwok Leung Tsui. Design for sequential follow-up experiments in computer emulations. Technometrics, 60(1):61–69, 2018. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Ke:2010:BRB

[Ke10] Weiming Ke. Book review: Automated Data Analysis Using Excel by Brian Bissett. Technometrics, 52(1):142, February 2010. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/40586692.

Ke:2011:BRB

[Ke11a] Weiming Ke. Book review: SAS for Data Analysis — Intermediate Statistical Methods by Mervyn Marasinghe; William Kennedy. Technometrics, 53(1):99–100, February 2011. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://

www.jstor.org/stable/40997298.

Ke:2011:PSE

[Ke11b] Weiming Ke. PROC SQL by example — using SQL within SAS by Howard Schreier. Technometrics, 53(1):105, February 2011. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/40997307.

Kazor:2019:MRM

[KH19] Karen Kazor and Amanda S. Hering. Mixture of regression models for large spatial datasets. Technometrics, 61(4):507–523, 2019. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Kang:2011:DMS

[KJB11] Lulu Kang, V. Roshan Joseph, and William A. Brenneman. Design and modeling strategies for mixture-of-mixtures experiments. Technometrics, 53(2):125–136, May 2011. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/23209725.

Konomi:2014:BTM

[KKS^{+}14] Bledar Konomi, Georgios Karagiannis, Avik Sarkar, Xin Sun, and Guang Lin. Bayesian treed multivariate Gaussian process with adaptive design: Application to a carbon capture unit. Technometrics, 56(2):145–158, 2014. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Kim:2017:HSV

[KL17] Heeyoung Kim and Jaehwan Lee. Hierarchical spatially varying coefficient process model. Technometrics, 59(4):521–527, 2017. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Kleijnen:2013:CDE

[Kle13] Jack P. C. Kleijnen. Comment: A discrete-event simulation perspective. Technometrics, 55(1):21–24, 2013. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Krishnamoorthy:2011:ILM

[KMM11] K. Krishnamoorthy, Avishek Mallick, and Thomas Mathew. Inference for the lognormal mean and quantiles based on samples with left and right Type I censoring. Technometrics, 53(1):72– 83, February 2011. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/

40997294. Supplementary materials available online.

Kang:2018:EBI

[KMQ18] Yicheng Kang, Partha Sarathi Mukherjee, and Peihua Qiu. Efficient blind image deblurring using nonparametric regression and local pixel clustering. Technometrics, 60(4):522–531, 2018. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Kang:2014:JDB

[KQ14] Yicheng Kang and Peihua Qiu. Jump detection in blurred regression surfaces. Technometrics, 56(4):539–550, 2014. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Kao:2015:MDU

[KRSA15] Yimin Kao, Brian Reich, Curtis Storlie, and Blake Anderson. Malware detection using nonparametric Bayesian clustering and classification techniques. Technometrics, 57(4):535–546, 2015. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00401706.2014.958916.

Kiriliouk:2019:PTM

[KRSW19] Anna Kiriliouk, Holger Rootzén, Johan Segers, and Jennifer L. Wadsworth. Peaks over thresholds modeling with multivariate generalized Pareto distributions. Technometrics, 61(1):123–135, 2019. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Kang:2016:CSV

[KSB16] Lulu Kang, Javier Cruz Salgado, and William A. Brenneman. Comparing the slack-variable mixture model with other alternatives. Technometrics, 58(2):255–268, 2016. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Kuhn:2010:BRB

[Kuh10] Max Kuhn. Book review: Lattice: Multivariate Data Visualization With R by Deepayan Sarkar. Technometrics, 52(3):363– 364, August 2010. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
27867260.

Kontar:2018:NMP

[KZS^{+}18] Raed Kontar, Shiyu Zhou, Chaitanya Sankavaram, Xinyu Du, and Yilu Zhang. Nonparametric modeling and prognosis of condition monitoring signals using multivariate Gaussian convolution processes. Technometrics, 60(4):484–496, 2018. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Lu:2011:ODE

[LACR11] Lu Lu, Christine M. Anderson-Cook, and Timothy J. Robinson. Optimization of designed experiments based on multiple criteria utilizing a Pareto frontier. Technometrics, 53(4):353–365, November 2011. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/41714949.

Lainez-Aguirre:2016:DSV

[LAMO^{+}16] José M. Laínez-Aguirre, Linas Mockus, Seza Orçun, Gary Blau, and Gintaras V. Reklaitis. A decomposition strategy for the variational inference of complex systems. Technometrics, 58(1):84– 94, 2016. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00401706.2014.995833.

Lazar:2010:BRB

[Laz10] Nicole Lazar. Book review: Elementary Probability for Applications by Rick Durrett. Technometrics, 52(4):467–468, November 2010. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/40997264.

Lazar:2011:BRB

[Laz11] Nicole Lazar. Book review: Introduction to Probability Simulation and Gibbs Sampling With R by Eric A. Suess; Bruce E. Trumbo. Technometrics, 53(3):327, August 2011. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/23210415.

Gratiet:2015:CBS

[LC15] Loic Le Gratiet and Claire Cannamela. Cokriging-based sequential design strategies using fast cross-validation techniques for multi-fidelity computer codes. Technometrics, 57(3):418–427, 2015. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00401706.2014.928233.

Lu:2013:CSS

[LCAc13] Lu Lu, Jessica L. Chapman, and Christine M. Anderson-cook. A case study on selecting a best allocation of new data for improving the estimation precision of system and subsystem reliability using Pareto fronts. Technometrics, 55(4):473–487, 2013. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Lawless:2012:TMT

[LÇC12] Jerald F. Lawless, Candemır Çiğşar, and Richard J. Cook. Testing for monotone trend in recurrent event processes. Technometrics, 54(2):147–158, May 2012. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/41714863.

Lawless:2012:MWC

[LCL12] J. F. Lawless, M. J. Crowder, and K.-A. Lee. Monitoring warranty claims with cusums. Technometrics, 54(3):269–278, August 2012. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/41714895.

Lim:2011:AFI

[LD11] Chae Young Lim and Sarat C. Dass. Assessing fingerprint individuality using EPIC: A case study in the analysis of spatially dependent marked processes. Technometrics, 53(2):112–124, May 2011. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/23209724.

Lee:2012:CDD

[Lee12] Herbert K. H. Lee. [Comment: DoIt and Do It Well]: Comment [MR2967968]. Technometrics, 54(3):231–232, August 2012. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/41714889.

Lennox:2013:BME

[LG13] Kristin P. Lennox and Lee G. Glascoe. A Bayesian measurement error model for misaligned radiographic data. Technometrics, 55 (4):450–460, 2013. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Li:2014:BNT

[LHDP14] Li Li, Timothy Hanson, Paul Damien, and Elmira Popova. A Bayesian nonparametric test for minimal repair. Technometrics, 56(3):393–406, 2014. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Lee:2018:SBD

[LHTD18] I-Chen Lee, Yili Hong, Sheng-Tsaing Tseng, and Tirthankar Dasgupta. Sequential Bayesian design for accelerated life tests. Technometrics, 60(4):472–483, 2018. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Li:2010:BRB

[Li10] Hongfei Li. Book review: Bayesian Disease Mapping — Hierarchical Modeling in Spatial Epidemiology by Andrew B. Lawson. Technometrics, 52(4):463, November 2010. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://

www.jstor.org/stable/40997256.

Lipovetsky:2010:BRBa

[Lip10a] Stan Lipovetsky. Book review: Computational Methods of Feature Selection by Huan Liu; Hiroshi Motoda. Technometrics, 52(4): 462–463, November 2010. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/40997255.

Lipovetsky:2010:BRBb

[Lip10b] Stan Lipovetsky. Book review: Introduction to Multivariate Statistical Analysis in Chemometrics by Kurt Varmuza; Peter Filzmoser. Technometrics, 52(4):468–469, November 2010. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/40997265.

Lipovetsky:2011:BRBa

[Lip11a] Stan Lipovetsky. Book review: Epidemics and Rumours in Complex Networks by Moez Draief; Laurent Massoulie. Technometrics, 53(2):219–220, May 2011. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/23209746.

Lipovetsky:2011:BRBc

[Lip11b] Stan Lipovetsky. Book review: Networks, Crowds, and Markets: Reasoning About a Highly Connected World by David Easley; Jon Kleinberg. Technometrics, 53(3):329–330, August 2011. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/23210420.

Lipovetsky:2011:BRBb

[Lip11c] Stan Lipovetsky. Book review: Regression With Linear Predictors by Per Kragh Andersen; Lene Theil Skovgaard. Technometrics, 53(3):326, August 2011. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/23210413.

Lipovetsky:2012:BRBd

[Lip12a] Stan Lipovetsky. Book review: Bayesian Logical Data Analysis for the Physical Sciences: A Comparative Approach With Mathematica^{T M} Support by Phil Gregory. Technometrics, 54(4): 443–444, November 2012. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/41714935.

Lipovetsky:2012:BRBa

[Lip12b] Stan Lipovetsky. Book review: Ranking and Prioritization for Multi-indicator Systems: Introduction to Partial Order Applications by Rainer Bruggemann; Ganapati P. Patil. Technometrics, 54(2):203–204, May 2012. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/41714878.

Lipovetsky:2012:BRBc

[Lip12c] Stan Lipovetsky. Book review: Statistical Image Processing and Multidimensional Modeling by Paul Fieguth. Technometrics, 54 (4):440–441, November 2012. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/41714929.

Lipovetsky:2012:BRBb

[Lip12d] Stan Lipovetsky. Book review: Understanding Violence: The Intertwining of Morality, Religion and Violence: A Philosophical Stance by Lorenzo Magnani. Technometrics, 54(3):326, August 2012. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/41714906.

Lipovetsky:2019:BRAa

[Lip19a] Stan Lipovetsky. Book review: Advances in Latent Variables: Methods, Models and Applications. Technometrics, 61(3):423– 424, 2019. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Lipovetsky:2019:BRAb

[Lip19b] Stan Lipovetsky. Book review: Applied Data-Centric Social Sciences: Concepts, Data, Computation, and Theory. Technometrics, 61(3):424–425, 2019. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Lipovetsky:2019:BRD

[Lip19c] Stan Lipovetsky. Book review: Digital Humanities and Film Studies: Visualising Dziga Vertov’s Work. Technometrics, 61 (4):562–563, 2019. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Lipovetsky:2019:BRM

[Lip19d] Stan Lipovetsky. Book review: Machine Learning in Medicine — A Complete Overview. Technometrics, 61(3):425–426, 2019. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Lipovetsky:2019:BRSb

[Lip19e] Stan Lipovetsky. Book review: Simulating Societal Change: Counterfactual Modelling for Social and Policy Inquiry. Technometrics, 61(4):563–565, 2019. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Lipovetsky:2019:BRSa

[Lip19f] Stan Lipovetsky. Book review: Strategic Economic Decision-Making: Using Bayesian Belief Networks to Solve Complex Problems. Technometrics, 61(3):422, 2019. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Liu:2012:PAL

[Liu12] Xiao Liu. Planning of accelerated life tests with dependent failure modes based on a gamma frailty model. Technometrics, 54(4):

398–409, November 2012. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/41714924.

Liang:2016:BMH

[LKS16] Faming Liang, Jinsu Kim, and Qifan Song. A bootstrap Metropolis–Hastings algorithm for Bayesian analysis of big data. Technometrics, 58(3):304–??, 2016. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Li:2016:NFD

[LL16] William Li and Dennis K. J. Lin. A note on foldover of 2^{k−p} designs with column permutations. Technometrics, 58(4):508– 512, 2016. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Lock:2013:SAS

[LM13] Amy B. Lock and Max D. Morris. Significance of angle in the statistical comparison of forensic tool marks. Technometrics, 55 (4):548–561, 2013. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Liu:2015:SME

[LM15] Shiyao Liu and William Q. Meeker. Statistical methods for estimating the minimum thickness along a pipeline. Technometrics, 57(2):164–179, 2015. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Liu:2015:ASS

[LMS15] Kaibo Liu, Yajun Mei, and Jianjun Shi. An adaptive sampling strategy for online high-dimensional process monitoring. Technometrics, 57(3):305–319, 2015. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/00401706.2014.947005.

Li:2014:PMA

[LMT14] Ming Li, William Q. Meeker, and R. Bruce Thompson. Physical model-assisted probability of detection of flaws in titanium forgings using ultrasonic nondestructive evaluation. Technometrics, 56(1):78–91, 2014. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Lukemire:2019:QQB

[LMW19] Joshua Lukemire, Abhyuday Mandal, and Weng Kee Wong. d-QPSO: A quantum-behaved particle swarm technique for finding D-optimal designs with discrete and continuous factors and a binary response. Technometrics, 61(1):77–87, 2019. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Li:2019:UIF

[LMZ19] William Li, Robert W. Mee, and Qi Zhou. Using individual factor information in fractional factorial designs. Technometrics, 61(1): 38–49, 2019. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Leyva:2013:BSA

[LPVW13] Norma Leyva, Garritt L. Page, Stephen B. Vardeman, and Joanne R. Wendelberger. Bayes statistical analyses for particle sieving studies. Technometrics, 55(2):224–231, 2013. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Li:2014:SPC

[LQ14] Zhonghua Li and Peihua Qiu. Statistical process control using a dynamic sampling scheme. Technometrics, 56(3):325–335, 2014. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Linkletter:2012:ECT

[LRL^{+}12a] C. D. Linkletter, P. Ranjan, C. D. Lin, D. R. Bingham, W. A. Brenneman, R. A. Lockhart, and T. M. Loughin. Erratum: “Compliance Testing for Random Effects Models with Joint Acceptance Criteria” [MR2967975]. Technometrics, 54(4):450, November 2012. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/

41714944. See [LRL^{+}12b].

Linkletter:2012:CTR

[LRL^{+}12b] Crystal D. Linkletter, Pritam Ranjan, C. Devon Lin, Derek R. Bingham, William A. Brenneman, Richard A. Lockhart, and Thomas M. Loughin. Compliance testing for random effects models with joint acceptance criteria. Technometrics, 54(3):243– 255, August 2012. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/

41714893. See erratum [LRL^{+}12a].

Lin:2013:CSS

[LS13] Dennis K. J. Lin and Weijie Shen. Comment: Some statistical concerns on dimensional analysis. Technometrics, 55(3):281–285, 2013. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Lim:2013:RAH

[LSP13] Changwon Lim, Pranab K. Sen, and Shyamal D. Peddada. Robust analysis of high throughput screening (HTS) assay data. Technometrics, 55(2):150–160, 2013. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Lu:2011:BRB

[Lu11] Z. Q. John Lu. Book review: Perturbation Bounds for Matrix Eigenvalues by Rajendra Bhatia. Technometrics, 53(4):443, November 2011. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
41714960.

Lawrence:2013:MBS

[LVB13] Earl Lawrence, Scott Vander Wiel, and Russell Bent. Model bank state estimation for power grids using importance sampling. Technometrics, 55(4):426–435, 2013. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Li:2012:CRC

[LW12] Bo Li and Xiao Wang. [Clustering Random Curves Under Spatial Interdependence With Application to Service Accessibility]: Comment. Technometrics, 54(2):127–128, May 2012. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/41714859. See [JS12a].

Loeppky:2013:GSA

[LWM13] Jason L. Loeppky, Brian J. Williams, and Leslie M. Moore. Global sensitivity analysis for mixture experiments. Technometrics, 55(1):68–78, 2013. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Lee:2018:NFM

[LX18] Kevin H. Lee and Lingzhou Xue. Nonparametric finite mixture of Gaussian graphical models. Technometrics, 60(4):511–521, 2018. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Li:2014:CMS

[LZ14] William Li and Ji Zhu. Comment: Model selection with strong and weak heredity constraints. Technometrics, 56(1):21–22, 2014. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Li:2015:MBS

[LZ15] Jie Li and Dale L. Zimmerman. Model-based sampling design for multivariate geostatistics. Technometrics, 57(1):75–86, 2015. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Li:2016:PMM

[LZ16] Yongxiang Li and Qiang Zhou. Pairwise meta-modeling of multivariate output computer models using nonseparable covariance function. Technometrics, 58(4):483–494, 2016. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Li:2018:PEM

[LZHZ18] Yongxiang Li, Qiang Zhou, Xiaohu Huang, and Li Zeng. Pairwise estimation of multivariate Gaussian process models with replicated observations: Application to multivariate profile monitoring. Technometrics, 60(1):70–78, 2018. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Li:2015:RSE

[LZW15] Yun Li, Ji Zhu, and Naisyin Wang. Regularized semiparametric estimation for ordinary differential equations. Technometrics, 57(3):341–350, 2015. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00401706.2015.1006338.

Magnan:2011:BRB

[Mag11] Shon Magnan. Book review: Bayesian Analysis for the Social Sciences by Simon Jackman. Technometrics, 53(2):218, May 2011. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/23209744.

Maitra:2011:BRB

[Mai11] Ranjan Maitra. Book review: The Statistical Analysis of Functional MRI Data by Nicole A. Lazar. Technometrics, 53(2):211– 212, May 2011. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
23209735.

Mallik:2019:WPT

[Mal19] Abhirup Mallik. Wavelet packets and their statistical applications. Technometrics, 61(3):426, 2019. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Maronna:2011:RRR

[Mar11] Ricardo A. Maronna. Robust ridge regression for highdimensional data. Technometrics, 53(1):44–53, February 2011. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/40997291. Sup-

plementary materials available online.

Marzjarani:2019:BRE

[Mar19a] Morteza Marzjarani. Book review: Exploratory Data Analysis With MATLAB by Wendy L. Martinez, Angel R. Martinez, and Jeffery L. Solka. Boca Raton, FL: CRC Press, 2017, xxv + 590 pp., $100.00 (Hardback), $46.36 (e-book), ISBN-13: 978-1-4987-7606-6. Technometrics, 61(4):565–566, 2019. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Marzjarani:2019:BRS

[Mar19b] Morteza Marzjarani. Book review: Simulation and the Monte Carlo Method (3rd ed.). Technometrics, 61(3):427–428, 2019. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Monteiro:2014:BMP

[MBR14] João V. D. Monteiro, Sudipto Banerjee, and Gurumurthy Ramachandran. Bayesian modeling for physical processes in industrial hygiene using misaligned workplace data. Technometrics, 56 (2):238–247, 2014. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

McComb:2010:BRB

[McC10] Mark A. McComb. Book review: Interpreting Economic and Social Data by Othmar W. Winkler. Technometrics, 52(4):466– 467, November 2010. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
40997263.

McComb:2012:BRBc

[McC12a] Mark A. McComb. Book review: Business Analytics for Managers by Wolfgang Jank. Technometrics, 54(4):442–443, November 2012. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/41714932.

McComb:2012:BRBa

[McC12b] Mark A. McComb. Book review: Fundamentals of Probability: A First Course by Anirban DasGupta. Technometrics, 54(1): 100, February 2012. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
41714838.

McComb:2012:BRBb

[McC12c] Mark A. McComb. Book review: Large Sample Techniques for Statistics by Jiming Jiang. Technometrics, 54(1):100–101, February 2012. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/41714840.

Moon:2012:TSS

[MDS12] Hyejung Moon, Angela M. Dean, and Thomas J. Santner. Two-stage sensitivity-based group screening in computer experiments. Technometrics, 54(4):376–387, November 2012. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/41714922.

McGree:2012:RDP

[ME12] James M. McGree and John A. Eccleston. Robust designs for Poisson regression models. Technometrics, 54(1):64–72, February 2012. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/41714832.

Mylona:2014:ODB

[MGJ14] Kalliopi Mylona, Peter Goos, and Bradley Jones. Optimal design of blocked and split-plot experiments for fixed effects and variance component estimation. Technometrics, 56(2):132–144, 2014. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Menafoglio:2018:PMP

[MGSC18] Alessandra Menafoglio, Marco Grasso, Piercesare Secchi, and Bianca Maria Colosimo. Profile monitoring of probability density functions via simplicial functional PCA with application to image data. Technometrics, 60(4):497–510, 2018. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Michelson:2010:BRB

[Mic10] Diane K. Michelson. Book review: Variations on Split Plot and Split Block Experiment Designs by Walter T. Federer; Freedom King. Technometrics, 52(2):258–259, May 2010. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/27867231.

Mitchell:2010:AHA

[Mit10] Robert H. Mitchell. Announcement of 2010 Hunter award. Technometrics, 52(2):147, 2010. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Ma:2019:SSD

[MKBN19] Pulong Ma, Emily L. Kang, Amy J. Braverman, and Hai M. Nguyen. Spatial statistical downscaling for constructing highresolution nature runs in global observing system simulation experiments. Technometrics, 61(3):322–340, 2019. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Martin:2016:PFP

[ML16] Ryan Martin and Rama T. Lingham. Prior-free probabilistic prediction of future observations. Technometrics, 58(2):225–235, 2016. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Mittman:2019:HMH

[MLBM19] Eric Mittman, Colin Lewis-Beck, and William Q. Meeker. A hierarchical model for heterogeneous reliability field data. Technometrics, 61(3):354–368, 2019. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Myers:2016:PLS

[MLF^{+}16] Kary Myers, Earl Lawrence, Michael Fugate, Claire McKay Bowen, Lawrence Ticknor, Jon Woodring, Joanne Wendelberger, and Jim Ahrens. Partitioning a large simulation as it runs. Technometrics, 58(3):329–??, 2016. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Marget:2019:CCE

[MM19] Wilmina M. Marget and Max D. Morris. Central composite experimental designs for multiple responses with different models. Technometrics, 61(4):524–532, 2019. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Mondal:2014:BUQ

[MMEDG14] Anirban Mondal, Bani Mallick, Yalchin Efendiev, and Akhil Datta-Gupta. Bayesian uncertainty quantification for subsurface inversion using a multiscale hierarchical model. Technometrics, 56(3):381–392, 2014. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Morris:2012:GSC

[Mor12] Max D. Morris. Gaussian surrogates for computer models with time-varying inputs and outputs. Technometrics, 54(1):42–50, February 2012. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
41714830.

Morris:2015:PED

[Mor15] Max D. Morris. Physical experimental design in support of computer model development. Technometrics, 57(1):45–53, 2015. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Mukherjee:2011:DID

[MQ11] Partha Sarathi Mukherjee and Peihua Qiu. 3-D image denoising by local smoothing and nonparametric regression. Technometrics, 53(2):196–208, May 2011. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/23209731.

Mee:2017:SON

[MSE17] Robert W. Mee, Eric D. Schoen, and David J. Edwards. Selecting an orthogonal or nonorthogonal two-level design for screening. Technometrics, 59(3):305–318, 2017. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Mak:2019:AMT

[MW19] Simon Mak and C. F. Jeff Wu. Analysis-of-marginal-tail-means (ATM): a robust method for discrete black-box optimization.

Technometrics, 61(4):545–559, 2019. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Myers:2011:BRBa

[Mye11a] Donald E. Myers. Book review: Quantitative Fund Management by M. A. H. Dempster; Gautam Mitra; Georg Pflug. Technometrics, 53(1):102–103, February 2011. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/40997303.

Myers:2011:BRBb

[Mye11b] Donald E. Myers. Book review: Regression Modeling With Actuarial and Financial Applications by Edward W. Frees. Technometrics, 53(2):220, May 2011. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/23209747.

Myers:2011:BRBc

[Mye11c] Donald E. Myers. Book review: Theory of Stochastic Processes: With Applications to Financial Mathematics and Risk Theory by Dmytro Gusak; Alexander Kukush; Alexey Kulik; Yuliya Mishura; Andrey Pilipenko. Technometrics, 53(3):324–325, August 2011. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/23210411.

Myers:2012:BRBa

[Mye12a] Donald Myers. Book review: Finite Mathematics for Business, Economics, Life Sciences, and Social Sciences by Raymond A. Barnett; Michael R. Ziegler; Karl E. Byleen. Technometrics, 54 (1):103–104, February 2012. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/41714845.

Myers:2012:BRBc

[Mye12b] Donald E. Myers. Book review: Forest Analytics With R by Andrew P. Robinson; Jeff D. Hamann. Technometrics, 54(3):320– 322, August 2012. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
41714901.

Myers:2012:BRBb

[Mye12c] Donald E. Myers. Book review: Numerical Ecology with R by Daniel Bocard; François Gillet; Pierre Legendre. Technometrics, 54(2):200–201, May 2012. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/41714872.

Myers:2012:BRBd

[Mye12d] Donald E. Myers. Book review: Smoothing Splines — Methods and Applications by Yuedong Wang. Technometrics, 54(3):324– 326, August 2012. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
41714905.

Mai:2013:NCE

[MZ13] Qing Mai and Hui Zou. A note on the connection and equivalence of three sparse linear discriminant analysis methods. Technometrics, 55(2):243–246, 2013. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Mai:2015:NVT

[MZ15] Qing Mai and Hui Zou. Nonparametric variable transformation in sufficient dimension reduction. Technometrics, 57(1):1–10, 2015. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Nam:2015:USS

[NAEK15] Christopher F. H. Nam, John A. D. Aston, Idris A. Eckley, and Rebecca Killick. The uncertainty of storm season changes: Quantifying the uncertainty of autocovariance changepoints. Technometrics, 57(2):194–206, 2015. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Nagaraja:2011:BSM

[Nag11] H. N. Nagaraja. Scan Statistics: Methods and Applications by Joseph Glaz; Vladimir Pozdnyakov; Sylvan Wallenstein. Technometrics, 53(1):103, February 2011. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/40997304.

Natarajan:2012:BRB

[Nat12] Ramesh Natarajan. Book review: Composite Sampling — A Novel Method to Accomplish Observational Economy in Environmental Studies by Ganapati P. Patil; Sharad D. Gore; Charles Taillie. Technometrics, 54(1):101–102, February 2012. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/41714841.

Neubauer:2010:BRBb

[Neu10a] Dean V. Neubauer. Book review: Design of Comparative Experiments by R. A. Bailey. Technometrics, 52(2):261–263, May 2010. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/27867237.

Neubauer:2010:BRBa

[Neu10b] Dean V. Neubauer. Book review: The Role of Statistics in Business and Industry by Gerald J. Hahn; Necip Doganaksoy. Technometrics, 52(1):139–140, February 2010. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://

www.jstor.org/stable/40586689.

Neubauer:2011:BRB

[Neu11] Dean V. Neubauer. Book review: Introducing Monte Carlo Methods With R by Christian P. Robert; George Casella. Technometrics, 53(2):217–218, May 2011. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/23209743.

Neubauer:2012:BRBc

[Neu12a] Dean V. Neubauer. Book review: A Concise Guide to Statistics by Hans-Michael Kaltenbach. Technometrics, 54(4):444–445, November 2012. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
41714936.

Neubauer:2012:BRBb

[Neu12b] Dean V. Neubauer. Book review: Nonparametric Statistical Tests: A Computational Approach by Markus Neuhauser. Technometrics, 54(3):329–330, August 2012. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://

www.jstor.org/stable/41714911.

Neubauer:2012:BRBa

[Neu12c] Dean V. Neubauer. Book review: The R Primer by Claus Thorn Ekstrøm. Technometrics, 54(3):326–327, August 2012. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/41714907.

Ng:2010:BRB

[Ng10] Hon Keung Tony Ng. Book review: Monte Carlo and Quasi-Monte Carlo Sampling by Christiane Lemieux. Technometrics, 52(1):140–141, February 2010. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/40586690.

Ng:2011:BRB

[Ng11] Hon Keung Tony Ng. Book review: Introduction to Scientific Programming and Simulation Using R by Owen Jones; Robert Maillardet; Andrew Robinson. Technometrics, 53(2):213–214, May 2011. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/23209737.

Ng:2012:BRB

[Ng12] Hon Keung Tony Ng. Book review: Statistical Programming in SAS by A. John Bailer. Technometrics, 54(4):441, November 2012. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/41714930.

Neil:2013:SSO

[NHB^{+}13] Joshua Neil, Curtis Hash, Alexander Brugh, Mike Fisk, and Curtis B. Storlie. Scan statistics for the online detection of locally anomalous subgraphs. Technometrics, 55(4):403–414, 2013. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Nie:2017:ALS

[NHQ17] Xiao Nie, Jared Huling, and Peter Z. G. Qian. Accelerating largescale statistical computation with the GOEM algorithm. Technometrics, 59(4):416–425, 2017. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Nguyen:2014:STD

[NKCB14] Hai Nguyen, Matthias Katzfuss, Noel Cressie, and Amy Braverman. Spatio-temporal data fusion for very large remote sensing datasets. Technometrics, 56(2):174–185, 2014. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Nkurunziza:2010:BRB

[Nku10] Sévérien Nkurunziza. Book review: Correlated Data Analysis: Modeling, Analytics, and Applications by Peter X.-K. Song. Technometrics, 52(2):259, May 2010. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/27867232.

Ngueyep:2015:LVA

[NS15] Rodrigue Ngueyep and Nicoleta Serban. Large-vector autoregression for multilayer spatially correlated time series. Technometrics, 57(2):207–216, 2015. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Nobre:2011:SVA

[NSS11] Aline A. Nobre, Bruno Sansó, and Alexandra M. Schmidt. Spatially varying autoregressive processes. Technometrics, 53(3):310– 321, August 2011. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
23210406.

OConnor:2010:RGM

[O’C10] Patrick D. T. O’Connor. [Reliability Growth Management Metrics and Statistical Methods for Discrete-Use Systems]: Comment. Technometrics, 52(4):391–392, November 2010. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/40997243.

Olive:2010:BRBa

[Oli10a] David J. Olive. Book review: Analysis of Variance Designs: A Conceptual and Computational Approach With SPSS and SAS by G. Gamst; L. S. Meyers; A. J. Guarino. Technometrics, 52(1): 139, February 2010. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
40586688.

Olive:2010:BRBc

[Oli10b] David J. Olive. Book review: Mixed Effects Models and Extensions in Ecology With R by A. F. Zuur; E. N. Ieno; N. J. Walker; A. A. Saveliev; G. M. Smith. Technometrics, 52(4):464–465, November 2010. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
40997258.

Olive:2010:BRBb

[Oli10c] David J. Olive. Book review: Software for Data Analysis: Programming With R by J. M. Chambers. Technometrics, 52(2):261, May 2010. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/27867236.

Olive:2011:BRBb

[Oli11a] David J. Olive. Book review: Fundamentals of Modern Statistical Methods (2nd ed.) by R. R. Wilcox. Technometrics, 53(2):221, May 2011. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/23209750.

Olive:2011:BRBa

[Oli11b] David J. Olive. Book review: Linear Model Methodology by A. I. Khuri. Technometrics, 53(2):214, May 2011. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://

www.jstor.org/stable/23209738.

Olive:2012:BRBb

[Oli12a] David J. Olive. Book review: A Practitioner’s Guide to Resampling for Data Analysis, Data Mining, and Modeling by P. I. Good. Technometrics, 54(3):327–328, August 2012. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/41714909.

Olive:2012:BRBc

[Oli12b] David J. Olive. Book review: Elements of Distribution Theory by T. A. Severini. Technometrics, 54(4):443, November 2012. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/41714934.

Olive:2012:BRBa

[Oli12c] David J. Olive. Book review: Projection Matrices, Generalized Inverse Matrices, and Singular Value Decomposition by H. Yanai; K. Takeuchi; Y. Takane. Technometrics, 54(2):201, May 2012. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/41714873.

Orsini:2019:BRA

[Ors19] Joyce Nilsson Orsini. Book review: Applied Quantitative Finance (3rd ed.). Technometrics, 61(2):273–280, 2019. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Ormerod:2012:CDD

[OW12] John T. Ormerod and M. P. Wand. [Comment: DoIt and Do It Well]: Comment [MR2967968]. Technometrics, 54(3):233– 236, August 2012. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
41714890.

Overstall:2017:BDE

[OW17] Antony M. Overstall and David C. Woods. Bayesian design of experiments using approximate coordinate exchange. Technometrics, 59(4):458–470, 2017. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Oakley:2017:CSC

[OY17] Jeremy E. Oakley and Benjamin D. Youngman. Calibration of stochastic computer simulators using likelihood emulation. Technometrics, 59(1):80–92, 2017. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Plumlee:2017:LBK

[PA17] Matthew Plumlee and Daniel W. Apley. Lifted Brownian kriging models. Technometrics, 59(2):165–177, 2017. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Park:2011:BRB

[Par11] Jin-Hong Park. Book review: Data Clustering: Theory, Algorithms, and Applications by Guojun Gan; Chaoqun Ma; Jianhong Wu. Technometrics, 53(2):210, May 2011. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://

www.jstor.org/stable/23209733.

Park:2014:EMP

[Par14] Chiwoo Park. Estimating multiple pathways of object growth using nonlongitudinal image data. Technometrics, 56(2):186–199, 2014. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Picard:2013:QER

[PBH13] Rick Picard, Tom Burr, and Michael S. Hamada. Quantile estimation for radiation portal monitoring. Technometrics, 55(1): 94–102, 2013. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Phoa:2016:OTL

[PCWW16] Frederick Kin Hing Phoa, Ray-Bing Chen, Weichung Wang, and Weng Kee Wong. Optimizing two-level supersaturated designs using swarm intelligence techniques. Technometrics, 58(1):43– 49, 2016. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00401706.2014.981346.

Page:2011:BLC

[PD11] Garritt L. Page and David B. Dunson. Bayesian local contamination models for multivariate outliers. Technometrics, 53(2): 152–162, May 2011. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
23209727.

Peng:2010:BRB

[Pen10] Yingwei Peng. Book review: Niche Modeling — Predictions From Statistical Distributions by David Stockwell. Technometrics, 52(3):363, August 2010. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/27867259.

Peng:2015:IGP

[Pen15] Chien-Yu Peng. Inverse Gaussian processes with random effects and explanatory variables for degradation data. Technometrics, 57(1):100–111, 2015. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Pfahler:2012:BRB

[Pfa12] Lori Pfahler. Book review: A Handbook of Statistical Analyses Using R (2nd ed.) by Brian S. Everitt; Torsten Hothorn. Technometrics, 54(1):100, February 2012. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://

www.jstor.org/stable/41714839.

Pan:2014:CDD

[PFF14] Jianxin Pan, Yu Fei, and Peter Foster. Case-deletion diagnostics for linear mixed models. Technometrics, 56(3):269–281, 2014. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Picheny:2016:CSE

[PGK16] Victor Picheny, David Ginsbourger, and Tipaluck Krityakierne. Comment: Some enhancements over the augmented Lagrangian approach. Technometrics, 58(1):17–21, 2016. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic).  URL http://www.tandfonline.com/doi/abs/10.
1080/00401706.2015.1079246. See [GGL^{+}16a].

Picheny:2013:QBO

[PGRC13a] Victor Picheny, David Ginsbourger, Yann Richet, and Gregory Caplin. Quantile-based optimization of noisy computer experiments with tunable precision. Technometrics, 55(1):2–13, 2013. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Picheny:2013:R

[PGRC13b] Victor Picheny, David Ginsbourger, Yann Richet, and Gregory Caplin. Rejoinder. Technometrics, 55(1):31–36, 2013. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Parker:2011:DEC

[PGS11] Ben M. Parker, Steven G. Gilmour, and John A. Schormans. Design of experiments for categorical repeated measurements in packet communication networks. Technometrics, 53(4):339–352, November 2011. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
41714948.

Pratola:2016:BAR

[PH16] M. T. Pratola and D. M. Higdon. Bayesian additive regression tree calibration of complex high-dimensional computer models. Technometrics, 58(2):166–179, 2016. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Pratola:2017:DAE

[PHBF17] Matthew T. Pratola, Ofir Harari, Derek Bingham, and Gwenn E. Flowers. Design and analysis of experiments on nonconvex regions. Technometrics, 59(1):36–47, 2017. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Pourhabib:2016:STW

[PHD16] Arash Pourhabib, Jianhua Z. Huang, and Yu Ding. Shortterm wind speed forecast using measurements from multiple turbines in a wind farm. Technometrics, 58(1):138–147, 2016. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00401706.2014.988291.

Park:2018:BSM

[PHK^{+}18] Eun Sug Park, Philip K. Hopke, Inyoung Kim, Shuman Tan, and Clifford H. Spiegelman. Bayesian spatial multivariate receptor modeling for multisite multipollutant data. Technometrics, 60 (3):306–318, 2018. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Piepel:2013:CSC

[Pie13] Greg F. Piepel. Comment: Spurious correlation and other observations on experimental design for engineering dimensional analysis. Technometrics, 55(3):286–289, 2013. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Paynabar:2015:ISF

[PJR15] Kamran Paynabar, Judy Jin, and Matthew P. Reed. Informative sensor and feature selection via hierarchical nonnegative garrote. Technometrics, 57(4):514–523, 2015. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/00401706.2014.947383.

Plumlee:2013:CAS

[PJW13] Matthew Plumlee, V. Roshan Joseph, and C. F. Jeff Wu. Comment: Alternative strategies for experimental design. Technometrics, 55(3):289–292, 2013. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Pronzato:2017:BLK

[PR17] Luc Pronzato and Maria-João Rendas. Bayesian local kriging. Technometrics, 59(3):293–304, 2017. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Pratola:2013:FSC

[PSB^{+}13] Matthew T. Pratola, Stephan R. Sain, Derek Bingham, Michael Wiltberger, and E. Joshua Rigler. Fast sequential computer model calibration of large nonstationary spatial-temporal processes. Technometrics, 55(2):232–242, 2013. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Pimentel:2017:ETF

[PSR17] Samuel D. Pimentel, Dylan S. Small, and Paul R. Rosenbaum. An exact test of fit for the Gaussian linear model using optimal nonbipartite matching. Technometrics, 59(3):330–337, 2017. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Plumlee:2014:BAE

[PT14] Matthew Plumlee and Rui Tuo. Building accurate emulators for stochastic simulations via quantile kriging. Technometrics, 56 (4):466–473, 2014. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Pena:2019:EDQ

[PTZ19] Daniel Peña, Ruey S. Tsay, and Ruben Zamar. Empirical dynamic quantiles for visualization of high-dimensional time series. Technometrics, 61(4):429–444, 2019. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Paynabar:2016:CPA

[PZQ16] Kamran Paynabar, Changliang Zou, and Peihua Qiu. A changepoint approach for Phase-I analysis in multivariate profile monitoring and diagnosis. Technometrics, 58(2):191–204, 2016. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Qian:2014:ANL

[QAHS14] Peter Z. G. Qian, Mingyao Ai, Youngdeok Hwang, and Heng Su. Asymmetric nested lattice samples. Technometrics, 56(1):46–54, 2014. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Qian:2010:BRBb

[Qia10a] Lianfen Qian. Book review: ROC Curves for Continuous Data by Wojtek J. Krzanowski; David J. Hand. Technometrics, 52 (4):466, November 2010. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/40997261.

Qian:2010:BRBa

[Qia10b] Lianfen Qian. Book review: Time Series Analysis With Applications in R (2nd ed.) by Jonathan D. Cryer; Kung-Sik Chan. Technometrics, 52(3):365, August 2010. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/27867262.

Qiu:2012:BRB

[Qiu12] Peihua Qiu. Book review: Microarray Image Analysis: An Algorithmic Approach by Karl Fraser; Zidong Wang; Xiaohui Liu. Technometrics, 54(2):199, May 2012. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/41714870.

Qiu:2014:ER

[Qiu14] Peihua Qiu. Editor’s report. Technometrics, 56(4):417–418, 2014. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Qiu:2015:ER

[Qiu15] Peihua Qiu. Editor’s report. Technometrics, 57(4):446–448, 2015. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00401706.2015.1116318.

Qiu:2016:BDM

[Qiu16a] Peihua Qiu. Big data? More challenges! Technometrics, 58 (3):283–??, 2016. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Qiu:2016:ER

[Qiu16b] Peihua Qiu. Editor’s report. Technometrics, 58(4):413–414, 2016. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Qiu:2011:NSP

[QL11] Peihua Qiu and Zhonghua Li. On nonparametric statistical process control of univariate processes. Technometrics, 53(4):390– 405, November 2011. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
41714952.

Qi:2019:CBL

[QL19] Zhengling Qi and Yufeng Liu. Convex bidirectional large margin classifiers. Technometrics, 61(2):176–186, 2019. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Qian:2019:IGA

[QLS^{+}19] Wei Qian, Wending Li, Yasuhiro Sogawa, Ryohei Fujimaki, Xitong Yang, and Ji Liu. An interactive greedy approach to group sparsity in high dimensions. Technometrics, 61(3):409–421, 2019. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Qu:2010:ORC

[Qu10] Xianggui Qu. Optimal row-column designs in high-throughput screening experiments. Technometrics, 52(4):409–420, November 2010. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/40997248.

Qu:2011:BRB

[Qu11] Xianggui Qu. Book review: Theoretical Statistics: Topics for a Core Course by Robert W. Keener. Technometrics, 53(3): 330, August 2011. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
23210421.

Qu:2012:BRB

[Qu12] Xianggui Qu. Book review: Negative Binomial Regression (2nd ed.) by Joseph M. Hilbe. Technometrics, 54(2):202–203, May 2012. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/41714876.

Qu:2014:TBE

[QWKR14] Liang Qu, Zhang Wu, Michael B. C. Khoo, and Abdur Rahim. Time-between-event control charts for sampling inspection. Technometrics, 56(3):336–346, 2014. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Qiu:2013:NIR

[QX13] Peihua Qiu and Chen Xing. On nonparametric image registration. Technometrics, 55(2):174–188, 2013. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Qiu:2014:UDS

[QX14] Peihua Qiu and Dongdong Xiang. Univariate dynamic screening system: An approach for identifying individuals with irregular longitudinal behavior. Technometrics, 56(2):248–260, 2014. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Qiu:2010:NPMa

[QZW10a] Peihua Qiu, Changliang Zou, and Zhaojun Wang. Nonparametric profile monitoring by mixed effects modeling. Technometrics, 52 (3):265–277, August 2010. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/27867245.

Qiu:2010:NPMb

[QZW10b] Peihua Qiu, Changliang Zou, and Zhaojun Wang. [Nonparametric Profile Monitoring by Mixed Effects Modeling]: Rejoinder. Technometrics, 52(3):288–293, August 2010. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://

www.jstor.org/stable/27867250.

Qiu:2018:NDC

[QZZ18] Peihua Qiu, Xuemin Zi, and Changliang Zou. Nonparametric dynamic curve monitoring. Technometrics, 60(3):386–397, 2018. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Ranjan:2013:CCN

[Ran13] Pritam Ranjan. Comment: EI criteria for noisy computer simulators. Technometrics, 55(1):24–28, 2013. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Ryan:2010:MAF

[RB10] Kenneth J. Ryan and Dursun A. Bulutoglu. Minimum aberration fractional factorial designs with large N. Technometrics, 52(2): 250–255, May 2010. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/

27867229. With supplementary material available online.

Ranjan:2008:SED

[RBM08] Pritam Ranjan, Derek Bingham, and George Michailidis. Sequential experiment design for contour estimation from complex computer codes. Technometrics, 50(4):527–541, November 2008. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25471535. See

errata [RBM11].

Ranjan:2011:ESE

[RBM11] P. Ranjan, D. Bingham, and G. Michailidis. Errata: “Sequential experiment design for contour estimation from complex computer codes” [MR2655651]. Technometrics, 53(1):109–110, February 2011. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/40997315.

See [RBM08].

Ranjan:2011:CSA

[RHK11] Pritam Ranjan, Ronald Haynes, and Richard Karsten. A computationally stable approach to Gaussian process interpolation of deterministic computer simulation data. Technometrics, 53(4): 366–378, November 2011. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/41714950.

Raftery:2010:OPU

[RKE10] Adrian E. Raftery, Miroslav Kárný, and Pavel Ettler. Online prediction under model uncertainty via dynamic model averaging: Application to a cold rolling mill. Technometrics, 52(1):52– 66, February 2010. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
40586680.

Reising:2011:MST

[RMVH11] Monica Reising, Max Morris, Stephen Vardeman, and Shawn Higbee. Modeling SpectralTemporal data from point source events. Technometrics, 53(2):183–195, May 2011. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://

www.jstor.org/stable/23209730.

Roberts:2015:AAM

[RMZ15] Steven Roberts, Michael A. Martin, and Letian Zheng. An adaptive, automatic multiple-case deletion technique for detecting influence in regression. Technometrics, 57(3):408–417, 2015. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00401706.2014.946152.

Rosenblad:2011:BRB

[Ros11] Andreas Rosenblad. Book review: Logistic Regression Models by Joseph M. Hilbe. Technometrics, 53(2):216, May 2011. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/23209741.

Ross:2011:NMD

[RTA11] Gordon J. Ross, Dimitris K. Tasoulis, and Niall M. Adams. Nonparametric monitoring of data streams for changes in location and scale. Technometrics, 53(4):379–389, November 2011. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/41714951.

Rukhin:2015:CRI

[Ruk15] Andrew L. Rukhin. Confidence regions and intervals for meta-analysis model parameters. Technometrics, 57(4):547–558, 2015. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00401706.2014.962707.

Rutherford:2012:BRB

[Rut12] Jim Rutherford. Book review: SAS(R) Programming for Enterprise Guide(R) Users (2nd ed.) by Neil Constable. Technometrics, 54(1):104, February 2012. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/41714846.

Rousseeuw:2018:DDD

[RV18] Peter J. Rousseeuw and Wannes Van Den Bossche. Detecting deviating data cells. Technometrics, 60(2):135–145, 2018. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Richards:2015:SNP

[RWP15] Sarah C. Richards, William H. Woodall, and Gregory Purdy. Surveillance of nonhomogeneous Poisson processes. Technometrics, 57(3):388–394, 2015. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00401706.2014.927790.

Rootzxen:2016:TEW

[RZ16] Holger Rootzén and Dmitrii Zholud. Tail estimation for windowcensored processes. Technometrics, 58(1):95–103, 2016. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic).  URL http://www.tandfonline.com/doi/abs/10.
1080/00401706.2014.995834.

Stevens:2017:CRR

[SAC17] Nathaniel T. Stevens and Christine M. Anderson-Cook. Comparing the reliability of related populations with the probability of agreement. Technometrics, 59(3):371–380, 2017. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Samarov:2017:CDB

[SAH^{+}17] Daniel V. Samarov, David Allen, Jeeseong Hwang, Young Jong Lee, and Maritoni Litorja. A coordinate-descent-based approach to solving the sparse group elastic net. Technometrics, 59(4):437– 445, 2017. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Sandry:2011:BRB

[San11] Thomas D. Sandry. Book review: An Intermediate Course in Probability (2nd ed.) by Allan Gut. Technometrics, 53(1):103– 104, February 2011. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
40997305.

Sarkar:2011:BRB

[Sar11] Pradipta Sarkar. Book review: Analyzing and Interpreting Continuous Data Using JMP: A Step-by-Step Guide by Jose G. Ramirez; Brenda S. Ramirez. Technometrics, 53(4):440–441, November 2011. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
41714957.

Shi:2016:DNV

[SAR16] Zhenyu Shi, Daniel W. Apley, and George C. Runger. Discovering the nature of variation in nonlinear profile data. Technometrics, 58(3):371–??, 2016. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Sauter:2010:BRB

[Sau10] Roger M. Sauter. Book review: Data Manipulation With R (1st ed.) by Phil Spector. Technometrics, 52(3):365, August 2010. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/27867263.

Sauter:2011:BRB

[Sau11] Roger M. Sauter. Book review: R for SAS and SPSS Users (1st ed) by Robert A. Muenchen. Technometrics, 53(1):101, February 2011. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/40997301.

Sauter:2012:BRB

[Sau12] Roger M. Sauter. Book review: Data Mining With Rattle and R (1st ed.) by Graham Williams. Technometrics, 54(3):323– 324, August 2012. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
41714903.

Shmueli:2010:SCF

[SB10] Galit Shmueli and Howard Burkom. Statistical challenges facing early outbreak detection in biosurveillance. Technometrics, 52 (1):39–51, February 2010. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/40586679.

Singpurwalla:2019:BRF

[SB19] Nozer D. Singpurwalla and Lai Boya. Book review: Foundations of Info-Metrics (Modeling, Inference and Imperfect Information). Technometrics, 61(4):566–567, 2019. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Schneider:2017:MLE

[SCH17] Grant Schneider, Peter F. Craigmile, and Radu Herbei. Maximum likelihood estimation for stochastic differential equations using sequential Gaussian-process-based optimization. Technometrics, 59(2):178–188, 2017. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Schoonhoven:2012:RSD

[SD12] Marit Schoonhoven and Ronald J. M. M. Does. A robust standard deviation control chart. Technometrics, 54(1):73–82, February 2012. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/41714833.

Smucker:2015:AMS

[SD15] Byran J. Smucker and Nathan M. Drew. Approximate model spaces for model-robust experiment design. Technometrics, 57 (1):54–63, 2015. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Smucker:2012:MRT

[SdCR12] Byran J. Smucker, Enrique del Castillo, and James L. Rosenberger. Model-robust two-level designs using coordinate exchange algorithms and a maximin criterion. Technometrics, 54(4):367– 375, November 2012. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
41714921.

Subrahmaniam:2015:SSR

[SDR15] Vignesh T. Subrahmaniam, Anup Dewanji, and Bimal K. Roy. A semiparametric software reliability model for analysis of a bugdatabase with multiple defect types. Technometrics, 57(4):576– 585, 2015. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00401706.2014.947004.

Seaver:2011:BRBa

[Sea11a] William Seaver. Book review: A SAS/IML Companion for Linear Models by Jamis J. Perret. Technometrics, 53(1):107, February 2011. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/40997311.

Seaver:2011:BRBb

[Sea11b] William Seaver. Book review: Time Series: Modeling, Computation, and Inference by Raquel Prado; Mike West. Technometrics, 53(3):328, August 2011. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/23210418.

Seaver:2012:BRB

[Sea12] W. L. Seaver. Book review: Statistical Programming with SAS/ IML Software by Rick Wicklin. Technometrics, 54(2):201–202, May 2012. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/41714874.

Sengupta:2011:BRB

[Sen11] Sailes K. Sengupta. Book review: Digital Noise Monitoring of Defect Origin by Telman Aliev. Technometrics, 53(1):99, February 2011. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/40997296.

Sen:2010:RGM

[SF10] Ananda Sen and Arthur Fries. [Reliability Growth Management Metrics and Statistical Methods for Discrete-Use Systems]: Comment. Technometrics, 52(4):393–395, November 2010. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/40997244.

Storlie:2013:MCC

[SFH^{+}13] Curtis B. Storlie, Michael L. Fugate, David M. Higdon, Aparna V. Huzurbazar, Elizabeth G. François, and Douglas C. McHugh.

Methods for characterizing and comparing populations of shock wave curves. Technometrics, 55(4):436–449, 2013. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Sarkar:2018:SMT

[SG18] Soham Sarkar and Anil K. Ghosh. Some multivariate tests of independence based on ranks of nearest neighbors. Technometrics, 60(1):101–111, 2018. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Swihart:2014:RLR

[SGC14] Bruce J. Swihart, Jeff Goldsmith, and Ciprian M. Crainiceanu. Restricted likelihood ratio tests for functional effects in the functional linear model. Technometrics, 56(4):483–493, 2014. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Stanfill:2013:PEC

[SGH13] Bryan Stanfill, Ulrike Genschel, and Heike Hofmann. Point estimation of the central orientation of random rotations. Technometrics, 55(4):524–535, 2013. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Sartono:2015:CGO

[SGS15] Bagus Sartono, Peter Goos, and Eric Schoen. Constructing general orthogonal fractional factorial split-plot designs. Technometrics, 57(4):488–502, 2015. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00401706.2014.958198.

Smidl:2014:ESM

[SH14] Václav Smídl and Radek Hofman. Efficient sequential Monte Carlo sampling for continuous monitoring of a radiation situation. Technometrics, 56(4):514–528, 2014. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Sabbaghi:2018:BMB

[SHD18] Arman Sabbaghi, Qiang Huang, and Tirthankar Dasgupta. Bayesian model building from small samples of disparate data for capturing in-plane deviation in additive manufacturing. Technometrics, 60(4):532–544, 2018. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Shen:2017:RPD

[She17] Weijie Shen. Robust parameter designs in computer experiments using stochastic approximation. Technometrics, 59(4):471–483, 2017. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Sun:2012:NIP

[SHG12] Ying Sun, Jeffrey D. Hart, and Marc G. Genton. Nonparametric inference for periodic sequences. Technometrics, 54(1):83– 96, February 2012. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
41714834.

Samarov:2015:SLA

[SHL15] Daniel V. Samarov, Jeeseong Hwang, and Maritoni Litorja. The spatial LASSO with applications to unmixing hyperspectral biomedical images. Technometrics, 57(4):503–513, 2015. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00401706.2014.979950.

Singpurwalla:2010:RGM

[Sin10] Nozer D. Singpurwalla. [Reliability Growth Management Metrics and Statistical Methods for Discrete-Use Systems]: Comment. Technometrics, 52(4):396–397, November 2010. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/40997245.

Steinberg:2012:CDS

[SJ12] David M. Steinberg and Bradley Jones. Comment: DoIt — some thoughts on how to do it [MR2967968]. Technometrics, 54(3): 236–238, August 2012. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/41714891.

Shen:2018:CMD

[SL18] Weijie Shen and Dennis K. J. Lin. A conjugate model for dimensional analysis. Technometrics, 60(1):79–89, 2018. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Suzuki:2010:OSS

[SNM10] Kazuyuki Suzuki, Toshie Nakamoto, and Yohtaro Matsuo. Optimum specimen sizes and sample allocation for estimating Weibull shape parameters for two competing failure modes. Technometrics, 52(2):209–220, May 2010. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/27867225.

Sudjianto:2010:SMF

[SNY^{+}10] Agus Sudjianto, Sheela Nair, Ming Yuan, Aijun Zhang, Daniel Kern, and Fernando Cela-Díaz. Statistical methods for fighting financial crimes. Technometrics, 52(1):5–19, February 2010. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/40586676. See discussion

[Han10].

Singpurwalla:2018:LSS

[SPS18] Nozer D. Singpurwalla, Nicholas G. Polson, and Refik Soyer. From least squares to signal processing and particle filtering. Technometrics, 60(2):146–160, 2018. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Sosina:2019:RSO

[SRZ^{+}19] Sobambo Sosina, E. Marielle Remillard, Qiaoying Zhang, Chad Vecitis, and Tirthankar Dasgupta. Response surface optimization in the presence of internal noise with application to optimal alignment of carbon nanotubes. Technometrics, 61(1):50–65, 2019. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Somboonsavatdee:2015:SIP

[SS15] Anupap Somboonsavatdee and Ananda Sen. Statistical inference for power-law process with competing risks. Technometrics, 57 (1):112–122, 2015. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Sartono:2015:BOD

[SSG15] Bagus Sartono, Eric Schoen, and Peter Goos. Blocking orthogonal designs with mixed integer linear programming. Technometrics, 57(3):428–439, 2015. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00401706.2014.938832.

Steinberg:2010:ER

[Ste10a] David M. Steinberg. Editor’s report. Technometrics, 52(4):376– 378, November 2010. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
40997240.

Steinberg:2010:MIA

[Ste10b] David M. Steinberg. Mini-issue on anomaly detection. Technometrics, 52(1):4, February 2010. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/40586675.

Shen:2016:SSM

[STZW16] Xiaobei Shen, Kwok-Leung Tsui, Changliang Zou, and William H. Woodall. Self-starting monitoring scheme for Poisson count data with varying population sizes. Technometrics, 58(4):460–471, 2016. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Sun:2016:FAL

[SWF16] Ying Sun, Huixia J. Wang, and Montserrat Fuentes. Fused adaptive lasso for spatial and temporal quantile function estimation. Technometrics, 58(1):127–137, 2016. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/00401706.2015.1017115.

Sklar:2013:NRB

[SWMW13] Jeffrey C. Sklar, Junqing Wu, Wendy Meiring, and Yuedong Wang. Nonparametric regression with basis selection from multiple libraries. Technometrics, 55(2):189–201, 2013. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Schifano:2016:OUS

[SWW^{+}16] Elizabeth D. Schifano, Jing Wu, Chun Wang, Jun Yan, and Ming-Hui Chen. Online updating of statistical inference in the big data setting. Technometrics, 58(3):393–??, 2016. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Sikaroudi:2018:DSP

[SWW^{+}18] Ali Esmaieeli Sikaroudi, David A. Welch, Taylor J. Woehl, Roland Faller, James E. Evans, Nigel D. Browning, and Chiwoo Park. Directional statistics of preferential orientations of two shapes in their aggregate and its application to nanoparticle aggregation. Technometrics, 60(3):332–344, 2018. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Scinto:2014:CNG

[SWWR14] Philip R. Scinto, Robert G. Wilkinson, Zhen Wang, and Andrew Rose. Comment: Need for guidelines on appropriate screening designs for practitioners. Technometrics, 56(1):23–24, 2014. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Sylvestre:2012:MWH

[Syl12] Edward A. Sylvestre. In memoriam: William H. Lawton. Technometrics, 54(4):447, November 2012. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/41714943.

Symanzik:2011:BRB

[Sym11] Jürgen Symanzik. Book review: Introduction to Data Technologies by Paul Murrell. Technometrics, 53(2):214–215, May 2011. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/23209739.

Si:2019:MDM

[SYW19] Wujun Si, Qingyu Yang, and Xin Wu. Material degradation modeling and failure prediction using microstructure images. Technometrics, 61(2):246–258, 2019. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Smaga:2019:LHT

[SZ19] Lukasz Smaga and Jin-Ting Zhang. Linear hypothesis testing with functional data. Technometrics, 61(1):99–110, 2019. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Taddy:2013:MPS

[Tad13] Matt Taddy. Measuring political sentiment on Twitter: Factor optimal design for multinomial inverse regression. Technometrics, 55(4):415–425, 2013. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Tan:2013:MDF

[Tan13] Matthias H. Y. Tan. Minimax designs for finite design regions. Technometrics, 55(3):346–358, 2013. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Tan:2015:RPD

[Tan15a] Matthias Hwai Yong Tan. Robust parameter design with computer experiments using orthonormal polynomials. Technometrics, 57(4):468–478, 2015. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00401706.2014.969446.

Tan:2015:SPI

[Tan15b] Matthias Hwai Yong Tan. Stochastic polynomial interpolation for uncertainty quantification with computer experiments. Technometrics, 57(4):457–467, 2015. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/00401706.2014.950431.

Tan:2016:MQR

[Tan16] Matthias H. Y. Tan. Monotonic quantile regression with Bernstein polynomials for stochastic simulation. Technometrics, 58 (2):180–190, 2016. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Tan:2017:MMD

[Tan17] Matthias Hwai Yong Tan. Monotonic metamodels for deterministic computer experiments. Technometrics, 59(1):1–10, 2017. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Tan:2018:GPM

[Tan18] Matthias Hy Tan. Gaussian process modeling of a functional output with information from boundary and initial conditions and analytical approximations. Technometrics, 60(2):209–221, 2018. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Tang:2015:TSP

[TBMM15] Runlong Tang, Moulinath Banerjee, George Michailidis, and Shawn Mankad. Two-stage plans for estimating the inverse of a monotone function. Technometrics, 57(3):395–407, 2015. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00401706.2014.940773.

Tseng:2017:GQM

[TC17] Sheng-Tsaing Tseng and Pei-Yu Chen. A generalized quasi-MMSE controller for run-to-run dynamic models. Technometrics, 59(3):381–390, 2017. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Taylor:2014:TST

[TEN14] Sarah L. Taylor, Idris A. Eckley, and Matthew A. Nunes. A test of stationarity for textured images. Technometrics, 56(3):291–301, 2014. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Tsai:2010:GCF

[TG10] Pi-Wen Tsai and Steven G. Gilmour. A general criterion for factorial designs under model uncertainty. Technometrics, 52(2): 231–242, May 2010. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
27867227.

Trinca:2015:ISP

[TG15] Luzia A. Trinca and Steven G. Gilmour. Improved split-plot and multistratum designs. Technometrics, 57(2):145–154, 2015. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Trinca:2017:SPM

[TG17] Luzia A. Trinca and Steven G. Gilmour. Split-plot and multistratum designs for statistical inference. Technometrics, 59(4): 446–457, 2017. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Tsiamyrtzis:2010:BSP

[TH10] Panagiotis Tsiamyrtzis and Douglas M. Hawkins. Bayesian startup phase mean monitoring of an autocorrelated process that is subject to random sized jumps. Technometrics, 52(4):438–452, November 2010. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/

40997251. With supplementary material available online.

Tzeng:2018:RAF

[TH18] ShengLi Tzeng and Hsin-Cheng Huang. Resolution adaptive fixed rank kriging. Technometrics, 60(2):198–208, 2018. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Theiler:2013:MPM

[The13] James Theiler. Matched-pair machine learning. Technometrics, 55(4):536–547, 2013. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Tian:2015:SMI

[THS15] Siva Tian, Jianhua Z. Huang, and Haipeng Shen. Solving the MEG inverse problem: A robust two-way regularization method. Technometrics, 57(1):123–137, 2015. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Tibshirani:2011:NIR

[THT11] Ryan J. Tibshirani, Holger Hoefling, and Robert Tibshirani. Nearly-isotonic regression. Technometrics, 53(1):54–61, February 2011. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/40997292.

Tseng:2016:OAR

[TL16] Sheng-Tsaing Tseng and I-Chen Lee. Optimum allocation rule for accelerated degradation tests with a class of exponentialdispersion degradation models. Technometrics, 58(2):244–254, 2016. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Tan:2019:GPM

[TL19] Matthias H. Y. Tan and Guilin Li. Gaussian process modeling using the principle of superposition. Technometrics, 61(2):202– 218, 2019. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Tupper:2018:BDC

[TMAZ18] Laura L. Tupper, David S. Matteson, C. Lindsay Anderson, and Luckny Zephyr. Band depth clustering for nonstationary time series and wind speed behavior. Technometrics, 60(2):245–254, 2018. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Tseng:2016:MEC

[TML16] Sheng-Tsaing Tseng, Hsin-Chao Mi, and I.-Chen Lee. A multivariate EWMA controller for linear dynamic processes. Technometrics, 58(1):104–115, 2016. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/00401706.2015.1006795.

Tian:2017:SFI

[TMMH17] Ye Tian, Ranjan Maitra, William Q. Meeker, and Stephen D. Holland. A statistical framework for improved automatic flaw detection in nondestructive evaluation images. Technometrics, 59 (2):247–261, 2017. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Tuo:2013:CBM

[TQW13] Rui Tuo, Peter Z. G. Qian, and C. F. Jeff Wu. Comment: A Brownian motion model for stochastic simulation with tunable precision. Technometrics, 55(1):29–31, 2013. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Tan:2012:BAI

[TS12] Matthias H. Y. Tan and Jianjun Shi. A Bayesian approach for interpreting mean shifts in multivariate quality control. Technometrics, 54(3):294–307, August 2012. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/41714897.

Tibshirani:2016:OLS

[TS16] Robert Tibshirani and Xiaotong Suo. An ordered lasso and sparse time-lagged regression. Technometrics, 58(4):415–423, 2016. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Tsung:2010:NPM

[Tsu10] F. Tsung. [Nonparametric Profile Monitoring by Mixed Effects Modeling]: Comment. Technometrics, 52(3):283–285, August 2010. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/27867248.

Tan:2012:RDO

[TW12] Matthias H. Y. Tan and C. F. Jeff Wu. Robust design optimization with quadratic loss derived from Gaussian process models. Technometrics, 54(1):51–63, February 2012. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://

www.jstor.org/stable/41714831.

Tan:2013:BAM

[TW13] Matthias H. Y. Tan and C. F. Jeff Wu. A Bayesian approach for model selection in fractionated split plot experiments with applications in robust parameter design. Technometrics, 55(3): 359–372, 2013. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Tuo:2014:SMC

[TWY14] Rui Tuo, C. F. Jeff Wu, and Dan Yu. Surrogate modeling of computer experiments with different mesh densities. Technometrics, 56(3):372–380, 2014. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

VanBrackle:2010:BRB

[Van10] Lewis VanBrackle. Book review: A Comprehensive Guide to Factorial Two-Level Experimentation by Robert W. Mee. Technometrics, 52(4):465, November 2010. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/40997259.

Vazquez:2019:CTLa

[VGS19] Alan R. Vazquez, Peter Goos, and Eric D. Schoen. Constructing two-level designs by concatenation of strength-3 orthogonal arrays. Technometrics, 61(2):219–232, 2019. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Vaughan:2013:NWS

[VSM13] Joel Vaughan, Stilian Stoev, and George Michailidis. Networkwide statistical modeling, prediction, and monitoring of computer traffic. Technometrics, 55(1):79–93, 2013. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Wiel:2011:ROM

[VWGR11] Scott Vander Wiel, Alyson Wilson, Todd Graves, and Shane Reese. A random onset model for degradation of high-reliability systems. Technometrics, 53(2):163–172, May 2011. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/23209728.

Vazquez:2019:CTL

[VX19] Alan R. Vazquez and Hongquan Xu. Construction of two-level nonregular designs of strength three with large run sizes. Technometrics, 61(3):341–353, 2019. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Wilson:2010:RGM

[WAC10] Alyson G. Wilson and Christine M. Anderson-Cook. [Reliability Growth Management Metrics and Statistical Methods for Discrete-Use Systems]: Comment. Technometrics, 52(4):397–400, November 2010. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
40997246.

Wadsworth:2016:ESM

[Wad16] J. L. Wadsworth. Exploiting structure of maximum likelihood estimators for extreme value threshold selection. Technometrics, 58(1):116–126, 2016. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00401706.2014.998345.

Wolters:2011:SAM

[WB11] Mark A. Wolters and Derek Bingham. Simulated annealing model search for subset selection in screening experiments. Technometrics, 53(3):225–237, August 2011. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/23210399.

Wang:2018:DPG

[WB18] Zimo Wang and Satish T. S. Bukkapatnam. A Dirichlet process Gaussian state machine model for change detection in transient processes. Technometrics, 60(3):373–385, 2018. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Woodall:2010:NPM

[WBD10] William H. Woodall, Jeffrey B. Birch, and Pang Du. [Nonparametric Profile Monitoring by Mixed Effects Modeling]: Comment. Technometrics, 52(3):285–287, August 2010. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://

www.jstor.org/stable/27867249.

Winterfors:2012:BME

[WC12] Emanuel Winterfors and Andrew Curtis. A bifocal measure of expected ambiguity in Bayesian nonlinear parameter estimation. Technometrics, 54(2):179–190, May 2012. CODEN TCMTA2.

ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://

www.jstor.org/stable/41714866.

Wang:2019:CSI

[WH19] Wenjia Wang and Benjamin Haaland. Controlling sources of inaccuracy in stochastic kriging. Technometrics, 61(3):309–321, 2019. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Wilson:2018:EUF

[WHQ18] Kevin J. Wilson, Daniel A. Henderson, and John Quigley. Emulation of utility functions over a set of permutations: Sequencing reliability growth tasks. Technometrics, 60(3):273–285, 2018. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Wiener:2011:BRB

[Wie11] Matthew C. Wiener. Book review: Design & Analysis of Vaccine Studies by M. E. Halloran; I. M. Longini; C. J. Struchiner. Technometrics, 53(1):106–107, February 2011. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://

www.jstor.org/stable/40997310.

Woods:2014:R

[WLDD14] David C. Woods, Susan M. Lewis, Angela M. Dean, and Danel Draguljić. Rejoinder. Technometrics, 56(1):25–28, 2014. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Wan:2015:ECS

[WLHB15] Fang Wan, Wei Liu, Yang Han, and Frank Bretz. An exact confidence set for a maximum point of a univariate polynomial function in a given interval. Technometrics, 57(4):559–565, 2015. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00401706.2014.962708.

Wludyka:2012:BRB

[Wlu12] Peter Wludyka. Book review: JMP Means Business: Statistical Models for Management by Josef Schmee; Jane E. Oppenlander. Technometrics, 54(3):320, August 2012. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://

www.jstor.org/stable/41714900.

Wong:2016:QUL

[WLWZ16] Samuel W. K. Wong, Conroy Lum, Lang Wu, and James V. Zidek. Quantifying uncertainty in lumber grading and strength prediction: A Bayesian approach. Technometrics, 58(2):236–243, 2016. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Weaver:2013:MPR

[WMEW13] Brian P. Weaver, William Q. Meeker, Luis A. Escobar, and Joanne Wendelberger. Methods for planning repeated measures degradation studies. Technometrics, 55(2):122–134, 2013. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Wludyka:2011:BRB

[WMI11] Peter Wludyka and Carmen Masnita-Iusan. Book review: Mathematical and Statistical Estimation Approaches in Epidemiology by Gerardo Chowell; James M. Hyman; Luis M. A. Bettencourt; Carlos Castillo-Chavez. Technometrics, 53(1):105–106, February 2011. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/40997309.

Wang:2018:TMP

[WMP18] Yuan Wang, Yajun Mei, and Kamran Paynabar. Thresholded multivariate principal component analysis for Phase I multichannel profile monitoring. Technometrics, 60(3):360–372, 2018. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Wu:2012:FFD

[WMT12] Huaiqing Wu, Robert Mee, and Boxin Tang. Fractional factorial designs with admissible sets of clear two-factor interactions. Technometrics, 54(2):191–197, May 2012. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://

www.jstor.org/stable/41714867.

Wang:2018:UHM

[WND^{+}18] Xueou Wang, David J. Nott, Christopher C. Drovandi, Kerrie Mengersen, and Michael Evans. Using history matching for prior choice. Technometrics, 60(4):445–460, 2018. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Wang:2012:CRC

[WSZ12] Jiaping Wang, Haipeng Shen, and Hongtu Zhu. [Clustering Random Curves Under Spatial Interdependence With Application to Service Accessibility]: Comment. Technometrics, 54(2):129– 133, May 2012. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/

41714860. See [JS12a].

Witten:2014:CEN

[WSZ14] Daniela M. Witten, Ali Shojaie, and Fan Zhang. The cluster elastic net for high-dimensional regression with unknown variable grouping. Technometrics, 56(1):112–122, 2014. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Woods:2011:BDE

[Wv11] David C. Woods and Peter van de Ven. Blocked designs for experiments with correlated non-normal response. Technometrics, 53(2):173–182, May 2011. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/23209729.

Wang:2018:IGD

[WW18] Bing Xing Wang and Fangtao Wu. Inference on the gamma distribution. Technometrics, 60(2):235–244, 2018. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Wang:2010:IGP

[WX10] Xiao Wang and Dihua Xu. An inverse Gaussian process model for degradation data. Technometrics, 52(2):188–197, May 2010. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/27867223. With supple-

mentary material available online.

Wang:2014:PSI

[WX14] Tianhao Wang and Yingcun Xia. A piecewise single-index model for dimension reduction. Technometrics, 56(3):312–324, 2014. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Wang:2018:AFR

[WYHT18] Xin Wang, Zhi-Sheng Ye, Yi-Li Hong, and Loon-Ching Tang. Analysis of field return data with failed-but-not-reported events.

Technometrics, 60(1):90–100, 2018. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Wang:2010:IUP

[WYJ10] Bing Xing Wang, Keming Yu, and M. C. Jones. Inference under progressively Type II right-censored sampling for certain lifetime distributions. Technometrics, 52(4):453–460, November 2010. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/40997252.

Wylupek:2010:DDS

[Wył10] Grzegorz Wyłupek. Data-driven k-sample tests. Technometrics, 52(1):107–123, February 2010. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/40586684.

Wang:2019:MKD

[WZ19] Boxiang Wang and Hui Zou. A multicategory kernel distance weighted discrimination method for multiclass classification. Technometrics, 61(3):396–408, 2019. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Xu:2010:LAM

[XC10] Huiping Xu and Bruce A. Craig. Likelihood analysis of multivariate probit models using a parameter expanded MCEM algorithm. Technometrics, 52(3):340–348, August 2010. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://

www.jstor.org/stable/27867256.

Xiong:2016:OED

[XDHQ16] Shifeng Xiong, Bin Dai, Jared Huling, and Peter Z. G. Qian. Orthogonalizing EM: A design-based least squares algorithm. Technometrics, 58(3):285–??, 2016. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Xu:2017:MLT

[XHM^{+}17] Zhibing Xu, Yili Hong, William Q. Meeker, Brock E. Osborn, and Kati Illouz. A multi-level trend-renewal process for modeling systems with recurrence data. Technometrics, 59(2):225–236, 2017. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Xu:2017:VCM

[XHX17] Maochao Xu, Lei Hua, and Shouhuai Xu. A vine copula model for predicting the effectiveness of cyber defense early-warning. Technometrics, 59(4):508–520, 2017. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Xiong:2010:SNN

[Xio10] Shifeng Xiong. Some notes on the nonnegative garrote. Technometrics, 52(3):349–361, August 2010. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/27867257.

Xie:2018:SMA

[XKHY18] Yimeng Xie, Caleb B. King, Yili Hong, and Qingyu Yang. Semiparametric models for accelerated destructive degradation test data analysis. Technometrics, 60(2):222–234, 2018. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Xia:2016:MIU

[XMM16] Donggeng Xia, Shawn Mankad, and George Michailidis. Measuring influence of users in Twitter ecosystems using a counting process modeling framework. Technometrics, 58(3):360–??, 2016. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Xiong:2013:SDA

[XQW13] Shifeng Xiong, Peter Z. G. Qian, and C. F. Jeff Wu. Sequential design and analysis of high-accuracy and low-accuracy computer codes. Technometrics, 55(1):37–46, 2013. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Xiang:2017:SPC

[XTP17] Dongdong Xiang, Fugee Tsung, and Xiaolong Pu. Statistical process control for latent quality characteristics using the upand-down test. Technometrics, 59(4):496–507, 2017. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Xian:2018:NAS

[XWL18] Xiaochen Xian, Andi Wang, and Kaibo Liu. A nonparametric adaptive sampling strategy for online monitoring of big data streams. Technometrics, 60(1):14–25, 2018. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Ye:2014:IGP

[YC14] Zhi-Sheng Ye and Nan Chen. The inverse Gaussian process as a degradation model. Technometrics, 56(3):302–311, 2014. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Yuan:2017:MRQ

[YCZ17] Yuan Yuan, Nan Chen, and Shiyu Zhou. Modeling regression quantile process using monotone B-splines. Technometrics, 59 (3):338–350, 2017. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Ye:2010:BRB

[Ye10] Keying Ye. Book review: Statistical Design by George Casella. Technometrics, 52(4):462, November 2010. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://

www.jstor.org/stable/40997254.

Young:2010:BRB

[You10] Derek S. Young. Book review: Statistical Tolerance Regions: Theory, Applications, and Computation by K. Krishnamoorthy; T. Mathew. Technometrics, 52(1):143–144, February 2010. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/40586694.

Yan:2019:SPC

[YPP19] Hao Yan, Kamran Paynabar, and Massimo Pacella. Structured point cloud data analysis via regularized tensor regression for process modeling and optimization. Technometrics, 61(3):385– 395, 2019. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Yan:2017:ADI

[YPS17] Hao Yan, Kamran Paynabar, and Jianjun Shi. Anomaly detection in images with smooth background via smooth-sparse decomposition. Technometrics, 59(1):102–114, 2017. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Yan:2018:RTM

[YPS18] Hao Yan, Kamran Paynabar, and Jianjun Shi. Real-time monitoring of high-dimensional functional data streams via spatiotemporal smooth sparse decomposition. Technometrics, 60(2):

181–197, 2018. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Young:2011:SEL

[YR11] P. C. Young and M. Ratto. Statistical emulation of large linear dynamic models. Technometrics, 53(1):29–43, February 2011. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/40997290. Sup-

plementary materials available online.

Ye:2016:AUF

[YT16] Zhi-Sheng Ye and Loon-Ching Tang. Augmenting the unreturned for field data with information on returned failures only. Technometrics, 58(4):513–523, 2016. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Yu:2012:BRB

[Yu12] Jihnhee Yu. Book review: An Introduction to Applied Multivariate Analysis With R by B. Everitt; T. Hothorn. Technometrics, 54(4):445, November 2012. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/41714937.

Ye:2014:SEG

[YXTC14] Zhi-Sheng Ye, Min Xie, Loon-Ching Tang, and Nan Chen. Semiparametric estimation of gamma processes for deteriorating products. Technometrics, 56(4):504–513, 2014. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Ye:2012:DBB

[YXTS12] Zhi-Sheng Ye, Min Xie, Loon-Ching Tang, and Yan Shen. Degradation-based burn-in planning under competing risks. Technometrics, 54(2):159–168, May 2012. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://

www.jstor.org/stable/41714864.

Yu:2012:ODF

[YZW12] Guan Yu, Changliang Zou, and Zhaojun Wang. Outlier detection in functional observations with applications to profile monitoring. Technometrics, 54(3):308–318, August 2012. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://

www.jstor.org/stable/41714898.

Yang:2019:BAA

[YZW19] Chun-Hao Yang, James V. Zidek, and Samuel W. K. Wong. Bayesian analysis of accumulated damage models in lumber reliability. Technometrics, 61(2):233–245, 2019. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Yang:2018:FER

[YZZ18] Yi Yang, Teng Zhang, and Hui Zou. Flexible expectile regression in reproducing kernel Hilbert spaces. Technometrics, 60(1):26– 35, 2018. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Zhu:2018:MPO

[ZC18] Guangyu Zhu and Jiahua Chen. Multi-parameter one-sided monitoring tests. Technometrics, 60(3):398–407, 2018. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Zhang:2016:SLN

[ZCG^{+}16] Chunming Zhang, Yi Chai, Xiao Guo, Muhong Gao, David Devilbiss, and Zhengjun Zhang. Statistical learning of neuronal functional connectivity. Technometrics, 58(3):350–??, 2016. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Zhu:2018:UAS

[ZCM^{+}18] Hongxiao Zhu, Philip Caspers, Jeffrey S. Morris, Xiaowei Wu, and Rolf Müller. A unified analysis of structured sonar-terrain data using Bayesian functional mixed models. Technometrics, 60 (1):112–123, 2018. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Zhang:2019:IEV

[ZCW19] Bohai Zhang, Noel Cressie, and Debra Wunch. Inference for errors-in-variables models in the presence of systematic errors with an application to a satellite remote sensing campaign. Technometrics, 61(2):187–201, 2019. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Zhu:2014:DOD

[ZDH14] Li Zhu, Tirthankar Dasgupta, and Qiang Huang. A D-optimal design for estimation of parameters of an exponential-linear growth curve of nanostructures. Technometrics, 56(4):432–442, 2014.

CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Zhang:2010:IEG

[Zha10] Jin Zhang. Improving on estimation for the generalized Pareto distribution. Technometrics, 52(3):335–339, August 2010. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/27867255.

Zhang:2011:TWM

[Zha11] Jin-Ting Zhang. Two-way MANOVA with unequal cell sizes and unequal cell covariance matrices. Technometrics, 53(4):426–439, November 2011. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
41714955.

Zhang:2013:AHD

[ZHOY13] Ke Zhang, Jacqueline M. Hughes-Oliver, and S. Stanley Young. Analysis of high-dimensional structure-activity screening datasets using the optimal bit string tree. Technometrics, 55(2):161–173, 2013. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Zou:2011:LBD

[ZJT11] Changliang Zou, Wei Jiang, and Fugee Tsung. A LASSO-based diagnostic framework for multivariate statistical process control. Technometrics, 53(3):297–309, August 2011. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://

www.jstor.org/stable/23210405.

Zhang:2017:TEP

[ZL17] Xin Zhang and Lexin Li. Tensor envelope partial least-squares regression. Technometrics, 59(4):426–436, 2017. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Zhang:2016:VSL

[ZLD16] Yukun Zhang, Xuewen Lu, and Anthony F. Desmond. Variable selection in a log-linear Birnbaum–Saunders regression model for high-dimensional survival data via the elastic-net and stochastic EM. Technometrics, 58(3):383–??, 2016. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Zhang:2019:EIS

[ZM19] Xin Zhang and Qing Mai. Efficient integration of sufficient dimension reduction and prediction in discriminant analysis. Technometrics, 61(2):259–272, 2019. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Zimmer:2014:PST

[ZPM14] Zachary Zimmer, DoHwan Park, and Thomas Mathew. Pointwise and simultaneous tolerance limits under logistic regression. Technometrics, 56(3):282–290, 2014. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Zang:2018:PMS

[ZQ18] Yangyang Zang and Peihua Qiu. Phase I monitoring of spatial surface data from 3D printing. Technometrics, 60(2):169–180, 2018. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Zhou:2011:SAE

[ZQZ11] Qiang Zhou, Peter Z. G. Qian, and Shiyu Zhou. A simple approach to emulation for computer models with qualitative and quantitative factors. Technometrics, 53(3):266–273, August 2011. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/23210402.

Zhong:2015:MDA

[ZS15] Wenxuan Zhong and Kenneth S. Suslick. Matrix discriminant analysis with application to colorimetric sensor array data. Technometrics, 57(4):524–534, 2015. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/00401706.2014.965347.

Zhou:2014:FTW

[ZSGM14] Rensheng R. Zhou, Nicoleta Serban, Nagi Gebraeel, and Hans-Georg Müller. A functional time warping approach to modeling and monitoring truncated degradation signals. Technometrics, 56(1):67–77, 2014. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Zou:2011:MSE

[ZT11] Changliang Zou and Fugee Tsung. A multivariate sign EWMA control chart. Technometrics, 53(1):84–97, February 2011. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/40997295. Supplemen-

tary materials available online.

Zhang:2017:EPC

[ZTEM17] Wei Zhang, Ye Tian, Luis A. Escobar, and William Q. Meeker. Estimating a parametric component lifetime distribution from a collection of superimposed renewal processes. Technometrics, 59 (2):202–214, 2017. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Zhang:2017:SRC

[ZW17] Xiang Zhang and Alyson Wilson. System reliability and component importance under dependence: A copula approach. Technometrics, 59(2):215–224, 2017. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Zou:2015:EOM

[ZWZJ15] Changliang Zou, Zhaojun Wang, Xuemin Zi, and Wei Jiang. An efficient online monitoring method for high-dimensional data streams. Technometrics, 57(3):374–387, 2015. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic).  URL http://www.tandfonline.com/doi/abs/10.
1080/00401706.2014.940089.

Zhang:2017:AIF

[ZX17] Mimi Zhang and Min Xie. An ameliorated improvement factor model for imperfect maintenance and its goodness of fit. Technometrics, 59(2):237–246, 2017. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Zhang:2017:BCT

[ZY17] Tonglin Zhang and Baijian Yang. Box–Cox transformation in big data. Technometrics, 59(2):189–201, 2017. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Zhai:2018:DCD

[ZY18] Qingqing Zhai and Zhi-Sheng Ye. Degradation in common dynamic environments. Technometrics, 60(4):461–471, 2018. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Zhu:2014:PEW

[ZYH14] Yada Zhu, Emmanuel Yashchin, and J. R. M. Hosking. Parametric estimation for window censored recurrence data. Technometrics, 56(1):55–66, 2014. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Zhou:2014:DSC

[ZZD^{+}14] Qiang Zhou, Junyi Zhou, Michael De Cicco, Shiyu Zhou, and Xiaochun Li. Detecting 3D spatial clustering of particles in nanocomposites based on cross-sectional images. Technometrics, 56(2):212–224, 2014. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Zou:2015:TLA

[ZZZ^{+}15] Na Zou, Yun Zhu, Ji Zhu, Mustafa Baydogan, Wei Wang, and Jing Li. A transfer learning approach for predictive modeling of degenerate biological systems. Technometrics, 57(3):362–373, 2015. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00401706.2015.1044117.